### v1

# Geometric Integration by Parts and Sobolev Spaces on Vector Bundles: A Unified Global Approach

Abstract
This article develops a unified framework for the theory of Sobolev spaces on vector bundles over riemannian manifolds. The analytical core of our approach is a rigorous higher-order geometric integration by parts formula, which characterizes the formal adjoint of the covariant derivative. This identity is established for arbitrary manifolds, requiring no assumptions on completeness or compactness. While these results are fundamental to global analysis, explicit and direct proofs are often elusive in the literature or rely on overly sophisticated machinery that overshadows the underlying geometry. To bridge this gap, we establish sharp local-to-global norm equivalence estimates and provide streamlined, self-contained proofs for the Meyers-Serrin theorem on general manifolds, as well as the Sobolev embedding and Rellich-Kondrashov theorems for the compact case. By prioritizing intrinsic global arguments over ad hoc coordinate patching, this work provides a modern and accessible foundation for the study of Sobolev spaces on bundles.
Geometric Integration by Parts and Sobolev Spaces on Vector Bundles: A Unified Global Approach
Velázquez-Mendoza Carlos Daniel
Department of Mathematics. Facultad de Ciencias, UNAM. Circuito Exterior s/n. México City. 04510. México.
danielc21@ciencias.unam.mx
Sandoval-Romero María de los Ángeles (Corresponding Author)
Department of Mathematics. Facultad de Ciencias, UNAM. Circuito Exterior s/n. México City. 04510. México.
selegna@ciencias.unam.mx
Keywords: Sobolev spaces, Vector bundles, Geometric integration by parts, Embedding Theorems, Rellich-Kondrashov Theorem.
1.  Introduction
The analysis of Sobolev spaces on vector bundles is a foundational pillar of modern geometric analysis. A central issue in this theory is ensuring the equivalence between the two standard definitions of these spaces: the completion of the space of smooth sections with respect to the Sobolev norm, denoted by $H^{m,p}(E)$, and the space of sections possessing weak covariant derivatives in $L^{p}$, denoted by $W^{m,p}(E)$.
In Riemannian manifolds, the treatment of this equivalence often varies significantly across the literature. Standard texts such as [1] and [4] typically define Sobolev spaces directly as completions of smooth sections, largely bypassing the intrinsic difficulties of weak derivatives. In contrast, approaches that prioritize intrinsic geometry, such as [7], define them via distributional derivatives. Establishing the identity $H^{m,p}(E)=W^{m,p}(E)$—the Meyers-Serrin theorem—is therefore essential to reconcile these perspectives. While this equivalence is well-documented for scalar functions or under specific geometric assumptions (see [2]), the treatment for general vector bundles requires rigorous handling of the fiber metric and the connection.
Existing proofs of the fundamental structural theorems—density, embedding, and compactness—often suffer from a dichotomy between generality and geometric clarity. For instance, the density of smooth sections is established in [3] for arbitrary vector bundles, but the framework relies on the theory of general linear partial differential operators and requires local elliptic regularity. Conversely, regarding the Sobolev embedding and Rellich-Kondrashov theorems, the seminal work of Palais [8] remains the most general reference for compact manifolds. Despite its importance, the proofs in [8] involve a complex reduction to the $n$-dimensional torus that can obscure the intrinsic geometry of the manifold.
The purpose of this paper is to provide a unified, constructive, and strictly geometric framework for the theory of Sobolev spaces on vector bundles. We avoid reliance on elliptic regularity or indirect embeddings into tori. Instead, we develop the theory from the intrinsic definition of the weak covariant derivative, utilizing a rigorous integration by parts formula as the analytical core.
Our approach yields three main contributions: First, in Section 3, we derive a higher-order integration by parts formula for the covariant derivative acting on sections of vector bundles (Theorem 1). Beyond its global significance, this identity provides a constructive path to characterize the formal adjoint of the covariant derivative. By working in arbitrary local frames, we expose the precise dependence of the adjoint on the structural properties of the connection and the fiber metrics, resolving the combinatorial complexity that arises at higher orders. Second, in Section 5, we present a self-contained proof of the Meyers–Serrin theorem valid for general Riemannian manifolds. Unlike previous approaches, our proof rests solely on the geometric integration by parts formula and standard convolution arguments within local charts, bridged by sharp local-to-global norm equivalence estimates. Finally, for the case of compact manifolds, in Section 6, we utilize this framework to provide streamlined proofs of the continuous Sobolev embedding theorem and the Rellich–Kondrashov compactness theorem. By prioritizing intrinsic global arguments over ad hoc coordinate patching, this work offers a rigorous foundation for global analysis compatible with modern geometric applications.
2.  Notation and Preliminaries
2.1 Notation and Conventions
Throughout this paper, $(M,g)$ denotes a smooth Riemannian manifold (not necessarily complete or compact), and $d\lambda_{g}$ denotes the Riemannian-Lebesgue measure induced by the metric $g$. If $M$ is a manifold with boundary, its boundary is denoted by $\partial M$.
Let $E\to M$ be a smooth vector bundle of finite rank equipped with a smooth fiber metric $h_{E}$ and a compatible connection $\nabla^{E}$. Following the notation in [5], $\Gamma(E)$ denotes the space of smooth sections of $E$, while $\Gamma_{c}(E)$ denotes the subspace of smooth sections with compact support. For the particular case of scalar-valued functions on a domain $\Omega\subseteq\mathbb{R}^{n}$, we employ the standard notation $C_{c}^{\infty}(\Omega)$.
For any integer $s\geq 0$, we denote by $T^{(0,s)}(TM)\otimes E$ the bundle of $E$-valued covariant tensor fields of rank $s$. Any connection on $TM$ together with $\nabla^{E}$ induces a connection on these tensor bundles, also denoted by $\nabla$, which is uniquely characterized by the Leibniz rule. We write $|\cdot|_{g,h_{E}}$ to denote the pointwise norm on the fibers induced by $g$ and $h_{E}$; note that for any section $\sigma$, $|\sigma|_{g,h_{E}}:M\to[0,\infty)$ is a non-negative function on $M$, which allows us to define the global norm of a section in the corresponding function spaces.
Finally, we adopt the Einstein summation convention, although summations will be written explicitly whenever deemed necessary for clarity or when the convention is not applicable.
2.2 Sobolev Spaces
We begin with the precise definitions for the Sobolev spaces used in this work.
Definition 1 (The space $H^{m,p}$).
Let $m\geq 0$ be an integer and $1\leq p<\infty$. We define the space of smooth sections with finite Sobolev norm, denoted by $S^{m,p}(E)$, as:
$S^{m,p}(E):=\left\{u\in\Gamma(E)\mathrel{\Big|}\|u\|_{m,p}<\infty\right\},$
where the norm is given by
$\|u\|_{m,p}:=\left(\sum_{s=0}^{m}\int_{M}|\nabla^{s}u|_{h_{E}}^{p}\,d\lambda_{g}\right)^{1/p}.$
Here, $\nabla^{s}u$ denotes the $s$-th iterated covariant derivative of $u$, viewed as a section of the bundle $T^{(0,s)}(TM)\otimes E$. The Sobolev space $H^{m,p}(E)$ is defined as the completion of $S^{m,p}(E)$ with respect to this norm.
To introduce the definition via weak covariant derivatives, we first recall the notion of the formal adjoint.
Definition 2 (Weak Covariant Derivative and the space $W^{m,p}$).
Let $s\geq 0$. The formal adjoint of the iterated covariant derivative $\nabla^{s}$, denoted by $(\nabla^{s})^{*}$, is the unique differential operator characterized by the identity
$\int_{M}\langle\nabla^{s}u,v\rangle_{h_{E}}\,d\lambda_{g}=\int_{M}\langle u,(\nabla^{s})^{*}v\rangle_{h_{E}}\,d\lambda_{g},$ (1)
for all $u\in\Gamma_{c}(E)$ and $v\in\Gamma_{c}(T^{(0,s)}(TM)\otimes E)$.
A section $u\in L^{1}_{\operatorname{loc}}(E)$ is said to admit a weak covariant derivative of order $s$ if there exists a section $v\in L^{1}_{\operatorname{loc}}(T^{(0,s)}(TM)\otimes E)$ such that
$\int_{M}\langle v,\psi\rangle_{h_{E}}\,d\lambda_{g}=\int_{M}\langle u,(\nabla^{s})^{*}\psi\rangle_{h_{E}}\,d\lambda_{g}$
holds for all test sections $\psi\in\Gamma_{c}(T^{(0,s)}(TM)\otimes E)$. If such a section $v$ exists, it is unique almost everywhere; we denote it by $\nabla_{w}^{s}u$.
Finally, the Sobolev space $W^{m,p}(E)$ is defined as:
[…]
This space is equipped with the same norm $\|\cdot\|_{m,p}$ defined in Definition 1.
To establish an explicit local characterization of the formal adjoint $(\nabla^{s})^{*}$, we first derive the local coordinate expression of the covariant derivative acting on tensor-valued sections.
Lemma 1.
Let $(M,g)$ be a Riemannian manifold and let $\nabla$ denote a connection on $TM$. Let $E\longrightarrow M$ be a smooth vector bundle with a connection $\nabla^{E}$. We denote by $\nabla$ the induced connection on the tensor product bundle $T^{(k,l)}(TM)\otimes E$, defined uniquely by the Leibniz rule:
$\nabla_{X}(F\otimes\sigma)=(\nabla_{X}F)\otimes\sigma+F\otimes\nabla^{E}_{X}\sigma,$
for all $X\in\Gamma(TM)$, $F\in\Gamma(T^{(k,l)}(TM))$, and $\sigma\in\Gamma(E)$.
Let $(E_{1},\dots,E_{n})$ be a local frame for $TM$ over an open set $U$, with dual coframe $(\varepsilon^{1},\dots,\varepsilon^{n})$, and let $(e_{1},\dots,e_{r})$ be a local frame for $E$ over $U$. If a section $F\in\Gamma(T^{(k,l)}(TM)\otimes E)$ is expressed locally as
$F=F^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}\,E_{i_{1}}\otimes\cdots\otimes E_{i_{k}}\otimes\varepsilon^{j_{1}}\otimes\cdots\otimes\varepsilon^{j_{l}}\otimes e_{a},$
then the components of the covariant derivative $\nabla_{E_{c}}F$ with respect to these frames are given by
$(\nabla_{E_{c}}F)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}$ $=E_{c}\big(F^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}\big)+\sum_{t=1}^{k}\Gamma^{i_{t}}_{cs}\,F^{i_{1}\dots s\dots i_{k}a}_{j_{1}\dots j_{l}}$
$\quad-\sum_{t=1}^{l}\Gamma^{s}_{cj_{t}}\,F^{i_{1}\dots i_{k}\,a}_{j_{1}\dots s\dots j_{l}}+\sum_{b=1}^{r}(A_{c})^{a}_{b}\,F^{i_{1}\dots i_{k}b}_{j_{1}\dots j_{l}},$
where the connection coefficients are defined by $\nabla_{E_{c}}E_{j}=\Gamma^{s}_{cj}E_{s}$ and $\nabla^{E}_{E_{c}}e_{b}=(A_{c})^{a}_{b}\,e_{a}$.
Proof.
We have
$\nabla_{E_{c}}F=\nabla_{E_{c}}\left(\sum_{a=1}^{r}F^{(a)}\otimes e_{a}\right),$
where $F^{(a)}\in\Gamma(T^{(k,l)}(TM))$ denotes the tensor field with components $(F^{(a)})^{i_{1}\dots i_{k}}_{j_{1}\dots j_{l}}=F^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}$. Applying the Leibniz rule for the induced connection yields
$\nabla_{E_{c}}F=\sum_{a=1}^{r}\left((\nabla_{E_{c}}F^{(a)})\otimes e_{a}+F^{(a)}\otimes\nabla^{E}_{E_{c}}e_{a}\right).$ (2)
Recall that for the tensor field $F^{(a)}$, the connection components are given by (see, e.g., [6, Proposition 4.16]):
$(\nabla_{E_{c}}F^{(a)})^{i_{1}\dots i_{k}}_{j_{1}\dots j_{l}}=E_{c}(F^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}})+\sum_{t=1}^{k}\Gamma^{i_{t}}_{cs}\,F^{i_{1}\dots s\dots i_{k}a}_{j_{1}\dots j_{l}}-\sum_{t=1}^{l}\Gamma^{s}_{cj_{t}}\,F^{i_{1}\dots i_{k}\,a}_{j_{1}\dots s\dots j_{l}}.$
For the second term in (2), we substitute $\nabla^{E}_{E_{c}}e_{b}=(A_{c})^{a}_{b}e_{a}$. Relabeling the summation index for the bundle frame as $b$, we obtain
$\sum_{b=1}^{r}F^{(b)}\otimes\nabla^{E}_{E_{c}}e_{b}=\sum_{b=1}^{r}F^{(b)}\otimes\left(\sum_{a=1}^{r}(A_{c})^{a}_{b}e_{a}\right)=\sum_{a=1}^{r}\left(\sum_{b=1}^{r}(A_{c})^{a}_{b}F^{(b)}\right)\otimes e_{a}.$
The component of this term along the basis element $e_{a}$ is $\sum_{b=1}^{r}(A_{c})^{a}_{b}F^{i_{1}\dots i_{k}b}_{j_{1}\dots j_{l}}$. Combining these contributions yields the asserted formula. ∎
Remark 1.
We regard the total covariant derivative $\nabla F$ as a section of $T^{(k,l+1)}(TM)\otimes E$. In terms of the local frames defined above, we write
$\nabla F=\sum_{c=1}^{n}(\nabla_{E_{c}}F)\otimes\varepsilon^{c}.$
Consequently, the components of $\nabla F$ are defined by $(\nabla F)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}c}:=(\nabla_{E_{c}}F)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}$. Explicitly,
$(\nabla F)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}c}$ $=E_{c}\big(F^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}\big)+\sum_{t=1}^{k}\Gamma^{i_{t}}_{cs}\,F^{i_{1}\dots s\dots i_{k}a}_{j_{1}\dots j_{l}}$
$\quad-\sum_{t=1}^{l}\Gamma^{s}_{cj_{t}}\,F^{i_{1}\dots i_{k}\,a}_{j_{1}\dots s\dots j_{l}}+\sum_{b=1}^{r}(A_{c})^{a}_{b}\,F^{i_{1}\dots i_{k}b}_{j_{1}\dots j_{l}}.$
We emphasize that we adopt the convention where the differentiation index (here $c$) corresponds to the last covariant index in the component representation.
3.  The Geometric Integration by Parts Formula
With the local properties of the connection established, we now turn to the analytical core of this work. The following theorem provides a rigorous global integration by parts formula for higher-order covariant derivatives. Unlike standard Euclidean formulas, this result accounts for the intrinsic curvature of the bundle through the connection and explicitly identifies the boundary pairings required for the study of boundary value problems.
Theorem 1 (Integration by Parts for Higher-Order Derivatives).
Let $(M,g)$ be a Riemannian manifold with boundary $\partial M$, and let $\nabla$ denote the Levi-Civita connection on $TM$. Let $E\to M$ be a smooth vector bundle equipped with a fiber metric $h_{E}$ and a connection $\nabla^{E}$ compatible with $h_{E}$. We continue to denote by $\nabla$ the induced connection on the tensor bundles $T^{(k,l)}(TM)\otimes E$.
For any integer $s\geq 1$, and sections $F\in\Gamma_{c}\bigl(T^{(k,l)}(TM)\otimes E\bigr)$ and $G\in\Gamma_{c}\bigl(T^{(k,l+s)}(TM)\otimes E\bigr)$, the following identity holds:
$\int_{M}\langle\nabla^{s}F,G\rangle_{g,h_{E}}\,d\lambda_{g}$ $=(-1)^{s}\int_{M}\big\langle F,(\operatorname{tr}_{g}\circ\nabla)^{s}G\big\rangle_{g,h_{E}}\,d\lambda_{g}$
$\quad+\sum_{j=0}^{s-1}(-1)^{\,s-1-j}\int_{\partial M}\Big\langle\nabla^{j}F,\,\iota_{\nu}\bigl((\operatorname{tr}_{g}\circ\nabla)^{\,s-1-j}G\bigr)\Big\rangle_{g,h_{E}}\,d\lambda_{\widetilde{g}}$,
where $\nu$ denotes the outward unit normal to $\partial M$, and $\iota_{\nu}$ denotes the contraction (interior multiplication) of $\nu$ into the last covariant index of the tensor factor. Explicitly, if $H\in\Gamma\bigl(T^{(k,m+1)}(TM)\otimes E\bigr)$, then $\iota_{\nu}H\in\Gamma\bigl(T^{(k,m)}(TM)\otimes E\bigr)$ is given in local coordinates by
$\bigl(\iota_{\nu}H\bigr)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{m}}:=\nu^{q}\,H^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{m}q},$
where $\nu=\nu^{q}\partial_{q}$.
Here, $\operatorname{tr}_{g}$ denotes the metric contraction of the last two covariant indices of $\nabla H$. Specifically, if $H\in\Gamma\bigl(T^{(k,m)}(TM)\otimes E\bigr)$, then $\nabla H\in\Gamma\bigl(T^{(k,m+1)}(TM)\otimes E\bigr)$ and locally
$\bigl(\operatorname{tr}_{g}(\nabla H)\bigr)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{m-1}}:=g^{pq}\,(\nabla H)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{m-1}pq}.$
That is, the derivative index is contracted with the last covariant index inherited from $H$.
Proof.
We proceed by induction on $s$. The strategy is to shift derivatives from $F$ to $G$ via covariant integration by parts, tracking the boundary terms that arise at each step.
Base Case ($s=1$). Fix a point $p\in M$. Since $\nabla$ on $TM$ is the Levi-Civita connection, we can choose geodesic normal coordinates centered at $p$, with coordinate frame $(\partial_{1},\dots,\partial_{n})$. In these coordinates, $g_{ij}(p)=\delta_{ij}$ and all connection coefficients vanish at $p$. In particular, $\nabla_{\partial_{i}}\partial_{j}(p)=0$. Furthermore, we choose a local orthonormal frame $(e_{1},\dots,e_{r})$ for the bundle $E$ defined in a neighborhood of $p$, such that $h_{E}(e_{a},e_{b})(p)=\delta_{ab}$.
Let $F\in\Gamma_{c}\bigl(T^{(k,l)}(TM)\otimes E\bigr)$ and $G\in\Gamma_{c}\bigl(T^{(k,l+1)}(TM)\otimes E\bigr)$. We define a map $\Phi:\mathfrak{X}(M)\longrightarrow C^{\infty}(M)$ by
$\Phi(Y):=\big\langle F,\,\iota_{Y}G\big\rangle_{g,h_{E}}.$
The map $Y\mapsto\iota_{Y}G$ is $C^{\infty}(M)$-linear in $Y$ and the inner product is bilinear, so $\Phi$ defines a 1-form. Let $X:=\Phi^{\sharp}$ be the dual vector field, characterized by $\langle X,Y\rangle_{g}=\Phi(Y)=\big\langle F,\,\iota_{Y}G\big\rangle_{g,h_{E}}$. In particular, for each coordinate vector $\partial_{c}$,
$\langle X,\partial_{c}\rangle_{g}=\big\langle F,\,\iota_{\partial_{c}}G\big\rangle_{g,h_{E}}.$ (3)
At $p$, the divergence of $X$ is given by
$\operatorname{div}X(p)=\sum_{c=1}^{n}\big\langle\nabla_{\partial_{c}}X,\partial_{c}\big\rangle_{g}(p).$
Using the compatibility of $\nabla$ with $g$ and the fact that $\nabla_{\partial_{c}}\partial_{c}(p)=0$, we obtain
$\operatorname{div}X(p)=\sum_{c=1}^{n}\partial_{c}\big(\langle X,\partial_{c}\rangle_{g}\big)(p).$ (4)
Substituting (3) into (4) and using the product rule:
$\operatorname{div}X(p)=\sum_{c=1}^{n}\Big(\big\langle\nabla_{\partial_{c}}F,\,\iota_{\partial_{c}}G\big\rangle_{g,h_{E}}(p)+\big\langle F,\,\nabla_{\partial_{c}}\bigl(\iota_{\partial_{c}}G\bigr)\big\rangle_{g,h_{E}}(p)\Big).$
Using the Leibniz rule for contraction $\nabla_{\partial_{c}}\bigl(\iota_{\partial_{c}}G\bigr)=\iota_{\partial_{c}}\bigl(\nabla_{\partial_{c}}G\bigr)+\iota_{\nabla_{\partial_{c}}\partial_{c}}G$, and noting that the second term vanishes at $p$, we arrive at:
$\operatorname{div}X(p)=\sum_{c=1}^{n}\Big(\big\langle\nabla_{\partial_{c}}F,\,\iota_{\partial_{c}}G\big\rangle_{g,h_{E}}(p)+\big\langle F,\,\iota_{\partial_{c}}(\nabla_{\partial_{c}}G)\big\rangle_{g,h_{E}}(p)\Big).$
We now identify these sums using local coordinates. Let
$F=F^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}\partial_{i_{1}}\otimes\dots\otimes\partial_{i_{k}}\otimes dx^{j_{1}}\otimes\cdots\otimes dx^{j_{l}}\otimes e_{a},$
$G=G^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}t}\partial_{i_{1}}\otimes\dots\otimes\partial_{i_{k}}\otimes dx^{j_{1}}\otimes\cdots\otimes dx^{j_{l}}\otimes dx^{t}\otimes e_{a}.$
At $p$, where $g_{ij}=\delta_{ij}$ and $h_{ab}=\delta_{ab}$, the first sum becomes:
$\sum_{c=1}^{n}\big\langle\nabla_{\partial_{c}}F,\,\iota_{\partial_{c}}G\big\rangle_{g,h_{E}}(p)$ $=\sum_{c=1}^{n}\sum_{a=1}^{r}\sum_{\begin{subarray}{c}1\leq i_{1},\dots,i_{k}\leq n\\ 1\leq j_{1},\dots,j_{l}\leq n\end{subarray}}(\nabla_{\partial_{c}}F)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}(p)\,G^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}c}(p)$
$=\big\langle\nabla F,\,G\big\rangle_{g,h_{E}}(p).$
For the second sum, noting that $(\nabla_{\partial_{c}}G)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}t}=(\nabla G)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}t\,c}$, we have:
$\sum_{c=1}^{n}\big\langle F,\,\iota_{\partial_{c}}(\nabla_{\partial_{c}}G\bigr)\big\rangle_{g,h_{E}}(p)$ $=\sum_{c=1}^{n}\sum_{a=1}^{r}\sum_{\begin{subarray}{c}1\leq i_{1},\dots,i_{k}\leq n\\ 1\leq j_{1},\dots,j_{l}\leq n\end{subarray}}F^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}(p)\,(\nabla G)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}c\,c}(p).$
Since $g^{pq}(p)=\delta^{pq}$ and $\bigl(\operatorname{tr}_{g}(\nabla G)\bigr)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}=g^{pq}\,(\nabla G)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}pq}$, we have
$\bigl(\operatorname{tr}_{g}(\nabla G)\bigr)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}(p)=\sum_{c=1}^{n}(\nabla G)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}c\,c}(p).$
Thus,
$\sum_{c=1}^{n}\big\langle F,\,\iota_{\partial_{c}}(\nabla_{\partial_{c}}G)\big\rangle_{g,h_{E}}(p)=\big\langle F,\,\operatorname{tr}_{g}(\nabla G)\big\rangle_{g,h_{E}}(p).$
Combining these results, we have $\operatorname{div}X=\langle\nabla F,G\rangle_{g,h_{E}}+\langle F,\operatorname{tr}_{g}(\nabla G)\rangle_{g,h_{E}}$ globally. Integrating and applying the Divergence Theorem yields:
$\int_{M}\big\langle\nabla F,G\big\rangle_{g,h_{E}}\,d\lambda_{g}=-\int_{M}\big\langle F,\operatorname{tr}_{g}(\nabla G)\big\rangle_{g,h_{E}}\,d\lambda_{g}+\int_{\partial M}\big\langle F,\,\iota_{\nu}G\big\rangle_{g,h_{E}}\,d\lambda_{\widetilde{g}}$.
Inductive Step. Assume the identity holds for some $s\geq 1$. Let $G\in\Gamma_{c}\bigl(T^{(k,l+s+1)}(TM)\otimes E\bigr)$. Applying the base case ($s=1$) to $(\nabla^{s}F,G)$, where $\nabla^{s}F\in\Gamma_{c}(T^{(k,l+s)}(TM)\otimes E)$:
$\int_{M}\langle\nabla^{s+1}F,G\rangle_{g,h_{E}}\,d\lambda_{g}=-\int_{M}\big\langle\nabla^{s}F,\operatorname{tr}_{g}(\nabla G)\big\rangle_{g,h_{E}}\,d\lambda_{g}+\int_{\partial M}\big\langle\nabla^{s}F,\,\iota_{\nu}G\big\rangle_{g,h_{E}}\,d\lambda_{\widetilde{g}}$.
Define $H:=\operatorname{tr}_{g}(\nabla G)$. Applying the inductive hypothesis to $(F,H)$, we obtain:
$\int_{M}\langle\nabla^{s}F,H\rangle_{g,h_{E}}\,d\lambda_{g}$ $=(-1)^{s}\int_{M}\big\langle F,(\operatorname{tr}_{g}\circ\nabla)^{s}H\big\rangle_{g,h_{E}}\,d\lambda_{g}$
$\quad+\sum_{j=0}^{s-1}(-1)^{\,s-1-j}\int_{\partial M}\Big\langle\nabla^{j}F,\,\iota_{\nu}\bigl((\operatorname{tr}_{g}\circ\nabla)^{\,s-1-j}H\bigr)\Big\rangle_{g,h_{E}}\,d\lambda_{\widetilde{g}}$.
Substituting $H=\operatorname{tr}_{g}(\nabla G)$, we observe that $(\operatorname{tr}_{g}\circ\nabla)^{s}H=(\operatorname{tr}_{g}\circ\nabla)^{s+1}G$ and $(\operatorname{tr}_{g}\circ\nabla)^{\,s-1-j}H=(\operatorname{tr}_{g}\circ\nabla)^{\,s-j}G$. Substituting this result back into the expansion of $\int_{M}\langle\nabla^{s+1}F,G\rangle_{g,h_{E}}\,d\lambda_{g}$, the interior integral term becomes $(-1)^{s+1}\int_{M}\big\langle F,(\operatorname{tr}_{g}\circ\nabla)^{s+1}G\big\rangle_{g,h_{E}}\,d\lambda_{g}$. The boundary sum becomes:
[…]
$=\sum_{j=0}^{s-1}(-1)^{\,s-j}\int_{\partial M}\Big\langle\nabla^{j}F,\,\iota_{\nu}\bigl((\operatorname{tr}_{g}\circ\nabla)^{\,s-j}G\bigr)\Big\rangle_{g,h_{E}}\,d\lambda_{\widetilde{g}}$.
Finally, adding the isolated boundary term from the first step $\int_{\partial M}\big\langle\nabla^{s}F,\,\iota_{\nu}G\big\rangle_{g,h_{E}}\,d\lambda_{\widetilde{g}}$ (which corresponds to $j=s$ since $(-1)^{s-s}=1$ and $(\operatorname{tr}\circ\nabla)^{0}=\operatorname{Id}$), we extend the sum index to $s$:
$\sum_{j=0}^{s}(-1)^{\,s-j}\int_{\partial M}\Big\langle\nabla^{j}F,\,\iota_{\nu}\bigl((\operatorname{tr}_{g}\circ\nabla)^{\,s-j}G\bigr)\Big\rangle_{g,h_{E}}\,d\lambda_{\widetilde{g}}$.
This yields exactly the formula for $s+1$, completing the induction. ∎
To handle the multi-indices rigorously, for a fixed $s\geq 1$, we denote ordered tuples (blocks) of indices in $\{1,\dots,n\}$ as $I_{s}=(i_{1},\dots,i_{s})$, $J_{s}=(j_{1},\dots,j_{s})$, etc. We adopt the Einstein summation convention for these blocks. We define:
$\partial_{I_{s}}:=\partial_{i_{s}}\cdots\partial_{i_{1}},\quad g^{I_{s}J_{s}}:=\prod_{l=1}^{s}g^{i_{l}j_{l}},\quad g_{J_{s}T_{s}}:=\prod_{l=1}^{s}g_{j_{l}t_{l}}.$
Also, let $\beta_{I_{s}}\in\mathbb{N}_{0}^{n}$ be the multi-index associated with the block $I_{s}$ such that $\partial^{\beta_{I_{s}}}\equiv\partial_{I_{s}}$.
The integration by parts formula derived above is of fundamental importance because it allows us to derive an explicit coordinate expression for the formal adjoint of the $s$-th covariant derivative.
Corollary 1 (Local Expression of the Formal Adjoint).
Let $(M,g)$ be a Riemannian manifold without boundary, and let $E\to M$ be a vector bundle equipped with a fiber metric $h_{E}$ and a compatible connection $\nabla^{E}$. For $s\geq 1$, the formal adjoint of
$\nabla^{s}:\Gamma(E)\longrightarrow\Gamma\bigl(T^{(0,s)}(TM)\otimes E\bigr)$
is given globally by
$(\nabla^{s})^{*}=(-1)^{s}(\operatorname{tr}_{g}\circ\nabla)^{s}.$
Furthermore, in a local chart $(U,x^{1},\dots,x^{n})$, if a section $G\in\Gamma\bigl(T^{(0,s)}(TM)\otimes E\bigr)$ has components $G^{a}_{J_{s}}:=G^{a}_{j_{1}\dots j_{s}}$, then
$\begin{split}\bigl((\nabla^{s})^{*}G\bigr)^{a}&=(-1)^{s}\Bigg[\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(\sqrt{\det(g)}\,g^{I_{s}J_{s}}\,G^{a}_{J_{s}}\Big)\\ &\quad+\sum_{b=1}^{r}\sum_{|\beta|\leq s-1}\bigl(C_{\beta}\bigr)^{a\,J_{s}}_{b}\,\partial^{\beta}\!\bigl(G^{b}_{J_{s}}\bigr)\Bigg],\end{split}$ (5)
where the coefficients $\bigl(C_{\beta}\bigr)^{a\,J_{s}}_{b}$ are smooth functions on $U$ depending solely on the metric components $g_{ij}$ and its inverse $g^{ij}$, the connection coefficients $\Gamma^{k}_{ij}$ and $(A_{i})^{a}_{b}$, and their partial derivatives up to order s1s-1.
Proof.
Since $\partial M=\emptyset$, the boundary terms in Theorem 1 vanish. Consequently, for all $F\in\Gamma_{c}(E)$ and $G\in\Gamma_{c}\bigl(T^{(0,s)}(TM)\otimes E\bigr)$, we have
$\int_{M}\langle\nabla^{s}F,G\rangle_{g,h_{E}}\,d\lambda_{g}=(-1)^{s}\int_{M}\big\langle F,(\operatorname{tr}_{g}\circ\nabla)^{s}G\big\rangle_{g,h_{E}}\,d\lambda_{g}.$
By the uniqueness of the formal adjoint (see, e.g., [7]), it follows that
$(\nabla^{s})^{*}=(-1)^{s}(\operatorname{tr}_{g}\circ\nabla)^{s}.$
It remains to verify the local formula (5). Fix a chart $(U,x^{1},\dots,x^{n})$ and a local frame $(e_{1},\dots,e_{r})$ for $E|_{U}$. Let $m\geq 0$ and let $H\in\Gamma\bigl(T^{(0,m+1)}(TM)\otimes E\bigr)$ with components $H^{a}_{J_{m}q}=H^{a}_{j_{1}\dots j_{m}q}$. By definition, the components of the trace are
$((\operatorname{tr}_{g}\circ\nabla)H)^{a}_{J_{m}}=g^{pq}\,(\nabla_{\partial_{p}}H)^{a}_{J_{m}q}.$
Applying Lemma 1 (with $k=0$, $l=m+1$), we expand the covariant derivative:
$\begin{split}((\operatorname{tr}_{g}\circ\nabla)H)^{a}_{J_{m}}&=g^{pq}\partial_{p}\bigl(H^{a}_{J_{m}q}\bigr)-\sum_{\ell=1}^{m}\Gamma^{\alpha}_{pj_{\ell}}\,g^{pq}\,H^{a}_{j_{1}\dots j_{\ell-1}\alpha j_{\ell+1}\dots j_{m}q}\\ &\quad-\Gamma^{\alpha}_{pq}\,g^{pq}\,H^{a}_{J_{m}\alpha}+\sum_{b=1}^{r}(A_{p})^{a}_{b}\,g^{pq}H^{b}_{J_{m}q}.\end{split}$ (6)
We recall the identity $\Gamma_{\alpha p}^{\alpha}=\frac{1}{\sqrt{\det(g)}}\partial_{p}\left(\sqrt{\det(g)}\right)$ and the formula for the derivative of the inverse metric, $\partial_{p}(g^{pq})=-\Gamma^{p}_{p\alpha}g^{\alpha q}-\Gamma^{q}_{p\alpha}g^{p\alpha}$. These imply the following divergence identity for any collection of functions $(Y_{q})_{q=1}^{n}$:
$g^{pq}\partial_{p}\bigl(Y_{q}\bigr)-\Gamma^{\alpha}_{pq}g^{pq}Y_{\alpha}=\frac{1}{\sqrt{\det(g)}}\,\partial_{p}\Big(\sqrt{\det(g)}\,g^{pq}Y_{q}\Big).$ (7)
Indeed, expanding the right-hand side of (7):
$\frac{1}{\sqrt{\det(g)}}\partial_{p}$ $\left(\sqrt{\det(g)}g^{pq}Y_{q}\right)$
$=\frac{1}{\sqrt{\det(g)}}\bigg[\partial_{p}\left(\sqrt{\det(g)}\right)g^{pq}Y_{q}+\sqrt{\det(g)}\,\partial_{p}(g^{pq})Y_{q}+\sqrt{\det(g)}\,g^{pq}\partial_{p}(Y_{q})\bigg]$
$=\Gamma_{\alpha p}^{\alpha}g^{pq}Y_{q}+\partial_{p}(g^{pq})Y_{q}+g^{pq}\partial_{p}(Y_{q})$
$=\Gamma_{\alpha p}^{\alpha}g^{pq}Y_{q}+\left(-\Gamma_{p\alpha}^{p}g^{\alpha q}Y_{q}-\Gamma_{p\alpha}^{q}g^{p\alpha}Y_{q}\right)+g^{pq}\partial_{p}(Y_{q}).$
Observing that the first term cancels with the second (after relabeling indices), and relabeling $\alpha\leftrightarrow q$ in the third term, we recover $g^{pq}\partial_{p}(Y_{q})-\Gamma_{pq}^{\alpha}g^{pq}Y_{\alpha}$.
Applying (7) with $Y_{q}=H^{a}_{J_{m}q}$, equation (6) simplifies to:
$\begin{split}((\operatorname{tr}_{g}\circ\nabla)H)^{a}_{J_{m}}&=\frac{1}{\sqrt{\det(g)}}\,\partial_{p}\Big(\sqrt{\det(g)}\,g^{pq}H^{a}_{J_{m}q}\Big)\\ &\quad-\sum_{\ell=1}^{m}\Gamma^{\alpha}_{pj_{\ell}}\,g^{pq}\,H^{a}_{j_{1}\dots j_{\ell-1}\alpha j_{\ell+1}\dots j_{m}q}+\sum_{b=1}^{r}(A_{p})^{a}_{b}\,g^{pq}H^{b}_{J_{m}q}.\end{split}$ (8)
We now prove by induction on $s\geq 1$ that for any $G\in\Gamma\bigl(T^{(0,s)}(TM)\otimes E\bigr)$, the following holds:
$\begin{split}\bigl((\operatorname{tr}_{g}\circ\nabla)^{s}G\bigr)^{a}&=\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(\sqrt{\det(g)}\,g^{I_{s}J_{s}}\,G^{a}_{J_{s}}\Big)\\ &\quad+\sum_{b=1}^{r}\sum_{|\beta|\leq s-1}\bigl(C_{\beta}\bigr)^{a\,J_{s}}_{b}\,\partial^{\beta}\!\bigl(G^{b}_{J_{s}}\bigr).\end{split}$ (9)
For $s=1$, setting $m=0$ in (8) yields
$((\operatorname{tr}_{g}\circ\nabla)G)^{a}=\frac{1}{\sqrt{\det(g)}}\,\partial_{p}\Big(\sqrt{\det(g)}\,g^{pq}G^{a}_{q}\Big)+\sum_{b=1}^{r}(A_{p})^{a}_{b}\,g^{pq}G^{b}_{q},$
which matches the form of (9) with $|\beta|=0$.
Now, assume (9) holds for some $s\geq 1$. We prove the case for $s+1$. Let $G\in\Gamma\bigl(T^{(0,s+1)}(TM)\otimes E\bigr)$ and define $H:=(\operatorname{tr}_{g}\circ\nabla)G\in\Gamma\bigl(T^{(0,s)}(TM)\otimes E\bigr)$. By the inductive hypothesis applied to $H$:
$\begin{split}\bigl((\operatorname{tr}_{g}\circ\nabla)^{s}H\bigr)^{a}&=\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(\sqrt{\det(g)}\,g^{I_{s}J_{s}}\,H^{a}_{J_{s}}\Big)\\ &\quad+\sum_{b=1}^{r}\sum_{|\beta|\leq s-1}\bigl(C_{\beta}\bigr)^{a\,J_{s}}_{b}\,\partial^{\beta}\!\bigl(H^{b}_{J_{s}}\bigr).\end{split}$ (10)
Since $(\operatorname{tr}_{g}\circ\nabla)^{s}H=(\operatorname{tr}_{g}\circ\nabla)^{s+1}G$, we expand $H^{a}_{J_{s}}$ in terms of $G$ using (8) with $m=s$:
$\begin{split}H^{a}_{J_{s}}&=\frac{1}{\sqrt{\det(g)}}\,\partial_{p}\Big(\sqrt{\det(g)}\,g^{pq}G^{a}_{J_{s}q}\Big)\\ &\quad-\sum_{\ell=1}^{s}\Gamma^{\alpha}_{pj_{\ell}}\,g^{pq}\,G^{a}_{j_{1}\dots j_{\ell-1}\alpha j_{\ell+1}\dots j_{s}q}+\sum_{b=1}^{r}(A_{p})^{a}_{b}\,g^{pq}G^{b}_{J_{s}q}.\end{split}$ (11)
Substituting (11) into the principal term of (10) splits it into three parts:
$\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(\sqrt{\det(g)}\,g^{I_{s}J_{s}}\,H^{a}_{J_{s}}\Big)=\mathrm{(I)}+\mathrm{(II)}+\mathrm{(III)},$ (12)
where
$\mathrm{(I)}$ $:=\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(g^{I_{s}J_{s}}\partial_{p}\big(\sqrt{\det(g)}\,g^{pq}G^{a}_{J_{s}q}\big)\Big),$
$\mathrm{(II)}$ $:=-\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(\sqrt{\det(g)}\,g^{I_{s}J_{s}}\sum_{\ell=1}^{s}\Gamma^{\alpha}_{pj_{\ell}}\,g^{pq}\,G^{a}_{j_{1}\dots j_{\ell-1}\alpha j_{\ell+1}\dots j_{s}q}\Big),$
$\mathrm{(III)}$ $:=\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(\sqrt{\det(g)}\,g^{I_{s}J_{s}}\sum_{b=1}^{r}(A_{p})^{a}_{b}\,g^{pq}G^{b}_{J_{s}q}\Big).$
For term $\mathrm{(I)}$, we utilize the identity $\phi\partial_{p}f=\partial_{p}(\phi f)-(\partial_{p}\phi)f$. Applying this with $\phi=g^{I_{s}J_{s}}$ and $f=\sqrt{\det(g)}\,g^{pq}G^{a}_{J_{s}q}$, we have:
$g^{I_{s}J_{s}}\partial_{p}\big(\sqrt{\det(g)}\,g^{pq}G^{a}_{J_{s}q}\big)=\partial_{p}\Big(g^{I_{s}J_{s}}\sqrt{\det(g)}\,g^{pq}G^{a}_{J_{s}q}\Big)-\partial_{p}\big(g^{I_{s}J_{s}}\big)\sqrt{\det(g)}\,g^{pq}G^{a}_{J_{s}q}.$
Substituting this back into $\mathrm{(I)}$:
$\mathrm{(I)}$ $=\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\partial_{p}\Big(\sqrt{\det(g)}\,g^{I_{s}J_{s}}g^{pq}\,G^{a}_{J_{s}q}\Big)$ (13)
$\quad-\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(\partial_{p}\big(g^{I_{s}J_{s}}\big)\sqrt{\det(g)}\,g^{pq}G^{a}_{J_{s}q}\Big).$
The first term of (13) is exactly the principal term required for the order $s+1$. If we define the block indices for order $s+1$ as $I_{s+1}=(I_{s},p)$ and $J_{s+1}=(J_{s},q)$, then $g^{I_{s+1}J_{s+1}}=g^{I_{s}J_{s}}g^{pq}$ and $\partial_{I_{s+1}}=\partial_{I_{s}}\partial_{p}$. Thus, this term is
$\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s+1}}\Big(\sqrt{\det(g)}\,g^{I_{s+1}J_{s+1}}\,G^{a}_{J_{s+1}}\Big).$
The second term of (13) involves $\partial_{I_{s}}$ acting on a term containing $G$ with no derivatives. By the Leibniz rule, the highest order of derivatives acting on $G$ in this term is $s$. Thus, there exist smooth coefficients $\bigl(B_{\beta}\bigr)$ such that this remainder term is of the form:
$\sum_{c=1}^{r}\sum_{|\beta|\leq s}\bigl(B_{\beta}\bigr)^{a\,J_{s+1}}_{c}\,\partial^{\beta}\!\bigl(G^{c}_{J_{s+1}}\bigr).$ (14)
For $\mathrm{(II)}$, the term inside $\partial_{I_{s}}$ contains $G$ algebraically (zero derivatives). Thus, applying $\partial_{I_{s}}$ yields terms with at most $s$ derivatives on $G$. Similarly for $\mathrm{(III)}$, the dependence on $G$ is linear without additional derivatives, so expanding yields terms of order at most $s$. Combining these, there exist coefficients $\bigl(\widetilde{B}_{\beta}\bigr)$ such that:
$\mathrm{(II)}+\mathrm{(III)}=\sum_{c=1}^{r}\sum_{|\beta|\leq s}\bigl(\widetilde{B}_{\beta}\bigr)^{a\,J_{s+1}}_{c}\,\partial^{\beta}\!\bigl(G^{c}_{J_{s+1}}\bigr).$ (15)
Finally, consider the lower-order term from (10):
$\sum_{b=1}^{r}\sum_{|\beta|\leq s-1}\bigl(C_{\beta}\bigr)^{a\,J_{s}}_{b}\,\partial^{\beta}\!\bigl(H^{b}_{J_{s}}\bigr).$
Substituting (11) into $H^{b}_{J_{s}}$, we see that $H$ involves first-order derivatives of $G$. Therefore, applying $\partial^{\beta}$ (where $|\beta|\leq s-1$) to $H$ yields derivatives of $G$ of order at most (s1)+1=s(s-1)+1=s. Thus, there exist coefficients $\bigl(\overline{D}_{\beta}\bigr)$ such that this sum reduces to:
$\sum_{c=1}^{r}\sum_{|\beta|\leq s}\bigl(\overline{D}_{\beta}\bigr)^{a\,J_{s+1}}_{c}\,\partial^{\beta}\!\bigl(G^{c}_{J_{s+1}}\bigr).$ (16)
Combining the principal term from (13) with the remainder terms (14), (15), and (16), we conclude that there exist smooth coefficients $\bigl(C_{\beta}\bigr)$ with $|\beta|\leq s$ such that
$\bigl((\operatorname{tr}_{g}\circ\nabla)^{s+1}G\bigr)^{a}$ $=\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s+1}}\Big(\sqrt{\det(g)}\,g^{I_{s+1}J_{s+1}}\,G^{a}_{J_{s+1}}\Big)$
$\quad+\sum_{c=1}^{r}\sum_{|\beta|\leq s}\bigl(C_{\beta}\bigr)^{a\,J_{s+1}}_{c}\,\partial^{\beta}\!\bigl(G^{c}_{J_{s+1}}\bigr),$
which is exactly the form of (9) for $s+1$. This completes the induction. ∎
4.  Local-to-Global Estimates and Structural Theorems
Having established the global integration by parts formula and the local structure of the formal adjoint, we now turn to the fundamental structural theorems of the theory. Our primary objective in this section is to establish a characterization of Sobolev sections on a coordinate chart in terms of their component functions with respect to a local frame (Lemma 3); this allows us to reduce intrinsic analytic problems to their Euclidean counterparts. To achieve this, we require a result ensuring norm equivalence for fiber metrics on vector bundles over compact manifolds. A similar equivalence is well known for the tangent bundle (see, e.g., [6]). This behavior extends to the Sobolev setting, where different choices of metric and connection yield equivalent Sobolev norms on compact manifolds, as we have previously established in [9, Theorem 2] for Sobolev spaces defined via completions.
Lemma 2.
Let $M$ be a smooth compact manifold (with or without boundary) and let $E\to M$ be a smooth vector bundle of finite rank. Let $h$ and $\widetilde{h}$ be two smooth fiber metrics on $E$. Then, there exist constants $c,C>0$ such that for all $p\in M$ and all $v\in E_{p}$,
$c\,h_{p}(v,v)\leq\widetilde{h}_{p}(v,v)\leq C\,h_{p}(v,v).$
Proof.
Let $p\in M$. Consider the unit sphere in the fiber $E_{p}$ with respect to the metric $h$:
$S_{p}:=\{v\in E_{p}\mid h_{p}(v,v)=1\}.$
Since $E_{p}$ is a finite-dimensional vector space, $S_{p}$ is a closed and bounded set, and thus compact. The function $f_{p}:S_{p}\longrightarrow\mathbb{R}$ defined by $f_{p}(v)=\widetilde{h}_{p}(v,v)$ is continuous, so it attains a maximum. Consequently, the function $f:M\longrightarrow\mathbb{R}$ given by
$f(p):=\max_{v\in S_{p}}f_{p}(v)$
is well-defined. Since the bundle and the metrics are smooth, $f$ is continuous on $M$. As $M$ is compact, there exists a point $p_{0}\in M$ such that $f(p)\leq f(p_{0})=:C$ for all $p\in M$.
Moreover,
$C=f(p_{0})=\max_{v\in S_{p_{0}}}\widetilde{h}_{p_{0}}(v,v)>0,$
since $\widetilde{h}$ is positive definite and $v\neq 0$ for all $v\in S_{p_{0}}$. Thus, for all $p\in M$ and all $v\in S_{p}$, we have
$\widetilde{h}_{p}(v,v)\leq C.$
By an analogous argument, considering the minimum, there exists a constant $c>0$ such that for all $p\in M$ and all $v\in S_{p}$,
$c\leq\widetilde{h}_{p}(v,v).$
Now, let $v\in E_{p}\setminus\{0\}$. By homogeneity, the vector $\frac{v}{|v|_{h_{p}}}$ belongs to $S_{p}$ (where $|v|_{h_{p}}=\sqrt{h_{p}(v,v)}$). Therefore,
$c\leq\widetilde{h}_{p}\!\left(\frac{v}{|v|_{h_{p}}},\frac{v}{|v|_{h_{p}}}\right)\leq C.$
Using the bilinearity of the metric, this implies
$c\,|v|^{2}_{h_{p}}\leq\frac{1}{|v|^{2}_{h_{p}}}\widetilde{h}_{p}(v,v)\leq C,$
or equivalently,
$c\,h_{p}(v,v)\leq\widetilde{h}_{p}(v,v)\leq C\,h_{p}(v,v).$
The case $v=0$ holds trivially, as all terms vanish. ∎
With the independence of the fiber metric established for compact domains, we now turn to the local analytical structure of the Sobolev spaces. The following lemma provides the crucial bridge between the intrinsic definition of $W^{m,p}(E)$—defined via weak covariant derivatives—and the standard Sobolev spaces of component functions in local coordinates. This equivalence allows us to import local analytic results, such as approximation theorems and embeddings, from Euclidean space into the manifold setting.
Lemma 3 (Local Norm Equivalence).
Let $(M,g)$ be a Riemannian manifold with Levi–Civita connection $\nabla$. Let $m\geq 1$, let $(U,\phi)$ be a regular coordinate ball in $M$, and let $E\to M$ be a smooth vector bundle of rank $r$ equipped with a fiber metric $h_{E}$ and a compatible connection $\nabla^{E}$. Let $F\in L^{1}_{\operatorname{loc}}(E)$ and let $\{e_{1},\dots,e_{r}\}$ be a local frame for $E$ over $U$, such that $F$ is expressed locally as $F=\displaystyle\sum_{a=1}^{r}F^{a}e_{a}$.
Then,
$F|_{U}\in W^{m,p}(E|_{U})\quad\text{if and only if}\quad F^{a}\circ\phi^{-1}\in W^{m,p}(\phi(U))$
for all $1\leq a\leq r$.
Furthermore, there exist constants $C_{m},c_{m}>0$ (independent of $F$) such that for all $F\in W^{m,p}(E|_{U})$,
$c_{m}\sum_{a=1}^{r}\left\|F^{a}\circ\phi^{-1}\right\|_{W^{m,p}(\phi(U))}\leq\|F\|_{W^{m,p}(E|_{U})}\leq C_{m}\sum_{a=1}^{r}\left\|F^{a}\circ\phi^{-1}\right\|_{W^{m,p}(\phi(U))}.$
Proof.
($\Rightarrow$) First, assume that $F|_{U}\in W^{m,p}(E|_{U})$. Then, for each $s\leq m$, the $s$-th weak covariant derivative of $F|_{U}$ exists; let us denote it by $G\in L^{p}\big(T^{(0,s)}(TU)\otimes E|_{U}\big)$. By definition, for all $\Psi\in\Gamma_{c}\big(T^{(0,s)}(TU)\otimes E|_{U}\big)$, the following holds:
$\int_{U}\langle F,(\nabla^{s})^{*}\Psi\rangle_{g,h_{E}}\,d\lambda_{g}=\int_{U}\langle G,\Psi\rangle_{g,h_{E}}\,d\lambda_{g}.$
In local coordinates, writing $F=F^{a}e_{a}$ and $\Psi=\Psi^{c}_{J_{s}}\,dx^{J_{s}}\otimes e_{c}$, this identity reads
$\int_{U}h_{bc}\,F^{b}\,\big((\nabla^{s})^{*}\Psi\big)^{c}\,d\lambda_{g}=\int_{U}g^{I_{s}J_{s}}\,h_{bc}\,G^{b}_{I_{s}}\,\Psi^{c}_{J_{s}}\,d\lambda_{g}.$
Fix a target block $T_{s}=(t_{1},\dots,t_{s})$ and a fixed fiber index $a$. Given $\psi\in C_{c}^{\infty}(\phi(U))$, we define a convenient test tensor field $\Psi$ by its components:
$\Psi^{c}_{J_{s}}=\frac{1}{\sqrt{\det(g)}}\,h^{ca}\,g_{J_{s}T_{s}}\,(\psi\circ\phi).$ (17)
By Corollary 1, the formal adjoint acts on $\Psi$ as
$\big((\nabla^{s})^{*}\Psi\big)^{c}=(-1)^{s}\left[\frac{1}{\sqrt{\det(g)}}\partial_{I_{s}}\left(\sqrt{\det(g)}g^{I_{s}J_{s}}\Psi^{c}_{J_{s}}\right)+\sum_{|\beta|\leq s-1}\big(D_{\beta}\big)^{c\,J_{s}}_{d}\partial^{\beta}\Psi^{d}_{J_{s}}\right].$
Substituting (17) into this expression and using the facts that $\frac{\sqrt{\det(g)}}{\sqrt{\det(g)}}=1$ and $g^{I_{s}J_{s}}g_{J_{s}T_{s}}=\delta^{I_{s}}_{T_{s}}$, we obtain
$\big((\nabla^{s})^{*}\Psi\big)^{c}=(-1)^{s}\left[\frac{1}{\sqrt{\det(g)}}\partial_{T_{s}}\left(h^{ca}(\psi\circ\phi)\right)+\sum_{|\beta|\leq s-1}\big(D_{\beta}\big)^{c\,J_{s}}_{d}\partial^{\beta}\left(\frac{h^{da}g_{J_{s}T_{s}}}{\sqrt{\det(g)}}(\psi\circ\phi)\right)\right].$
Let $I_{1}:=\int_{U}\langle F,(\nabla^{s})^{*}\Psi\rangle_{g,h_{E}}\,d\lambda_{g}$ and $I_{2}:=\int_{U}\langle G,\Psi\rangle_{g,h_{E}}\,d\lambda_{g}$. Substituting (17) into $I_{2}$ and noting that $g^{I_{s}J_{s}}g_{J_{s}T_{s}}=\delta^{I_{s}}_{T_{s}}$, we obtain
$I_{2}=\int_{U}\frac{1}{\sqrt{\det(g)}}\delta^{I_{s}}_{T_{s}}\delta_{b}^{a}G^{b}_{I_{s}}(\psi\circ\phi)\,d\lambda_{g}=\int_{\phi(U)}(G^{a}_{T_{s}}\circ\phi^{-1})\psi\,d\lambda_{n}.$ (18)
For $I_{1}$, we use the Leibniz formula on $\partial_{T_{s}}$:
$\partial_{T_{s}}\left(h^{ca}(\psi\circ\phi)\right)=h^{ca}\partial_{T_{s}}(\psi\circ\phi)+\sum_{\alpha<\beta_{T_{s}}}\binom{\beta_{T_{s}}}{\alpha}\partial^{\alpha}(\psi\circ\phi)\partial^{\beta_{T_{s}}-\alpha}\left(h^{ca}\right).$
Substituting back into the integral $I_{1}$, we obtain
$I_{1}$ $=(-1)^{s}\int_{U}\frac{1}{\sqrt{\det(g)}}F^{a}\partial_{T_{s}}(\psi\circ\phi)\,d\lambda_{g}$
$\quad+(-1)^{s}\sum_{\alpha<\beta_{T_{s}}}\int_{U}\frac{1}{\sqrt{\det(g)}}h_{bc}F^{b}\binom{\beta_{T_{s}}}{\alpha}\partial^{\alpha}(\psi\circ\phi)\partial^{\beta_{T_{s}}-\alpha}\left(h^{ca}\right)\,d\lambda_{g}$
$\quad+\sum_{|\beta|\leq s-1}(-1)^{s}\int_{U}h_{bc}F^{b}\big(D_{\beta}\big)^{c\,J_{s}}_{d}\partial^{\beta}\left(\frac{h^{da}g_{J_{s}T_{s}}}{\sqrt{\det(g)}}(\psi\circ\phi)\right)\,d\lambda_{g}.$
We now proceed by induction on $s$. For $s=1$, we have $T_{1}=(t_{1})$, $J_{1}=(j_{1})$, so the expression reduces to
$I_{1}$ $=-\int_{\phi(U)}(F^{a}\circ\phi^{-1})\partial_{t_{1}}\psi\,d\lambda_{n}-\int_{\phi(U)}((h_{bc}F^{b})\circ\phi^{-1})\partial_{t_{1}}\left(h^{ca}\circ\phi^{-1}\right)\psi\,d\lambda_{g}$
$\quad-\int_{\phi(U)}\left(\left(h_{bc}F^{b}(D_{0})^{c\,j_{1}}_{d}h^{da}g_{j_{1}t_{1}}\right)\circ\phi^{-1}\right)\psi\,d\lambda_{g}.$
Define
$H_{t_{1}}^{a}:=h_{bc}F^{b}\partial_{t_{1}}\left(h^{ca}\right)+h_{bc}F^{b}(D_{0})^{c\,j_{1}}_{d}h^{da}g_{j_{1}t_{1}}.$
Since $F|_{U}\in W^{m,p}(E|_{U})$ implies that $F^{b}\circ\phi^{-1}\in L^{p}(\phi(U))$, and the metric coefficients are smooth on the compact set $\overline{\phi(U)}$, it follows that $H_{t_{1}}^{a}\circ\phi^{-1}\in L^{p}(\phi(U))$. Since $I_{1}=I_{2}$, we have
$\int_{\phi(U)}(F^{a}\circ\phi^{-1})\partial_{t_{1}}\psi\,d\lambda_{n}=\int_{\phi(U)}\big((G_{t_{1}}^{a}\circ\phi^{-1})+(H^{a}_{t_{1}}\circ\phi^{-1})\big)\psi\,d\lambda_{n},$
which shows that $F^{a}\circ\phi^{-1}\in W^{1,p}(\phi(U))$.
Now assume the statement holds up to order s1s-1 for $s\leq m$. The inductive hypothesis guarantees that the components of $F$ in $U$ have weak derivatives of order $\leq s-1$, so we can integrate by parts in the weak sense for the lower-order terms.
First, consider the terms coming from the Leibniz expansion in $I_{1}$:
$(-1)^{s}$ $\sum_{\alpha<\beta_{T_{s}}}\int_{U}\frac{h_{bc}F^{b}}{\sqrt{\det(g)}}\binom{\beta_{T_{s}}}{\alpha}\partial^{\alpha}(\psi\circ\phi)\partial^{\beta_{T_{s}}-\alpha}\left(h^{ca}\right)\,d\lambda_{g}$
$=(-1)^{s}\sum_{\alpha<\beta_{T_{s}}}\int_{\phi(U)}\left[\left(h_{bc}F^{b}\binom{\beta_{T_{s}}}{\alpha}\partial^{\beta_{T_{s}}-\alpha}\left(h^{ca}\right)\right)\circ\phi^{-1}\right]\partial^{\alpha}\psi\,d\lambda_{n}.$
Since $|\alpha|<|\beta_{I_{s}}|=s$, we have $|\alpha|\leq s-1$. Integrating by parts (weakly) $|\alpha|$ times, we transfer the derivatives from $\psi$ to the coefficient (which involves $F^{b}$), so the previous term equals
$\sum_{\alpha<\beta_{T_{s}}}(-1)^{s+|\alpha|}\int_{\phi(U)}\partial^{\alpha}\left[\left(h_{bc}F^{b}\binom{\beta_{T_{s}}}{\alpha}\partial^{\beta_{T_{s}}-\alpha}\left(h^{ca}\right)\right)\circ\phi^{-1}\right]\psi\,d\lambda_{n}.$
Second, consider the terms in $I_{1}$ involving the connection coefficients ($D_{\beta}$). Using the same change of variables and recalling that $|\beta|\leq s-1$, we get
$\sum_{|\beta|\leq s-1}(-1)^{s}\int_{U}h_{bc}F^{b}\big(D_{\beta}\big)^{c\,J_{s}}_{d}\partial^{\beta}\left(\frac{h^{da}g_{J_{s}T_{s}}}{\sqrt{\det(g)}}(\psi\circ\phi)\right)\,d\lambda_{g}$
$=(-1)^{s}\sum_{|\beta|\leq s-1}\int_{\phi(U)}\left[\left(h_{bc}F^{b}\big(D_{\beta}\big)^{c\,J_{s}}_{d}\sqrt{\det(g)}\right)\circ\phi^{-1}\right]\partial^{\beta}\left(\left(\frac{h^{da}g_{J_{s}T_{s}}}{\sqrt{\det(g)}}\circ\phi^{-1}\right)\psi\right)\,d\lambda_{n}.$
Integrating by parts $|\beta|$ times yields
$(-1)^{s}\sum_{|\beta|\leq s-1}(-1)^{|\beta|}\int_{\phi(U)}\partial^{\beta}\left[\left(h_{bc}F^{b}\big(D_{\beta}\big)^{c\,J_{s}}_{d}\sqrt{\det(g)}\right)\circ\phi^{-1}\right]\left(\frac{h^{da}g_{J_{s}T_{s}}}{\sqrt{\det(g)}}\circ\phi^{-1}\right)\psi\,d\lambda_{n}.$
We now define $\widetilde{G}^{a}_{T_{s}}$ on $\phi(U)$ by grouping $G^{a}_{T_{s}}\circ\phi^{-1}$ (from $I_{2}$) and moving the terms calculated above from $I_{1}$ to the other side (introducing an extra factor of 1-1, which becomes $(-1)^{s+1}$):
$\widetilde{G}^{a}_{T_{s}}$ $:=G^{a}_{T_{s}}\circ\phi^{-1}+\sum_{\alpha<\beta_{T_{s}}}(-1)^{s+1+|\alpha|}\partial^{\alpha}\left[\left(h_{bc}F^{b}\binom{\beta_{T_{s}}}{\alpha}\partial^{\beta_{T_{s}}-\alpha}\left(h^{ca}\right)\right)\circ\phi^{-1}\right]$
$\quad+\sum_{|\beta|\leq s-1}(-1)^{s+1+|\beta|}\partial^{\beta}\left[\left(h_{bc}F^{b}\big(D_{\beta}\big)^{c\,J_{s}}_{d}\sqrt{\det(g)}\right)\circ\phi^{-1}\right]\left(\frac{h^{da}g_{J_{s}T_{s}}}{\sqrt{\det(g)}}\circ\phi^{-1}\right).$
The inductive hypothesis states that $F^{b}\circ\phi^{-1}\in W^{s-1,p}(\phi(U))$ for all $b\in\{1,\dots,r\}$. Since $\widetilde{G}_{T_{s}}^{a}$ consists of smooth bounded functions (defined on $\overline{\phi(U)}$) and weak derivatives of $F^{b}\circ\phi^{-1}$ of order at most s1s-1, we conclude that $\widetilde{G}_{T_{s}}^{a}\in L^{p}(\phi(U))$. Using the equality $I_{1}=I_{2}$ and the definition of $\widetilde{G}$, we obtain
$(-1)^{s}\int_{\phi(U)}(F^{a}\circ\phi^{-1})\partial_{T_{s}}\psi\,d\lambda_{n}=\int_{\phi(U)}\widetilde{G}_{T_{s}}^{a}\psi\,d\lambda_{n}.$
This shows that $F^{a}\circ\phi^{-1}$ possesses a weak derivative $\partial_{T_{s}}(F^{a}\circ\phi^{-1})=\widetilde{G}_{T_{s}}^{a}\in L^{p}(\phi(U))$. Thus, $F^{a}\circ\phi^{-1}\in W^{s,p}(\phi(U))$. Since $s\leq m$ was arbitrary, we conclude that $F^{a}\circ\phi^{-1}\in W^{m,p}(\phi(U))$ for all $a\in\{1,\dots,r\}$.
($\Leftarrow$) Now assume that $F^{a}\circ\phi^{-1}\in W^{m,p}(\phi(U))$ for all $a\in\{1,\dots,r\}$. Fix $s\leq m$ and let $\Psi\in\Gamma_{c}\big(T^{(0,s)}(TU)\otimes E|_{U}\big)$. We have
$\int_{U}\langle F,(\nabla^{s})^{*}\Psi\rangle_{g,h_{E}}\,d\lambda_{g}=\int_{U}h_{ab}\,F^{a}\big((\nabla^{s})^{*}\Psi\big)^{b}\,d\lambda_{g}.$
By Corollary 1, the adjoint in coordinates takes the form
$\big((\nabla^{s})^{*}\Psi\big)^{b}=(-1)^{s}\left[\frac{1}{\sqrt{\det(g)}}\partial_{I_{s}}\left(\sqrt{\det(g)}g^{I_{s}J_{s}}\Psi^{b}_{J_{s}}\right)+\sum_{|\beta|\leq s-1}\big(D_{\beta}\big)^{b\,J_{s}}_{c}\,\partial^{\beta}\Psi^{c}_{J_{s}}\right].$
Then the integral splits into two terms, $I_{1}$ and $I_{2}$:
$\int_{U}\langle F,(\nabla^{s})^{*}\Psi\rangle_{g,h_{E}}\,d\lambda_{g}$ $=\underbrace{(-1)^{s}\int_{U}h_{ab}F^{a}\frac{1}{\sqrt{\det(g)}}\partial_{I_{s}}\left(\sqrt{\det(g)}g^{I_{s}J_{s}}\Psi^{b}_{J_{s}}\right)\,d\lambda_{g}}_{I_{1}}$
$\quad+\underbrace{(-1)^{s}\sum_{|\beta|\leq s-1}\int_{U}h_{ab}F^{a}\big(D_{\beta}\big)^{b\,J_{s}}_{c}\,\partial^{\beta}\Psi^{c}_{J_{s}}\,d\lambda_{g}}_{I_{2}}.$
Changing the domain of integration to $\phi(U)$ and writing the compositions explicitly:
$I_{1}$ $=(-1)^{s}\int_{\phi(U)}\left[\left(h_{ab}F^{a}\right)\circ\phi^{-1}\right]\partial_{I_{s}}\left(\left(\sqrt{\det(g)}g^{I_{s}J_{s}}\Psi^{b}_{J_{s}}\right)\circ\phi^{-1}\right)\,d\lambda_{n},$
$I_{2}$ $=(-1)^{s}\sum_{|\beta|\leq s-1}\int_{\phi(U)}\left[\left(h_{ab}F^{a}(D_{\beta})^{b\,J_{s}}_{c}\sqrt{\det(g)}\right)\circ\phi^{-1}\right]\partial^{\beta}(\Psi^{c}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}.$
We now integrate by parts weakly in $I_{1}$ (valid since $F^{a}\circ\phi^{-1}\in W^{m,p}(\phi(U))$ and $s\leq m$):
$I_{1}=\int_{\phi(U)}\partial_{I_{s}}\left[\left(h_{ab}\circ\phi^{-1}\right)(F^{a}\circ\phi^{-1})\right]\left(\sqrt{\det(g)}g^{I_{s}J_{s}}\Psi^{b}_{J_{s}}\circ\phi^{-1}\right)\,d\lambda_{n}.$
Applying the Leibniz rule to the bracketed term:
$I_{1}$ $=\int_{\phi(U)}\Bigl[\bigl(h_{ab}g^{I_{s}J_{s}}\sqrt{\det(g)}\circ\phi^{-1}\bigr)\partial_{I_{s}}(F^{a}\circ\phi^{-1})\Bigr](\Psi^{b}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}$
$\quad+\int_{\phi(U)}\Bigg[\sum_{\begin{subarray}{c}I_{s}\\ \alpha<\beta_{I_{s}}\end{subarray}}\binom{\beta_{I_{s}}}{\alpha}\partial^{\beta_{I_{s}}-\alpha}\bigl(h_{ab}\circ\phi^{-1}\bigr)\partial^{\alpha}(F^{a}\circ\phi^{-1})$
$\qquad\qquad\qquad\times\left(\bigl(g^{I_{s}J_{s}}\sqrt{\det(g)}\Psi^{b}_{J_{s}}\bigr)\circ\phi^{-1}\right)\Bigg]\,d\lambda_{n}.$
Define the coefficient functions $(B_{\gamma})_{abI_{s}}$ on $\phi(U)$ for $|\gamma|\leq s-1$ as
$(B_{\gamma})_{ab}^{J_{s}}:=\sum_{\begin{subarray}{c}I_{s}\\ \gamma<\beta_{I_{s}}\end{subarray}}\binom{\beta_{I_{s}}}{\gamma}\partial^{\beta_{I_{s}}-\gamma}\bigl(h_{ab}\circ\phi^{-1}\bigr)\left(\bigl(g^{I_{s}J_{s}}\sqrt{\det(g)}\bigr)\circ\phi^{-1}\right).$
Substituting this definition into $I_{1}$:
$I_{1}$ $=\int_{\phi(U)}\Bigg[\bigl(h_{ab}g^{I_{s}J_{s}}\sqrt{\det(g)}\circ\phi^{-1}\bigr)\partial_{I_{s}}(F^{a}\circ\phi^{-1})$
$\qquad\qquad+\sum_{|\gamma|\leq s-1}(B_{\gamma})_{ab}^{J_{s}}\partial^{\gamma}(F^{a}\circ\phi^{-1})\Bigg](\Psi^{b}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}.$
For $I_{2}$, integrating by parts $|\beta|$ times yields
$I_{2}$ $=\sum_{|\beta|\leq s-1}(-1)^{s+|\beta|}\int_{\phi(U)}\partial^{\beta}\Bigl[\Bigl(\bigl(h_{ab}(D_{\beta})^{b\,J_{s}}_{c}\sqrt{\det(g)}\bigr)\circ\phi^{-1}\Bigr)$
$\qquad\qquad\qquad\qquad\qquad\times(F^{a}\circ\phi^{-1})\Bigr](\Psi^{c}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}.$
Applying Leibniz rule again to expand the derivative $\partial^{\beta}$ of the product:
$I_{2}$ $=\int_{\phi(U)}\sum_{|\beta|\leq s-1}(-1)^{s+|\beta|}\sum_{\gamma\leq\beta}\binom{\beta}{\gamma}\partial^{\beta-\gamma}\Bigl(\bigl(h_{ab}(D_{\beta})^{b\,J_{s}}_{c}\sqrt{\det(g)}\bigr)\circ\phi^{-1}\Bigr)$
$\qquad\qquad\qquad\times\partial^{\gamma}(F^{a}\circ\phi^{-1})(\Psi^{c}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}.$
Define the coefficient functions $(C_{\gamma})^{J_{s}}_{ac}$ on $\phi(U)$ for each $|\gamma|\leq s-1$ as
$(C_{\gamma})^{J_{s}}_{ac}:=\sum_{\begin{subarray}{c}|\beta|\leq s-1\\ \gamma\leq\beta\end{subarray}}(-1)^{s+|\beta|}\binom{\beta}{\gamma}\partial^{\beta-\gamma}\left(\left(h_{ab}(D_{\beta})^{b\,J_{s}}_{c}\sqrt{\det(g)}\right)\circ\phi^{-1}\right).$
Substituting this definition into $I_{2}$:
$I_{2}=\int_{\phi(U)}\sum_{|\gamma|\leq s-1}(C_{\gamma})^{J_{s}}_{ac}\partial^{\gamma}(F^{a}\circ\phi^{-1})(\Psi^{c}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}.$
Combining both results, and renaming the fiber index $b\to c$ in $I_{1}$ (using $h_{ac}$):
$\int_{U}\langle F,(\nabla^{s})^{*}\Psi\rangle\,d\lambda_{g}$ $=\int_{\phi(U)}\Bigg[\left(h_{ac}g^{I_{s}J_{s}}\sqrt{\det(g)}\circ\phi^{-1}\right)\partial_{I_{s}}(F^{a}\circ\phi^{-1})$
$\quad+\sum_{|\gamma|\leq s-1}\left((B_{\gamma})_{ac}^{J_{s}}+(C_{\gamma})^{J_{s}}_{ac}\right)\partial^{\gamma}(F^{a}\circ\phi^{-1})\Bigg](\Psi^{c}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}.$
We now define the correction coefficients $(\widehat{D}_{\gamma})$ on $U$ by pulling back $B$ and $C$ and inverting the metric weights:
$\big(\widehat{D}_{\gamma}\big)^{a}_{b\,I_{s}}:=\frac{g_{I_{s}\tilde{J}_{s}}h^{a\tilde{c}}}{\sqrt{\det(g)}}\Big[\big((B_{\gamma})^{\tilde{J}_{s}}_{b\tilde{c}}\circ\phi\big)+\big((C_{\gamma})^{\tilde{J}_{s}}_{b\tilde{c}}\circ\phi\big)\Big].$
Let $G\in L^{p}(T^{(0,s)}(TM)\otimes E|_{U})$ be given in coordinates by:
$G^{a}_{I_{s}}=\partial_{I_{s}}F^{a}+\sum_{|\gamma|\leq s-1}\big(\widehat{D}_{\gamma}\big)^{a}_{b\,I_{s}}\partial^{\gamma}F^{b}.$ (19)
We verify that $G$ is the required weak derivative:
$\int_{U}\langle G,\Psi\rangle_{g,h_{E}}\,d\lambda_{g}$ $=\int_{U}h_{ac}g^{I_{s}J_{s}}G_{I_{s}}^{a}\Psi_{J_{s}}^{c}\,d\lambda_{g}$
$=\int_{U}h_{ac}g^{I_{s}J_{s}}\left(\partial_{I_{s}}F^{a}+\sum_{|\gamma|\leq s-1}(\widehat{D}_{\gamma})^{a}_{b\,I_{s}}\partial^{\gamma}F^{b}\right)\Psi^{c}_{J_{s}}\,d\lambda_{g}.$
The first term matches the principal part. For the lower-order terms, substituting $\widehat{D}_{\gamma}$:
$\int_{U}\sum_{|\gamma|\leq s-1}h_{ac}g^{I_{s}J_{s}}\left(\frac{g_{I_{s}\tilde{J}_{s}}h^{a\tilde{c}}}{\sqrt{\det(g)}}\Big[\big((B_{\gamma})^{\tilde{J}_{s}}_{b\tilde{c}}\circ\phi\big)+\big((C_{\gamma})^{\tilde{J}_{s}}_{b\tilde{c}}\circ\phi\big)\Big]\right)\partial^{\gamma}F^{b}\Psi^{c}_{J_{s}}\,d\lambda_{g}$
$=\int_{U}\sum_{|\gamma|\leq s-1}\frac{1}{\sqrt{\det(g)}}(h_{ac}h^{a\tilde{c}})(g^{I_{s}J_{s}}g_{I_{s}\tilde{J}_{s}})\Big[\big((B_{\gamma})^{\tilde{J}_{s}}_{b\tilde{c}}\circ\phi\big)+\big((C_{\gamma})^{\tilde{J}_{s}}_{b\tilde{c}}\circ\phi\big)\Big]\partial^{\gamma}F^{b}\Psi^{c}_{J_{s}}\,d\lambda_{g}.$
Using the identities $h_{ac}h^{a\tilde{c}}=\delta^{\tilde{c}}_{c}$ and $g^{I_{s}J_{s}}g_{I_{s}\tilde{J}_{s}}=\delta^{J_{s}}_{\tilde{J}_{s}}$, we obtain
$=\int_{U}\sum_{|\gamma|\leq s-1}\frac{1}{\sqrt{\det(g)}}\Big[\big((B_{\gamma})^{J_{s}}_{bc}\circ\phi\big)+\big((C_{\gamma})^{J_{s}}_{bc}\circ\phi\big)\Big]\partial^{\gamma}F^{b}\Psi^{c}_{J_{s}}\,d\lambda_{g}$
$=\int_{\phi(U)}\sum_{|\gamma|\leq s-1}\left[(B_{\gamma})^{J_{s}}_{bc}+(C_{\gamma})^{J_{s}}_{bc}\right]\partial^{\gamma}(F^{b}\circ\phi^{-1})(\Psi^{c}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}.$
This expression matches the expansion of $\int_{U}\langle F,(\nabla^{s})^{*}\Psi\rangle\,d\lambda_{g}$ exactly. Therefore, $G$ is the $s$-th weak covariant derivative of $F$ in $U$. Moreover, $G\in L^{p}(T^{(0,s)}(TU)\otimes E|_{U})$ because each coordinate component is a linear combination of weak derivatives of the components of $F$ (up to order $s$) multiplied by smooth bounded functions. Since $s\leq m$ was arbitrary, $F|_{U}\in W^{m,p}\big(E|_{U}\big)$.
Finally, we establish the norm estimates. The Sobolev norm is defined as
$\|F\|_{W^{m,p}(E|_{U})}:=\left(\sum_{s=0}^{m}\int_{U}|\nabla^{s}_{w}F|_{g,h_{E}}^{p}\,d\lambda_{g}.\right)^{\frac{1}{p}}$
From (19), for any $s\leq m$ and multi-index $I_{s}=(i_{1},\dots,i_{s})$:
$(\nabla^{s}_{w}F)^{a}_{I_{s}}=\partial_{I_{s}}F^{a}+\sum_{|\gamma|\leq s-1}\big(\widehat{D}_{\gamma}\big)^{a}_{b\,I_{s}}\,\partial^{\gamma}F^{b}.$
Define a fiber metric $\overline{h}_{E}$ on $E|_{U}$ by $\overline{h}_{E}(u,v)=\displaystyle\sum_{a=1}^{r}u^{a}v^{a}$ for sections $u=\displaystyle\sum_{a=1}^{r}u^{a}e_{a},v=\sum_{a=1}^{r}v^{a}e_{a}$. Also consider the Riemannian metric $\delta:=\phi^{-1}\overline{g}$ on $\overline{U}$. Since $\overline{U}$ is compact, Lemma 2 provides constants $c_{0},C_{0}>0$ for the pointwise equivalence of the fiber norms, uniform for all $s\in\{0,\dots,m\}$. By integrating these estimates and observing that the Riemann–Lebesgue measures $d\lambda_{g}$ and $d\lambda_{\overline{g}}$ are comparable on compact sets through positive constants [9, Lemma 2], we conclude that there exist constants $c,C>0$ such that the following equivalence of Sobolev norms holds:
$c\|F\|_{W^{m,p}(E|_{U}),\overline{g},\overline{h}_{E}}\leq\|F\|_{W^{m,p}(E|_{U})}\leq C\|F\|_{W^{m,p}(E|_{U}),\overline{g},\overline{h}_{E}}.$ (20)
We express the auxiliary norm as:
$\|F\|_{W^{m,p}(E|_{U}),\overline{g},\overline{h}_{E}}$ $=\Bigg(\sum_{s=0}^{m}\int_{U}\Bigg(\sum_{a=1}^{r}\sum_{I_{s}}\bigg(\partial_{I_{s}}F^{a}+\sum_{|\gamma|\leq s-1}\big(\widehat{D}_{\gamma}\big)^{a}_{b\,I_{s}}\partial^{\gamma}F^{b}\bigg)^{2}\Bigg)^{\frac{p}{2}}d\lambda_{\overline{g}}\Bigg)^{\frac{1}{p}}$
$=\Bigg(\sum_{s=0}^{m}\int_{\phi(U)}\Bigg(\sum_{a=1}^{r}\sum_{I_{s}}\bigg(\partial_{I_{s}}(F^{a}\circ\phi^{-1})$
$\qquad\qquad+\sum_{|\gamma|\leq s-1}\left(\big(\widehat{D}_{\gamma}\big)^{a}_{b\,I_{s}}\circ\phi^{-1}\right)\partial^{\gamma}(F^{b}\circ\phi^{-1})\bigg)^{2}\Bigg)^{\frac{p}{2}}d\lambda_{n}\Bigg)^{\frac{1}{p}}.$
To bound this from above, using $(A+B)^{2}\leq 2A^{2}+2B^{2}$ and defining $K:=\max\|\widehat{D}\|_{\infty}$, we obtain:
$\|F\|_{W^{m,p}(E|_{U}),\overline{g},\overline{h}_{E}}$ $\leq\sqrt{2}\Bigg(\sum_{s=0}^{m}\int_{\phi(U)}\Bigg(\sum_{a=1}^{r}\sum_{|\alpha|=s}\frac{s!}{\alpha!}\left(\partial^{\alpha}(F^{a}\circ\phi^{-1})\right)^{2}$
$\quad+rn^{s}\left(\sum_{b=1}^{r}\sum_{|\gamma|\leq s-1}K\partial^{\gamma}(F^{b}\circ\phi^{-1})\right)^{2}\Bigg)^{\frac{p}{2}}d\lambda_{n}\Bigg)^{\frac{1}{p}}.$
Using $\|\cdot\|_{2}\leq\|\cdot\|_{1}$ to introduce square roots on coefficients, and then applying $(\sum|x_{i}|)^{p}\leq N^{p-1}\sum|x_{i}|^{p}$, all multiplicative constants absorb into a constant $C^{\prime}>0$:
$\leq C^{\prime}\left(\sum_{s=0}^{m}\int_{\phi(U)}\sum_{a=1}^{r}\left(\sum_{|\alpha|=s}\left|\partial^{\alpha}(F^{a}\circ\phi^{-1})\right|^{p}+\sum_{|\gamma|\leq s-1}\left|\partial^{\gamma}(F^{a}\circ\phi^{-1})\right|^{p}\right)d\lambda_{n}\right)^{\frac{1}{p}}$
$\leq C^{\prime\prime}\left(\sum_{a=1}^{r}\sum_{|\beta|\leq m}\int_{\phi(U)}\left|\partial^{\beta}(F^{a}\circ\phi^{-1})\right|^{p}d\lambda_{n}\right)^{\frac{1}{p}}=C^{\prime\prime}\left(\sum_{a=1}^{r}\left\|F^{a}\circ\phi^{-1}\right\|_{W^{m,p}(\phi(U))}^{p}\right)^{\frac{1}{p}}.$
Using $\|\cdot\|_{p}\leq\|\cdot\|_{1}$, we obtain the upper bound:
$\|F\|_{W^{m,p}(E|_{U})}\leq C_{m}\sum_{a=1}^{r}\left\|F^{a}\circ\phi^{-1}\right\|_{W^{m,p}(\phi(U))}.$
For the lower bound, from (19), we isolate the partial derivative:
$\partial_{I_{s}}(F^{a}\circ\phi^{-1})=(\nabla^{s}_{w}F)^{a}_{I_{s}}\circ\phi^{-1}-\sum_{|\gamma|\leq s-1}\left(\big(\widehat{D}_{\gamma}\big)^{a}_{b\,I_{s}}\circ\phi^{-1}\right)\partial^{\gamma}(F^{b}\circ\phi^{-1}).$ (21)
The $L^{p}$ norm of the first term is bounded by the bundle Sobolev norm. Taking $L^{p}(\phi(U))$ norms and summing over components and orders:
$\sum_{a=1}^{r}\|F^{a}\circ\phi^{-1}\|_{W^{m,p}(\phi(U))}\leq\widetilde{C}\left(\|F\|_{W^{m,p}(E|_{U}),\overline{g},\overline{h}_{E}}+\sum_{a=1}^{r}\|F^{a}\circ\phi^{-1}\|_{W^{m-1,p}(\phi(U))}\right).$ (22)
We proceed by induction on $m$. For $m=1$, the $L^{p}$ norms are clearly bounded. Assume the bound holds for m1m-1. Then
$\sum_{a=1}^{r}\|F^{a}\circ\phi^{-1}\|_{W^{m-1,p}(\phi(U))}\leq K_{m-1}\|F\|_{W^{m-1,p}(E|_{U}),\overline{g},\overline{h}_{E}}.$
Substituting this into (22):
$\sum_{a=1}^{r}\|F^{a}\circ\phi^{-1}\|_{W^{m,p}(\phi(U))}\leq\widetilde{C}\left(1+K_{m-1}\right)\|F\|_{W^{m,p}(E|_{U}),\overline{g},\overline{h}_{E}}.$
Combining this with (20), we obtain the lower bound constant $c_{m}>0$. ∎
Since the local characterization in Lemma 3 relies on the Sobolev spaces defined on the restricted bundle $E|_{U}$, we first ensure that the restriction of a global Sobolev section is well-defined and consistent with the local theory:
Proposition 1 (Restriction to open subsets).
Let $(M,g)$ be a Riemannian manifold and $E\to M$ a smooth vector bundle equipped with a fiber metric $h_{E}$. If $U\subseteq M$ is an open subset and $u\in W^{m,p}(E)$, then $u|_{U}\in W^{m,p}(E|_{U})$.
Proof.
Let $s\leq m$. To show that $u|_{U}$ possesses a weak covariant derivative of order $s$ in $L^{p}(T^{(0,s)}(TU)\otimes E|_{U})$, let $\sigma\in\Gamma_{c}(T^{(0,s)}(TU)\otimes E|_{U})$ be an arbitrary test section. We extend $\sigma$ to a section $\widetilde{\sigma}\in\Gamma_{c}(T^{(0,s)}(TM)\otimes E)$ by setting it to zero outside $U$.
Since $u\in W^{m,p}(E)$, it possesses a global $s$-th weak covariant derivative $\nabla_{w}^{s}u$ satisfying:
$\int_{M}\langle\nabla_{w}^{s}u,\widetilde{\sigma}\rangle_{g,h_{E}}d\lambda_{g}=\int_{M}\langle u,(\nabla^{s})^{*}\widetilde{\sigma}\rangle_{h_{E}}d\lambda_{g}.$
Because $(\nabla^{s})^{*}$ is a local operator, the support of $(\nabla^{s})^{*}\widetilde{\sigma}$ is contained in $\operatorname{supp}(\widetilde{\sigma})=\operatorname{supp}(\sigma)\subseteq U$. Furthermore, on the open set $U$, we have $(\nabla^{s})^{*}\widetilde{\sigma}\equiv(\nabla^{s})^{*}\sigma$. It follows that:
$\int_{U}\langle(\nabla_{w}^{s}u)|_{U},\sigma\rangle_{g,h_{E}}d\lambda_{g}$ $=\int_{M}\langle\nabla_{w}^{s}u,\widetilde{\sigma}\rangle_{g,h_{E}}d\lambda_{g}$
$=\int_{M}\langle u,(\nabla^{s})^{*}\widetilde{\sigma}\rangle_{h_{E}}d\lambda_{g}$
$=\int_{U}\langle u|_{U},(\nabla^{s})^{*}\sigma\rangle_{h_{E}}d\lambda_{g}.$
This identity identifies $(\nabla_{w}^{s}u)|_{U}$ as the $s$-th weak covariant derivative of $u|_{U}$. Since $\nabla_{w}^{s}u\in L^{p}$, its restriction remains in $L^{p}(T^{(0,s)}(TU)\otimes E|_{U})$. As $s\leq m$ was arbitrary, we conclude that $u|_{U}\in W^{m,p}(E|_{U})$. ∎
Another fundamental property is that the weak covariant derivative is a local operator:
Proposition 2 (Locality of the Weak Covariant Derivative).
Let $(M,g)$ be a Riemannian manifold and let $E\to M$ be a smooth vector bundle with fiber metric $h_{E}$ and a connection $\nabla^{E}$. Let $u\in L^{1}_{\operatorname{loc}}(E)$ be a section possessing a weak $s$-th covariant derivative $\nabla^{s}_{w}u$. Let $U\subseteq M$ be an open set. If $u=0$ a.e. in $U$, then $\nabla^{s}_{w}u=0$ a.e. in $U$.
Proof.
Let $\sigma\in\Gamma_{c}(T^{(0,s)}(TU)\otimes E|_{U})$ be an arbitrary test section with support in $U$. We extend $\sigma$ by zero to the whole manifold $M$, denoting the extension by $\widetilde{\sigma}$. Since $\operatorname{supp}(\sigma)\subset U$, $\widetilde{\sigma}$ is smooth and compactly supported on $M$. By the definition of the weak derivative,
$\int_{M}\langle\nabla_{w}^{s}u,\widetilde{\sigma}\rangle_{g,h_{E}}d\lambda_{g}=\int_{M}\langle u,(\nabla^{s})^{*}\widetilde{\sigma}\rangle_{h_{E}}d\lambda_{g}.$
Since $(\nabla^{s})^{*}$ is a local operator (being a differential operator with smooth coefficients), $\operatorname{supp}((\nabla^{s})^{*}\widetilde{\sigma})\subseteq\operatorname{supp}(\widetilde{\sigma})\subset U$. Since $u$ vanishes almost everywhere on $U$, the right-hand side is zero. Thus,
$\int_{U}\langle\nabla_{w}^{s}u,\sigma\rangle_{g,h_{E}}d\lambda_{g}=0$
for all $\sigma\in\Gamma_{c}(T^{(0,s)}(TU)\otimes E|_{U})$. By the fundamental lemma of the calculus of variations extended to vector bundles (density of test sections in $L^{1}_{loc}$), we conclude that $\nabla_{w}^{s}u=0$ almost everywhere in $U$. ∎
5.  Global Approximation Results
We are now in a position to prove the main approximation result of this paper. The following theorem generalizes the classical Meyers-Serrin theorem to the context of vector bundles over Riemannian manifolds, stating that smooth sections are dense in the Sobolev space $W^{m,p}(E)$.
Theorem 2 (Meyers–Serrin for Vector Bundles).
Let $(M,g)$ be a Riemannian manifold and let $\pi_{E}:E\to M$ be a smooth vector bundle equipped with a fiber metric $h_{E}$ and a compatible connection $\nabla^{E}$. Then
$H^{m,p}(E)=W^{m,p}(E).$
Proof.
Since $M$ is a Lindelöf space, it admits a countable open cover by regular coordinate balls $(U_{\alpha},\phi_{\alpha})_{\alpha\in\mathbb{N}}$. Let $(\psi_{\alpha})_{\alpha\in\mathbb{N}}$ be a smooth partition of unity subordinate to this cover.
Consider a section $F\in W^{m,p}(E)$. We can decompose it as
$F=\sum_{\alpha=1}^{\infty}\psi_{\alpha}F,$
where the sum is locally finite. For each $\alpha$, by Proposition 1, $\psi_{\alpha}F|_{U_{\alpha}}$ belongs to $W^{m,p}(E|_{U_{\alpha}})$, since $\psi_{\alpha}$ is smooth and has compact support contained in $U_{\alpha}$.
Fix $\alpha$ and work in the chart $(U_{\alpha},\phi_{\alpha})$, equipped with a local frame $(e_{1},\dots,e_{r})$ of $E$ over $U_{\alpha}$. Applying Lemma 3 to the section $(\psi_{\alpha}F)|_{U_{\alpha}}$, we deduce that for every fiber index $a\in\{1,\dots,r\}$,
$\big(\psi_{\alpha}F\big)^{a}\circ\phi_{\alpha}^{-1}\in W^{m,p}(\phi_{\alpha}(U_{\alpha})).$
Moreover, there exist constants $C_{m,\alpha}>0$ such that, for any section $H\in W^{m,p}(E|_{U_{\alpha}})$,
$\|H\|_{W^{m,p}(E|_{U_{\alpha}})}\leq C_{m,\alpha}\sum_{a=1}^{r}\big\|H^{a}\circ\phi_{\alpha}^{-1}\big\|_{W^{m,p}(\phi_{\alpha}(U_{\alpha}))}.$
Let $\Omega_{\alpha}:=\phi_{\alpha}(U_{\alpha})\subset\mathbb{R}^{n}$. By the classical Meyers–Serrin theorem on domains in $\mathbb{R}^{n}$, for each $\alpha\in\mathbb{N}$ and each component $a\in\{1,\dots,r\}$, there exists a sequence
$(\widetilde{G}_{\alpha,k})^{a}\in C_{c}^{\infty}(\Omega_{\alpha}),\qquad k\in\mathbb{N},$
such that
$\sum_{a=1}^{r}\left\|(\widetilde{G}_{\alpha,k})^{a}-\big(\psi_{\alpha}F\big)^{a}\circ\phi_{\alpha}^{-1}\right\|_{W^{m,p}(\Omega_{\alpha})}\xrightarrow[k\to\infty]{}0.$
Using these smooth functions, we define a smooth section $G_{\alpha,k}\in\Gamma(E|_{U_{\alpha}})$ by
$G_{\alpha,k}=\sum_{a=1}^{r}\Big[(\widetilde{G}_{\alpha,k})^{a}\circ\phi_{\alpha}\Big]\;e_{a}.$
Then, by the norm estimate from Lemma 3, it follows that
$\|G_{\alpha,k}-(\psi_{\alpha}F)|_{U_{\alpha}}\|_{W^{m,p}(E|_{U_{\alpha}})}\xrightarrow[k\to\infty]{}0.$
Extending each $G_{\alpha,k}$ by zero outside $U_{\alpha}$, we may view it as a global section with compact support, i.e., $G_{\alpha,k}\in\Gamma_{c}(E)$. Furthermore, the norm on the restriction is given by
$\|G_{\alpha,k}-(\psi_{\alpha}F)|_{U_{\alpha}}\|_{W^{m,p}(E|_{U_{\alpha}})}=\left(\sum_{s=0}^{m}\int_{U_{\alpha}}|\nabla_{w}^{s}(G_{\alpha,k}-(\psi_{\alpha}F)|_{U_{\alpha}})|^{p}_{g,h_{E}}d\lambda_{g}\right)^{\frac{1}{p}}.$
Since $\operatorname{supp}(G_{\alpha,k}-\psi_{\alpha}F)\subseteq U_{\alpha}$ and $\nabla^{s}_{w}$ is a local operator (Proposition 2), the integral over $U_{\alpha}$ coincides with the global integral:
$\left(\sum_{s=0}^{m}\int_{M}|\nabla_{w}^{s}(G_{\alpha,k}-\psi_{\alpha}F)|^{p}_{g,h_{E}}d\lambda_{g}\right)^{\frac{1}{p}}.$
Consequently,
$\|G_{\alpha,k}-\psi_{\alpha}F\|_{W^{m,p}(E)}\xrightarrow[k\to\infty]{}0.$
Now, fix $\alpha\in\mathbb{N}$ and an integer $q\in\mathbb{N}$. By the convergence established above, there exists an index $k_{\alpha,q}\in\mathbb{N}$ such that
$\|\psi_{\alpha}F-G_{\alpha,k_{\alpha,q}}\|_{W^{m,p}(E|_{U_{\alpha}})}<\frac{1}{2^{\alpha+q}}.$
Since $\operatorname{supp}(G_{\alpha,k})\subseteq U_{\alpha}$ for all $\alpha,k\in\mathbb{N}$ and the cover $(U_{\alpha})_{\alpha\in\mathbb{N}}$ is locally finite, the sum
$G_{q}:=\sum_{\alpha=1}^{\infty}G_{\alpha,k_{\alpha,q}}$
is locally finite and therefore defines a global smooth section $G_{q}\in\Gamma(E)$.
We now show that $G_{q}$ approximates $F$ in the Sobolev norm. Using the triangle inequality:
$\|F-G_{q}\|_{W^{m,p}(E)}\leq\sum_{\alpha=1}^{\infty}\|\psi_{\alpha}F-G_{\alpha,k_{\alpha,q}}\|_{W^{m,p}(E)}\leq\sum_{\alpha=1}^{\infty}\frac{1}{2^{\alpha+q}}=\frac{1}{2^{q}}.$
In particular,
$\|G_{q}\|_{W^{m,p}(E)}\leq\|F\|_{W^{m,p}(E)}+\|F-G_{q}\|_{W^{m,p}(E)}\leq\|F\|_{W^{m,p}(E)}+2^{-q}<\infty,$
which implies that $G_{q}\in W^{m,p}(E)\cap\Gamma(E)$ for all $q\in\mathbb{N}$.
Since $2^{-q}\to 0$ as $q\to\infty$, we conclude that $G_{q}\to F$ in $W^{m,p}(E)$. Thus, every section in $W^{m,p}(E)$ can be approximated in the Sobolev norm by smooth sections, proving that
$H^{m,p}(E)=W^{m,p}(E).$
Having established the equivalence between the completion-based spaces and those defined via weak derivatives, we conclude this section by addressing the geometric robustness of the theory. The following theorem ensures that, on compact manifolds, the Sobolev spaces $W^{m,p}(E)$ are intrinsic to the smooth structure of the bundle and the manifold, remaining invariant under changes of the underlying metrics and connection.
Theorem 3 (Geometric Invariance on Compact Manifolds).
Let $M$ be a compact Riemannian manifold and $E\to M$ a smooth vector bundle. The Sobolev spaces $W^{m,p}(E)$ are independent of the choice of the Riemannian metric $g$, the fiber metric $h_{E}$, and the compatible connection $\nabla^{E}$.
Proof.
By Theorem 2, $W^{m,p}(E)$ coincides with the completion-based space $H^{m,p}(E)$. The independence of $H^{m,p}(E)$ from the underlying geometric data on compact manifolds follows by generalizing the analytic arguments established for the scalar-valued case in [9, Theorem 2] to the vector bundle setting.∎
6.  Sobolev Embeddings and the Rellich–Kondrashov Theorem
Sobolev embeddings are of paramount importance when addressing partial differential equations. Just as in the case of scalar-valued functions, it is natural to inquire about the validity of these embeddings within the more general framework we have defined for smooth vector bundles. It turns out that if $M$ is a compact manifold, the classical results remain valid in this setting, as we demonstrate in this section.
Theorem 4 (Sobolev Embedding for Vector Bundles).
Let $(M,g)$ be a compact Riemannian manifold of dimension $n$, and let $E\longrightarrow M$ be a smooth vector bundle equipped with a fiber metric $h_{E}$ and a compatible connection $\nabla^{E}$. Then, for any integers $j\geq 0$ and $1\leq m<n$, and real numbers $p\in[1,n/m)$ and $q\in[1,p^{*}_{m}]$, we have
$W^{j+m,p}(E)\subseteq W^{j,q}(E),$
and this inclusion is continuous. Here $p^{*}_{m}=np/(n-mp)$ denotes the critical Sobolev exponent.
Proof.
Let $q\in[1,p_{m}^{*}]$ be arbitrary. Consider a finite open cover $(U_{\alpha},\phi_{\alpha})_{\alpha=1}^{N}$ of $M$ consisting of regular coordinate balls, each being the domain of a local frame $(e_{1,\alpha},\dots,e_{r,\alpha})$, and let $(\psi_{\alpha})_{\alpha=1}^{N}$ be a partition of unity subordinate to this cover. By the classical Sobolev embedding theorem in Euclidean domains, for each $\alpha\in\{1,\dots,N\}$, we have the continuous inclusion:
$W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))\subseteq W^{j,q}(\phi_{\alpha}(U_{\alpha})).$ (23)
Thus, there exists $C_{\alpha}>0$ such that $\|\cdot\|_{W^{j,q}(\phi_{\alpha}(U_{\alpha}))}\leq C_{\alpha}\|\cdot\|_{W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))}$.
Let $F\in W^{j+m,p}(E)$ and assume that $F=F^{a}_{\alpha}e_{a,\alpha}$ in $U_{\alpha}$. Note that by Proposition 1, $(\psi_{\alpha}F)|_{U_{\alpha}}\in W^{j+m,p}(E|_{U_{\alpha}})$. By Lemma 3, we know that for all $a\in\{1,\dots,r\}$ and all $\alpha\in\{1,\dots,N\}$, the function $\psi_{\alpha}F^{a}_{\alpha}\circ\phi_{\alpha}^{-1}$ belongs to $W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))$. Consequently, by (23), it also belongs to $W^{j,q}(\phi_{\alpha}(U_{\alpha}))$ and satisfies
$\|\psi_{\alpha}F^{a}_{\alpha}\circ\phi_{\alpha}^{-1}\|_{W^{j,q}(\phi_{\alpha}(U_{\alpha}))}\leq C_{\alpha}\|\psi_{\alpha}F^{a}_{\alpha}\circ\phi_{\alpha}^{-1}\|_{W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))}.$
Applying Lemma 3 again, since the component functions are in $W^{j,q}$, we conclude that $(\psi_{\alpha}F)|_{U_{\alpha}}\in W^{j,q}(E|_{U_{\alpha}})$. Furthermore, there exist constants $K_{\alpha,j,m},K_{\alpha,j}>0$ such that
$\|\psi_{\alpha}F\|_{W^{j,q}(E|_{U_{\alpha}})}\leq K_{\alpha,j}\sum_{a=1}^{r}\|\psi_{\alpha}F^{a}_{\alpha}\circ\phi^{-1}_{\alpha}\|_{W^{j,q}(\phi_{\alpha}(U_{\alpha}))},$
and
$K_{\alpha,j,m}\sum_{a=1}^{r}\|\psi_{\alpha}F^{a}_{\alpha}\circ\phi_{\alpha}^{-1}\|_{W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))}\leq\|\psi_{\alpha}F\|_{W^{j+m,p}(E|_{U_{\alpha}})}.$
Now, for each $s\leq m$, $(\psi_{\alpha}F)|_{U_{\alpha}}$ possesses a weak covariant derivative $\nabla_{w}^{s}(\psi_{\alpha}F)|_{U_{\alpha}}\in L^{q}(T^{(0,s)}(TU_{\alpha})\otimes E|_{U_{\alpha}})$. For each $\alpha$, we extend this derivative by zero outside $U_{\alpha}$ and denote the extension by $\nabla^{s}_{w}(\psi_{\alpha}F)$. By Proposition 2 (locality), $\operatorname{supp}(\nabla_{w}^{s}(\psi_{\alpha}F))\subseteq\operatorname{supp}((\psi_{\alpha}F)|_{U_{\alpha}})\subseteq U_{\alpha}$.
We claim that $\nabla^{s}_{w}(\psi_{\alpha}F)\in L^{q}(T^{(0,s)}(TM)\otimes E)$ is indeed the global weak derivative of $\psi_{\alpha}F$. Let $\Psi\in\Gamma_{c}(T^{(0,s)}(TM)\otimes E)$. Since $\operatorname{supp}(\psi_{\alpha}F)$ is a closed set contained in $U_{\alpha}$, there exists an open set $V_{\alpha}$ such that $\operatorname{supp}(\psi_{\alpha}F)\subseteq V_{\alpha}\subseteq\overline{V_{\alpha}}\subseteq U_{\alpha}$. Let $f_{\alpha}\in C^{\infty}(M)$ be a bump function such that $f_{\alpha}\equiv 1$ on $\overline{V_{\alpha}}$ and $\operatorname{supp}(f_{\alpha})\subseteq U_{\alpha}$. Define $\widetilde{\Psi}:=\Psi|_{U_{\alpha}}f_{\alpha}$. Note that $\operatorname{supp}(\widetilde{\Psi})\subseteq U_{\alpha}$ and $\widetilde{\Psi}\in\Gamma_{c}(T^{(0,s)}(TU_{\alpha})\otimes E|_{U_{\alpha}})$. Then
$\int_{M}\langle\nabla_{w}^{s}(\psi_{\alpha}F),\Psi\rangle_{g,h_{E}}d\lambda_{g}$ $=\int_{U_{\alpha}}\langle\nabla_{w}^{s}(\psi_{\alpha}F),\widetilde{\Psi}\rangle_{g,h_{E}}d\lambda_{g}$
$=\int_{U_{\alpha}}\langle\psi_{\alpha}F,(\nabla^{s})^{*}\widetilde{\Psi}\rangle_{h_{E}}d\lambda_{g}$
$=\int_{M}\langle\psi_{\alpha}F,(\nabla^{s})^{*}\Psi\rangle_{h_{E}}d\lambda_{g},$
where we used the fact that $\Psi\equiv\widetilde{\Psi}$ on $V_{\alpha}$ and the locality of $(\nabla^{s})^{*}$ implies $(\nabla^{s})^{*}\Psi\equiv(\nabla^{s})^{*}\widetilde{\Psi}$ on $V_{\alpha}$ (since $\psi_{\alpha}F$ vanishes outside $V_{\alpha}$). Thus, $\psi_{\alpha}F\in W^{j,q}(E)$. Since $F=\displaystyle\sum_{\alpha=1}^{N}\psi_{\alpha}F$, it follows that $F\in W^{j,q}(E)$.
Finally, we estimate the norm:
$\|F\|_{W^{j,q}(E)}$ $=\left\|\sum_{\alpha=1}^{N}\psi_{\alpha}F\right\|_{W^{j,q}(E)}\leq\sum_{\alpha=1}^{N}\|\psi_{\alpha}F\|_{W^{j,q}(E)}=\sum_{\alpha=1}^{N}\|\psi_{\alpha}F\|_{W^{j,q}(E|_{U_{\alpha}})}$
$\leq\sum_{\alpha=1}^{N}\sum_{a=1}^{r}K_{\alpha,j}\|\psi_{\alpha}F^{a}_{\alpha}\circ\phi_{\alpha}^{-1}\|_{W^{j,q}(\phi_{\alpha}(U_{\alpha}))}$
$\leq\sum_{\alpha=1}^{N}\sum_{a=1}^{r}K_{\alpha,j}C_{\alpha}\|\psi_{\alpha}F^{a}_{\alpha}\circ\phi_{\alpha}^{-1}\|_{W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))}.$
Defining $K:=\displaystyle\max_{1\leq\alpha\leq N}\{K_{\alpha,j},C_{\alpha}\}$ and using the lower bound for the local Sobolev norm:
$\|F\|_{W^{j,q}(E)}\leq K\sum_{\alpha=1}^{N}\frac{1}{K_{\alpha,j,m}}\|\psi_{\alpha}F\|_{W^{j+m,p}(E|_{U_{\alpha}})}\leq C\sum_{\alpha=1}^{N}\|\psi_{\alpha}F\|_{W^{j+m,p}(E)},$
where $C$ is a constant independent of $F$.
To conclude the continuity, we use the density of smooth sections (Theorem 2). Suppose $F\in\Gamma(E)\cap W^{j+m,p}(E)$. By the Leibniz rule for the tensor product connection:
$|\nabla^{s}(\psi_{\alpha}F)|_{g,h_{E}}\leq\sum_{t=0}^{s}\binom{s}{t}|\nabla^{s-t}\psi_{\alpha}|_{g}|\nabla^{t}F|_{g,h_{E}}.$
Taking $L^{p}$ norms, there exists $C^{\prime}>0$ such that
$\|\psi_{\alpha}F\|_{W^{j+m,p}(E)}\leq C^{\prime}\|F\|_{W^{j+m,p}(E)},\qquad\forall\alpha\in\{1,\dots,N\}.$ (24)
Thus,
$\|F\|_{W^{j,q}(E)}\leq NCC^{\prime}\|F\|_{W^{j+m,p}(E)}.$
Extending this result by density, we conclude that the inclusion is continuous. ∎
We also state and prove the Rellich–Kondrashov compactness theorem for vector bundles.
Theorem 5 (Rellich–Kondrashov for Vector Bundles).
Let $(M,g)$ be a compact Riemannian manifold of dimension $n$, and let $E\longrightarrow M$ be a smooth vector bundle with fiber metric $h_{E}$ and compatible connection $\nabla^{E}$. Then, for any integers $j\geq 0$ and $1\leq m<n$, and real numbers $p\in[1,n/m)$ and $q\in[1,p^{*}_{m})$, the inclusion
$W^{j+m,p}(E)\subseteq W^{j,q}(E)$
is compact.
Proof.
Let $q\in[1,p_{m}^{*})$ be arbitrary. Consider the same finite cover $(U_{\alpha},\phi_{\alpha})_{\alpha=1}^{N}$ and partition of unity $(\psi_{\alpha})_{\alpha=1}^{N}$ as before. By Theorem 4, we have the continuous embedding $W^{j+m,p}(E)\subseteq W^{j,q}(E)$.
Let $(F_{k})_{k\in\mathbb{N}}\subseteq W^{j+m,p}(E)$ be a bounded sequence; that is, there exists $C_{0}>0$ such that $\|F_{k}\|_{W^{j+m,p}(E)}\leq C_{0}$ for all $k$.
For each $k\in\mathbb{N}$ and each $\alpha\in\{1,\dots,N\}$, we have $(\psi_{\alpha}F_{k})|_{U_{\alpha}}\in W^{j+m,p}(E|_{U_{\alpha}})$ and $(\psi_{\alpha}F_{k})|_{U_{\alpha}}\in W^{j,q}(E|_{U_{\alpha}})$. By Lemma 3, there exist constants $K_{\alpha,j,m}>0$ and $K_{\alpha,j}>0$ such that the following norm estimates hold:
$\begin{split}K_{\alpha,j,m}\sum_{a=1}^{r}\|\psi_{\alpha}(F_{k})_{\alpha}^{a}\circ\phi_{\alpha}^{-1}\|_{W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))}&\leq\|\psi_{\alpha}F_{k}\|_{W^{j+m,p}(E|_{U_{\alpha}})},\\ \|\psi_{\alpha}F_{k}\|_{W^{j,q}(E|_{U_{\alpha}})}&\leq K_{\alpha,j}\sum_{a=1}^{r}\|\psi_{\alpha}(F_{k})_{\alpha}^{a}\circ\phi_{\alpha}^{-1}\|_{W^{j,q}(\phi_{\alpha}(U_{\alpha}))}.\end{split}$ (25)
By combining the first inequality of (25) with the estimate (24), it follows that for each $\alpha\in\{1,\dots,N\}$ and each component index $a\in\{1,\dots,r\}$, the sequence of functions $\left(\psi_{\alpha}(F_{k})_{\alpha}^{a}\circ\phi_{\alpha}^{-1}\right)_{k\in\mathbb{N}}$ is bounded in $W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))$:
$\|\psi_{\alpha}(F_{k})_{\alpha}^{a}\circ\phi_{\alpha}^{-1}\|_{W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))}\leq\frac{1}{K_{\alpha,j,m}}\|\psi_{\alpha}F_{k}\|_{W^{j+m,p}(E)}\leq\frac{C^{\prime}C_{0}}{K_{\alpha,j,m}}.$
By the classical Rellich–Kondrachov theorem for Euclidean domains, for each chart index $\alpha\in\{1,\dots,N\}$ and each fiber component index $a\in\{1,\dots,r\}$, the sequence $\left(\psi_{\alpha}(F_{k})_{\alpha}^{a}\circ\phi_{\alpha}^{-1}\right)_{k\in\mathbb{N}}$ admits a subsequence converging in $W^{j,q}(\phi_{\alpha}(U_{\alpha}))$. Since there are finitely many charts $\alpha\in\{1,\dots,N\}$ and finitely many fiber components $a\in\{1,\dots,r\}$, we may perform successive subsequence extractions. More precisely, by iteratively refining the sequence $(F_{k})$ a finite number of times, we obtain a subsequence (still denoted by $(F_{k})$) such that, for each chart $\alpha\in\{1,\dots,N\}$ and each fiber component $a\in\{1,\dots,r\}$, the sequence $\psi_{\alpha}(F_{k})_{\alpha}^{a}\circ\phi_{\alpha}^{-1}$ converges —and hence is Cauchy— in $W^{j,q}(\phi_{\alpha}(U_{\alpha}))$.
Let $\varepsilon>0$. Since the subsequence is Cauchy component-wise, there exists $N^{\prime}\in\mathbb{N}$ such that for all $k,l\geq N^{\prime}$,
$\sum_{a=1}^{r}\|\psi_{\alpha}(F_{k}-F_{l})_{\alpha}^{a}\circ\phi^{-1}_{\alpha}\|_{W^{j,q}(\phi_{\alpha}(U_{\alpha}))}<\frac{\varepsilon}{NK_{\alpha,j}},\qquad\forall\alpha\in\{1,\dots,N\}.$
Then, estimating the global norm:
$\|F_{k}-F_{l}\|_{W^{j,q}(E)}$ $=\left\|\sum_{\alpha=1}^{N}\psi_{\alpha}(F_{k}-F_{l})\right\|_{W^{j,q}(E)}\leq\sum_{\alpha=1}^{N}\|\psi_{\alpha}(F_{k}-F_{l})\|_{W^{j,q}(E)}$
$\leq\sum_{\alpha=1}^{N}K_{j,\alpha}\sum_{a=1}^{r}\|\psi_{\alpha}(F_{k}-F_{l})_{\alpha}^{a}\circ\phi_{\alpha}^{-1}\|_{W^{j,q}(\phi_{\alpha}(U_{\alpha}))}$
$<\sum_{\alpha=1}^{N}\frac{\varepsilon}{N}=\varepsilon.$
Thus, $(F_{k})_{k\in\mathbb{N}}$ is a Cauchy sequence in the Banach space $W^{j,q}(E)$, and therefore converges. We conclude that the embedding is compact. ∎
The results presented in this work establish the foundational properties of Sobolev spaces of sections of vector bundles over Riemannian manifolds. By bridging the gap between intrinsic geometric definitions and local analytical estimates, we have extended the classical Meyers–Serrin and Rellich–Kondrashov theorems to this more general setting. Finally, we must note that the global integration by parts formula derived in Section 3 (Theorem 1) remains valid for manifolds with boundary, provided the corresponding boundary integrals are taken into account. While the treatment of such boundary cases and the associated trace theorems lies beyond the immediate scope of this paper, the framework developed herein provides the necessary tools for addressing these problems in a purely intrinsic manner in future research.