### v2

Abstract
This article develops a unified and intrinsic framework for the theory of Sobolev spaces on vector bundles over Riemannian manifolds. The analytical core of our approach is an explicit higher-order geometric integration by parts formula, which characterizes the formal adjoint of the covariant derivative as a global differential operator. This identity is established on arbitrary Riemannian manifolds with boundary, without assuming completeness or compactness.
While first-order integration by parts identities are classical, explicit higher-order formulas with precise boundary terms are rarely stated in the literature. As applications of this framework, we recover the classical Meyers–Serrin theorem on arbitrary manifolds and, in the compact case, the Sobolev embedding and Rellich–Kondrachov compactness theorems, providing direct and self-contained proofs. At the end of this work we also stablish a Green Formula and we use it to stablish norm equivalence in Sobolev Spaces on vector bundles with closed manifold as base space and the Bochner laplacian operator. As a corollary we show that, in case of trivial vector bundles this equivalence reduces to a well known (but non proved rigorously in literature) result for closed manifolds and the Laplace-Beltrami operator.
By emphasizing intrinsic global arguments and sharp local-to-global norm equivalence estimates, rather than ad hoc coordinate patching, this work offers a transparent and accessible foundation for the study of Sobolev spaces on vector bundles, suitable for researchers in global analysis, differential geometry, and partial differential equations.
keywords:
Sobolev spaces, Vector bundles, Geometric integration by parts, Embedding theorems, Rellich-Kondrashov theorem
pacs:
MSC Classification]46E35, 58A99, 58C99
1 Introduction
The analysis of Sobolev spaces on vector bundles is a fundamental pillar of modern geometric analysis. A central issue in this theory is ensuring the equivalence between the two standard definitions of these spaces: the completion of the space of smooth sections with respect to the Sobolev norm, denoted by $H^{m,p}(E)$, and the space of sections possessing weak covariant derivatives in $L^{p}$, denoted by $W^{m,p}(E)$.
On Riemannian manifolds, the treatment of this equivalence and the associated structural results of Sobolev theory vary significantly across the literature. Standard references such as [1] and [2] often introduce Sobolev spaces as completions of smooth sections, while more intrinsically oriented approaches, such as [3], define them via distributional covariant derivatives. Establishing the identity $H^{m,p}(E)=W^{m,p}(E)$—the Meyers–Serrin theorem, which is classical in the scalar case—is therefore essential to reconcile these perspectives. For general vector bundles, this equivalence is also known and has been established in a very general setting by Guidetti [4] using the theory of linear partial differential operators and the theory of elliptic operators. We will also derive this result in an alternative manner, using only a geometrical intrinsic approach.
Beyond density, the fundamental theorems of Sobolev theory—embedding and compactness—are likewise well established in the literature for vector bundle–valued sections, particularly in the compact case. The work of Palais [5] remains a foundational reference for Sobolev embeddings and Rellich–Kondrachov compactness on compact manifolds. However, existing proofs are often scattered across different sources, rely on diverse analytical frameworks, or involve constructions that can obscure the underlying geometric structure, making them less accessible to a broader audience interested in both partial differential equations and differential geometry.
The primary purpose of this paper is not to reestablish these classical results, but rather to provide a global and intrinsic integration by parts formula for higher-order covariant derivatives acting on sections of smooth vector bundles over Riemannian manifolds with boundary (Theorem 2). This formula constitutes the analytical core of the work. Its proof is carried out in a fully geometric manner and does not rely on elliptic regularity, completeness, or compactness assumptions. By formulating the identity intrinsically and making all boundary terms explicit, we obtain a transparent characterization of the formal adjoint of the covariant derivative as a global differential operator.
As applications of this integration by parts formula, we revisit several classical results in the theory of Sobolev spaces on vector bundles. In Section 5, we recover the Meyers–Serrin theorem on general Riemannian manifolds by a direct argument based on the integration by parts identity and the classical scalar Meyers–Serrin theorem. The resulting statement for vector bundle–valued sections is in full agreement with the corresponding extensions available in the literature.
In the compact setting, treated in Section 6, we apply the same framework to provide streamlined proofs of the classical Sobolev embedding theorem and the Rellich–Kondrachov compactness theorem, fully consistent with known results.
By placing the integration by parts formula at the center of the theory, this work offers a unifying and geometrically transparent perspective on Sobolev spaces on vector bundles. The approach emphasizes global and intrinsic arguments and is designed to be accessible to a broad audience, including researchers in global analysis, differential geometry, and partial differential equations. We expect the framework developed here to serve both as a conceptual reference and as a flexible tool for further investigations involving vector bundles and boundary value problems.
2 Notation and Preliminaries
2.1 Notation and Conventions
Throughout this paper, $(M,g)$ denotes a smooth Riemannian manifold (not necessarily complete or compact), and $d\lambda_{g}$ denotes the Riemannian-Lebesgue measure induced by the metric $g$. If $M$ is a manifold with boundary, its boundary is denoted by $\partial M$.
Let $E\to M$ be a smooth vector bundle of finite rank equipped with a smooth fiber metric $h_{E}$ and a compatible connection $\nabla^{E}$. Following the notation in [6], $\Gamma(E)$ denotes the space of smooth sections of $E$, while $\Gamma_{c}(E)$ denotes the subspace of smooth sections with compact support. For the particular case of scalar-valued functions on a domain $\Omega\subseteq\mathbb{R}^{n}$, we employ the standard notation $C_{c}^{\infty}(\Omega)$.
For any integer $s\geq 0$, we denote by $T^{(0,s)}(TM)\otimes E$ the bundle of $E$-valued covariant tensor fields of rank $s$. Any connection on $TM$ together with $\nabla^{E}$ induces a connection on these tensor bundles, also denoted by $\nabla$, which is uniquely characterized by the Leibniz rule. We write $|\cdot|_{g,h_{E}}$ to denote the pointwise norm on the fibers induced by $g$ and $h_{E}$; note that for any section $\sigma$, $|\sigma|_{g,h_{E}}:M\to[0,\infty)$ is a non-negative function on $M$, which allows us to define the global norm of a section in the corresponding function spaces.
Finally, we adopt the Einstein summation convention, although summations will be written explicitly whenever deemed necessary for clarity or when the convention is not applicable.
2.2 Sobolev Spaces
We begin with the precise definitions of the Sobolev spaces used in this work.
Definition 1 (The space $H^{m,p}$).
Let $m\geq 0$ be an integer and $1\leq p<\infty$. We define the space of smooth sections with finite Sobolev norm, denoted by $S^{m,p}(E)$, as:
$S^{m,p}(E):=\left\{u\in\Gamma(E)\mathrel{\Big|}\|u\|_{m,p}<\infty\right\},$
where the norm is given by
$\|u\|_{m,p}:=\left(\sum_{s=0}^{m}\int_{M}|\nabla^{s}u|_{h_{E}}^{p}\,d\lambda_{g}\right)^{1/p}.$
Here, $\nabla^{s}u$ denotes the $s$-th iterated covariant derivative of $u$, viewed as a section of the bundle $T^{(0,s)}(TM)\otimes E$. The Sobolev space $H^{m,p}(E)$ is defined as the completion of $S^{m,p}(E)$ with respect to this norm.
To introduce the definition via weak covariant derivatives, we first recall the notion of the formal adjoint.
Definition 2 (Weak Covariant Derivative and the space $W^{m,p}$).
Let $s\geq 0$. The formal adjoint of the iterated covariant derivative $\nabla^{s}$, denoted by $(\nabla^{s})^{*}$, is the unique differential operator characterized by the identity
$\int_{M}\langle\nabla^{s}u,v\rangle_{h_{E}}\,d\lambda_{g}=\int_{M}\langle u,(\nabla^{s})^{*}v\rangle_{h_{E}}\,d\lambda_{g},$ (1)
for all $u\in\Gamma_{c}(E)$ and $v\in\Gamma_{c}(T^{(0,s)}(TM)\otimes E)$.
A section $u\in L^{1}_{\operatorname{loc}}(E)$ is said to admit a weak covariant derivative of order $s$ if there exists a section $v\in L^{1}_{\operatorname{loc}}(T^{(0,s)}(TM)\otimes E)$ such that
$\int_{M}\langle v,\psi\rangle_{h_{E}}\,d\lambda_{g}=\int_{M}\langle u,(\nabla^{s})^{*}\psi\rangle_{h_{E}}\,d\lambda_{g}$
holds for all test sections $\psi\in\Gamma_{c}(T^{(0,s)}(TM)\otimes E)$. If such a section $v$ exists, it is unique almost everywhere; we denote it by $\nabla_{w}^{s}u$.
Finally, the Sobolev space $W^{m,p}(E)$ is defined as:
[…]
This space is equipped with the same norm $\|\cdot\|_{m,p}$ defined in Definition 1.
To establish an explicit local characterization of the formal adjoint $(\nabla^{s})^{*}$, we first derive the local coordinate expression of the covariant derivative acting on tensor-valued sections.
Lemma 1.
Let $(M,g)$ be a Riemannian manifold and let $\nabla$ denote a connection on $TM$. Let $E\longrightarrow M$ be a smooth vector bundle with a connection $\nabla^{E}$. We denote by $\nabla$ the induced connection on the tensor product bundle $T^{(k,l)}(TM)\otimes E$, defined uniquely by the Leibniz rule:
$\nabla_{X}(F\otimes\sigma)=(\nabla_{X}F)\otimes\sigma+F\otimes\nabla^{E}_{X}\sigma,$
for all $X\in\Gamma(TM)$, $F\in\Gamma(T^{(k,l)}(TM))$, and $\sigma\in\Gamma(E)$.
Let $(E_{1},\dots,E_{n})$ be a local frame for $TM$ over an open set $U$, with dual coframe $(\varepsilon^{1},\dots,\varepsilon^{n})$, and let $(e_{1},\dots,e_{r})$ be a local frame for $E$ over $U$. If a section $F\in\Gamma(T^{(k,l)}(TM)\otimes E)$ is expressed locally as
$F=F^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}\,E_{i_{1}}\otimes\cdots\otimes E_{i_{k}}\otimes\varepsilon^{j_{1}}\otimes\cdots\otimes\varepsilon^{j_{l}}\otimes e_{a},$
then the components of the covariant derivative $\nabla_{E_{c}}F$ with respect to these frames are given by
$(\nabla_{E_{c}}F)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}$ $=E_{c}\big(F^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}\big)+\sum_{t=1}^{k}\Gamma^{i_{t}}_{cs}\,F^{i_{1}\dots s\dots i_{k}a}_{j_{1}\dots j_{l}}$
$\quad-\sum_{t=1}^{l}\Gamma^{s}_{cj_{t}}\,F^{i_{1}\dots i_{k}\,a}_{j_{1}\dots s\dots j_{l}}+\sum_{b=1}^{r}(A_{c})^{a}_{b}\,F^{i_{1}\dots i_{k}b}_{j_{1}\dots j_{l}},$
where the connection coefficients are defined by $\nabla_{E_{c}}E_{j}=\Gamma^{s}_{cj}E_{s}$ and $\nabla^{E}_{E_{c}}e_{b}=(A_{c})^{a}_{b}\,e_{a}$.
Proof.
We have
$\nabla_{E_{c}}F=\nabla_{E_{c}}\left(\sum_{a=1}^{r}F^{(a)}\otimes e_{a}\right),$
where $F^{(a)}\in\Gamma(T^{(k,l)}(TM))$ denotes the tensor field with components $(F^{(a)})^{i_{1}\dots i_{k}}_{j_{1}\dots j_{l}}=F^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}$. Applying the Leibniz rule for the induced connection yields
$\nabla_{E_{c}}F=\sum_{a=1}^{r}\left((\nabla_{E_{c}}F^{(a)})\otimes e_{a}+F^{(a)}\otimes\nabla^{E}_{E_{c}}e_{a}\right).$ (2)
Recall that for the tensor field $F^{(a)}$, the connection components are given by (see, e.g., [7, Proposition 4.16]):
$(\nabla_{E_{c}}F^{(a)})^{i_{1}\dots i_{k}}_{j_{1}\dots j_{l}}=E_{c}(F^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}})+\sum_{t=1}^{k}\Gamma^{i_{t}}_{cs}\,F^{i_{1}\dots s\dots i_{k}a}_{j_{1}\dots j_{l}}-\sum_{t=1}^{l}\Gamma^{s}_{cj_{t}}\,F^{i_{1}\dots i_{k}\,a}_{j_{1}\dots s\dots j_{l}}.$
For the second term in (2), we substitute $\nabla^{E}_{E_{c}}e_{b}=(A_{c})^{a}_{b}e_{a}$. Relabeling the summation index for the bundle frame as $b$, we obtain
$\sum_{b=1}^{r}F^{(b)}\otimes\nabla^{E}_{E_{c}}e_{b}=\sum_{b=1}^{r}F^{(b)}\otimes\left(\sum_{a=1}^{r}(A_{c})^{a}_{b}e_{a}\right)=\sum_{a=1}^{r}\left(\sum_{b=1}^{r}(A_{c})^{a}_{b}F^{(b)}\right)\otimes e_{a}.$
The component of this term along the basis element $e_{a}$ is $\sum_{b=1}^{r}(A_{c})^{a}_{b}F^{i_{1}\dots i_{k}b}_{j_{1}\dots j_{l}}$. Combining these contributions yields the asserted formula. ∎
Remark 1.
We regard the total covariant derivative $\nabla F$ as a section of $T^{(k,l+1)}(TM)\otimes E$. In terms of the local frames defined above, we write
$\nabla F=\sum_{c=1}^{n}(\nabla_{E_{c}}F)\otimes\varepsilon^{c}.$
Consequently, the components of $\nabla F$ are defined by $(\nabla F)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}c}:=(\nabla_{E_{c}}F)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}$. Explicitly,
$(\nabla F)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}c}$ $=E_{c}\big(F^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}\big)+\sum_{t=1}^{k}\Gamma^{i_{t}}_{cs}\,F^{i_{1}\dots s\dots i_{k}a}_{j_{1}\dots j_{l}}$
$\quad-\sum_{t=1}^{l}\Gamma^{s}_{cj_{t}}\,F^{i_{1}\dots i_{k}\,a}_{j_{1}\dots s\dots j_{l}}+\sum_{b=1}^{r}(A_{c})^{a}_{b}\,F^{i_{1}\dots i_{k}b}_{j_{1}\dots j_{l}}.$
We emphasize that we adopt the convention where the differentiation index (here $c$) corresponds to the last covariant index in the component representation.
3 The Geometric Integration by Parts Formula
With the local properties of the connection established, we now turn to the analytical core of this work. The following theorem provides a rigorous global integration by parts formula for higher-order covariant derivatives. Unlike standard Euclidean formulas, this result accounts for the intrinsic curvature of the bundle through the connection and explicitly identifies the boundary pairings required for the study of boundary value problems.
Theorem 2 (Integration by Parts for Higher-Order Derivatives).
Let $(M,g)$ be a Riemannian manifold with boundary $\partial M$, and let $\nabla$ denote the Levi-Civita connection on $TM$. Let $E\to M$ be a smooth vector bundle equipped with a fiber metric $h_{E}$ and a connection $\nabla^{E}$ that is compatible with $h_{E}$. We continue to denote by $\nabla$ the induced connection on the tensor bundles $T^{(k,l)}(TM)\otimes E$.
For any integer $s\geq 1$, and sections $F\in\Gamma_{c}\bigl(T^{(k,l)}(TM)\otimes E\bigr)$ and $G\in\Gamma_{c}\bigl(T^{(k,l+s)}(TM)\otimes E\bigr)$, the following identity holds:
$\int_{M}\langle\nabla^{s}F,G\rangle_{g,h_{E}}\,d\lambda_{g}$ $=(-1)^{s}\int_{M}\big\langle F,(\operatorname{tr}_{g}\circ\nabla)^{s}G\big\rangle_{g,h_{E}}\,d\lambda_{g}$
$\quad+\sum_{j=0}^{s-1}(-1)^{\,s-1-j}\int_{\partial M}\Big\langle\nabla^{j}F,\,\iota_{\nu}\bigl((\operatorname{tr}_{g}\circ\nabla)^{\,s-1-j}G\bigr)\Big\rangle_{g,h_{E}}\,d\lambda_{\widetilde{g}}$,
where $\nu$ denotes the outward unit normal to $\partial M$, and $\iota_{\nu}$ denotes the contraction (interior multiplication) of $\nu$ into the last covariant index of the tensor factor. Explicitly, if $H\in\Gamma\bigl(T^{(k,m+1)}(TM)\otimes E\bigr)$, then $\iota_{\nu}H\in\Gamma\bigl(T^{(k,m)}(TM)\otimes E\bigr)$ is given in local coordinates by
$\bigl(\iota_{\nu}H\bigr)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{m}}:=\nu^{q}\,H^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{m}q},$
where $\nu=\nu^{q}\partial_{q}$.
Here, $\operatorname{tr}_{g}$ denotes the metric contraction of the last two covariant indices of $\nabla H$. Specifically, if $H\in\Gamma\bigl(T^{(k,m)}(TM)\otimes E\bigr)$, then $\nabla H\in\Gamma\bigl(T^{(k,m+1)}(TM)\otimes E\bigr)$ and locally
$\bigl(\operatorname{tr}_{g}(\nabla H)\bigr)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{m-1}}:=g^{pq}\,(\nabla H)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{m-1}pq}.$
That is, the derivative index is contracted with the last covariant index inherited from $H$.
Proof.
We proceed by induction on $s$. The strategy is to shift derivatives from $F$ to $G$ via covariant integration by parts, tracking the boundary terms that arise at each step.
Base Case ($s=1$). Fix a point $p\in M$. Since $\nabla$ on $TM$ is the Levi-Civita connection, we can choose geodesic normal coordinates centered at $p$, with coordinate frame $(\partial_{1},\dots,\partial_{n})$. In these coordinates, $g_{ij}(p)=\delta_{ij}$ and all connection coefficients vanish at $p$. In particular, $\nabla_{\partial_{i}}\partial_{j}(p)=0$. Furthermore, we choose a local orthonormal frame $(e_{1},\dots,e_{r})$ for the bundle $E$ defined in a neighborhood of $p$, such that $h_{E}(e_{a},e_{b})(p)=\delta_{ab}$.
Let $F\in\Gamma_{c}\bigl(T^{(k,l)}(TM)\otimes E\bigr)$ and $G\in\Gamma_{c}\bigl(T^{(k,l+1)}(TM)\otimes E\bigr)$. We define a map $\Phi:\mathfrak{X}(M)\longrightarrow C^{\infty}(M)$ by
$\Phi(Y):=\big\langle F,\,\iota_{Y}G\big\rangle_{g,h_{E}}.$
The map $Y\mapsto\iota_{Y}G$ is $C^{\infty}(M)$-linear in $Y$ and the inner product is bilinear, so $\Phi$ defines a 1-form. Let $X:=\Phi^{\sharp}$ be the dual vector field, characterized by $\langle X,Y\rangle_{g}=\Phi(Y)=\big\langle F,\,\iota_{Y}G\big\rangle_{g,h_{E}}$. In particular, for each coordinate vector $\partial_{c}$,
$\langle X,\partial_{c}\rangle_{g}=\big\langle F,\,\iota_{\partial_{c}}G\big\rangle_{g,h_{E}}.$ (3)
At $p$, the divergence of $X$ is given by
$\operatorname{div}X(p)=\sum_{c=1}^{n}\big\langle\nabla_{\partial_{c}}X,\partial_{c}\big\rangle_{g}(p).$
Using the compatibility of $\nabla$ with $g$ and the fact that $\nabla_{\partial_{c}}\partial_{c}(p)=0$, we obtain
$\operatorname{div}X(p)=\sum_{c=1}^{n}\partial_{c}\big(\langle X,\partial_{c}\rangle_{g}\big)(p).$ (4)
Substituting (3) into (4) and using the product rule:
$\operatorname{div}X(p)=\sum_{c=1}^{n}\Big(\big\langle\nabla_{\partial_{c}}F,\,\iota_{\partial_{c}}G\big\rangle_{g,h_{E}}(p)+\big\langle F,\,\nabla_{\partial_{c}}\bigl(\iota_{\partial_{c}}G\bigr)\big\rangle_{g,h_{E}}(p)\Big).$
Using the Leibniz rule for contraction $\nabla_{\partial_{c}}\bigl(\iota_{\partial_{c}}G\bigr)=\iota_{\partial_{c}}\bigl(\nabla_{\partial_{c}}G\bigr)+\iota_{\nabla_{\partial_{c}}\partial_{c}}G$, and noting that the second term vanishes at $p$, we arrive at:
$\operatorname{div}X(p)=\sum_{c=1}^{n}\Big(\big\langle\nabla_{\partial_{c}}F,\,\iota_{\partial_{c}}G\big\rangle_{g,h_{E}}(p)+\big\langle F,\,\iota_{\partial_{c}}(\nabla_{\partial_{c}}G)\big\rangle_{g,h_{E}}(p)\Big).$
We now identify these sums using local coordinates. Let
$F=F^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}\partial_{i_{1}}\otimes\dots\otimes\partial_{i_{k}}\otimes dx^{j_{1}}\otimes\cdots\otimes dx^{j_{l}}\otimes e_{a},$
$G=G^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}t}\partial_{i_{1}}\otimes\dots\otimes\partial_{i_{k}}\otimes dx^{j_{1}}\otimes\cdots\otimes dx^{j_{l}}\otimes dx^{t}\otimes e_{a}.$
At $p$, where $g_{ij}=\delta_{ij}$ and $h_{ab}=\delta_{ab}$, the first sum becomes:
$\sum_{c=1}^{n}\big\langle\nabla_{\partial_{c}}F,\,\iota_{\partial_{c}}G\big\rangle_{g,h_{E}}(p)$ $=\sum_{c=1}^{n}\sum_{a=1}^{r}\sum_{\begin{subarray}{c}1\leq i_{1},\dots,i_{k}\leq n\\ 1\leq j_{1},\dots,j_{l}\leq n\end{subarray}}(\nabla_{\partial_{c}}F)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}(p)\,G^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}c}(p)$
$=\big\langle\nabla F,\,G\big\rangle_{g,h_{E}}(p).$
For the second sum, noting that $(\nabla_{\partial_{c}}G)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}t}=(\nabla G)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}t\,c}$, we have:
$\sum_{c=1}^{n}\big\langle F,\,\iota_{\partial_{c}}(\nabla_{\partial_{c}}G\bigr)\big\rangle_{g,h_{E}}(p)$ $=\sum_{c=1}^{n}\sum_{a=1}^{r}\sum_{\begin{subarray}{c}1\leq i_{1},\dots,i_{k}\leq n\\ 1\leq j_{1},\dots,j_{l}\leq n\end{subarray}}F^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}(p)\,(\nabla G)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}c\,c}(p).$
Since $g^{pq}(p)=\delta^{pq}$ and $\bigl(\operatorname{tr}_{g}(\nabla G)\bigr)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}=g^{pq}\,(\nabla G)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}pq}$, we have
$\bigl(\operatorname{tr}_{g}(\nabla G)\bigr)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}}(p)=\sum_{c=1}^{n}(\nabla G)^{i_{1}\dots i_{k}a}_{j_{1}\dots j_{l}c\,c}(p).$
Thus,
$\sum_{c=1}^{n}\big\langle F,\,\iota_{\partial_{c}}(\nabla_{\partial_{c}}G)\big\rangle_{g,h_{E}}(p)=\big\langle F,\,\operatorname{tr}_{g}(\nabla G)\big\rangle_{g,h_{E}}(p).$
Combining these results, we have $\operatorname{div}X=\langle\nabla F,G\rangle_{g,h_{E}}+\langle F,\operatorname{tr}_{g}(\nabla G)\rangle_{g,h_{E}}$ globally. Integrating and applying the Divergence Theorem yields:
$\int_{M}\big\langle\nabla F,G\big\rangle_{g,h_{E}}\,d\lambda_{g}=-\int_{M}\big\langle F,\operatorname{tr}_{g}(\nabla G)\big\rangle_{g,h_{E}}\,d\lambda_{g}+\int_{\partial M}\big\langle F,\,\iota_{\nu}G\big\rangle_{g,h_{E}}\,d\lambda_{\widetilde{g}}$.
Inductive Step. Assume the identity holds for some $s\geq 1$. Let $G\in\Gamma_{c}\bigl(T^{(k,l+s+1)}(TM)\otimes E\bigr)$. Applying the base case ($s=1$) to $(\nabla^{s}F,G)$, where $\nabla^{s}F\in\Gamma_{c}(T^{(k,l+s)}(TM)\otimes E)$:
$\int_{M}\langle\nabla^{s+1}F,G\rangle_{g,h_{E}}\,d\lambda_{g}=-\int_{M}\big\langle\nabla^{s}F,\operatorname{tr}_{g}(\nabla G)\big\rangle_{g,h_{E}}\,d\lambda_{g}+\int_{\partial M}\big\langle\nabla^{s}F,\,\iota_{\nu}G\big\rangle_{g,h_{E}}\,d\lambda_{\widetilde{g}}$.
Define $H:=\operatorname{tr}_{g}(\nabla G)$. Applying the inductive hypothesis to $(F,H)$, we obtain:
$\int_{M}\langle\nabla^{s}F,H\rangle_{g,h_{E}}\,d\lambda_{g}$ $=(-1)^{s}\int_{M}\big\langle F,(\operatorname{tr}_{g}\circ\nabla)^{s}H\big\rangle_{g,h_{E}}\,d\lambda_{g}$
$\quad+\sum_{j=0}^{s-1}(-1)^{\,s-1-j}\int_{\partial M}\Big\langle\nabla^{j}F,\,\iota_{\nu}\bigl((\operatorname{tr}_{g}\circ\nabla)^{\,s-1-j}H\bigr)\Big\rangle_{g,h_{E}}\,d\lambda_{\widetilde{g}}$.
Substituting $H=\operatorname{tr}_{g}(\nabla G)$, we observe that $(\operatorname{tr}_{g}\circ\nabla)^{s}H=(\operatorname{tr}_{g}\circ\nabla)^{s+1}G$ and $(\operatorname{tr}_{g}\circ\nabla)^{\,s-1-j}H=(\operatorname{tr}_{g}\circ\nabla)^{\,s-j}G$. Substituting this result back into the expansion of $\int_{M}\langle\nabla^{s+1}F,G\rangle_{g,h_{E}}\,d\lambda_{g}$, the interior integral term becomes $(-1)^{s+1}\int_{M}\big\langle F,(\operatorname{tr}_{g}\circ\nabla)^{s+1}G\big\rangle_{g,h_{E}}\,d\lambda_{g}$. The boundary sum becomes:
[…]
$=\sum_{j=0}^{s-1}(-1)^{\,s-j}\int_{\partial M}\Big\langle\nabla^{j}F,\,\iota_{\nu}\bigl((\operatorname{tr}_{g}\circ\nabla)^{\,s-j}G\bigr)\Big\rangle_{g,h_{E}}\,d\lambda_{\widetilde{g}}$.
Finally, adding the isolated boundary term from the first step $\int_{\partial M}\big\langle\nabla^{s}F,\,\iota_{\nu}G\big\rangle_{g,h_{E}}\,d\lambda_{\widetilde{g}}$ (which corresponds to $j=s$ since $(-1)^{s-s}=1$ and $(\operatorname{tr}\circ\nabla)^{0}=\operatorname{Id}$), we extend the sum index to $s$:
$\sum_{j=0}^{s}(-1)^{\,s-j}\int_{\partial M}\Big\langle\nabla^{j}F,\,\iota_{\nu}\bigl((\operatorname{tr}_{g}\circ\nabla)^{\,s-j}G\bigr)\Big\rangle_{g,h_{E}}\,d\lambda_{\widetilde{g}}$.
This yields exactly the formula for $s+1$, completing the induction. ∎
To handle the multi-indices rigorously, for a fixed $s\geq 1$, we denote ordered tuples (blocks) of indices in $\{1,\dots,n\}$ as $I_{s}=(i_{1},\dots,i_{s})$, $J_{s}=(j_{1},\dots,j_{s})$, etc. We adopt the Einstein summation convention for these blocks. We define:
$\partial_{I_{s}}:=\partial_{i_{s}}\cdots\partial_{i_{1}},\quad g^{I_{s}J_{s}}:=\prod_{l=1}^{s}g^{i_{l}j_{l}},\quad g_{J_{s}T_{s}}:=\prod_{l=1}^{s}g_{j_{l}t_{l}}.$
Also, let $\beta_{I_{s}}\in\mathbb{N}_{0}^{n}$ be the multi-index associated with the block $I_{s}$ such that $\partial^{\beta_{I_{s}}}\equiv\partial_{I_{s}}$.
The integration by parts formula derived above is of fundamental importance because it allows us to derive an explicit coordinate expression for the formal adjoint of the $s$-th covariant derivative.
Corollary 3 (Local Expression of the Formal Adjoint).
Let $(M,g)$ be a Riemannian manifold without boundary, and let $E\to M$ be a vector bundle equipped with a fiber metric $h_{E}$ and a compatible connection $\nabla^{E}$. For $s\geq 1$, the formal adjoint of
$\nabla^{s}:\Gamma(E)\longrightarrow\Gamma\bigl(T^{(0,s)}(TM)\otimes E\bigr)$
is given globally by
$(\nabla^{s})^{*}=(-1)^{s}(\operatorname{tr}_{g}\circ\nabla)^{s}.$
Furthermore, in a local chart $(U,x^{1},\dots,x^{n})$, if a section $G\in\Gamma\bigl(T^{(0,s)}(TM)\otimes E\bigr)$ has components $G^{a}_{J_{s}}:=G^{a}_{j_{1}\dots j_{s}}$, then
$\begin{split}\bigl((\nabla^{s})^{*}G\bigr)^{a}&=(-1)^{s}\Bigg[\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(\sqrt{\det(g)}\,g^{I_{s}J_{s}}\,G^{a}_{J_{s}}\Big)\\ &\quad+\sum_{b=1}^{r}\sum_{|\beta|\leq s-1}\bigl(C_{\beta}\bigr)^{a\,J_{s}}_{b}\,\partial^{\beta}\!\bigl(G^{b}_{J_{s}}\bigr)\Bigg],\end{split}$ (5)
where the coefficients $\bigl(C_{\beta}\bigr)^{a\,J_{s}}_{b}$ are smooth functions on $U$ depending solely on the metric components $g_{ij}$ and its inverse $g^{ij}$, the connection coefficients $\Gamma^{k}_{ij}$ and $(A_{i})^{a}_{b}$, and their partial derivatives up to order s1s-1.
Proof.
Since $\partial M=\emptyset$, the boundary terms in Theorem 2 vanish. Consequently, for all $F\in\Gamma_{c}(E)$ and $G\in\Gamma_{c}\bigl(T^{(0,s)}(TM)\otimes E\bigr)$, we have
$\int_{M}\langle\nabla^{s}F,G\rangle_{g,h_{E}}\,d\lambda_{g}=(-1)^{s}\int_{M}\big\langle F,(\operatorname{tr}_{g}\circ\nabla)^{s}G\big\rangle_{g,h_{E}}\,d\lambda_{g}.$
By the uniqueness of the formal adjoint (see, e.g., [3]), it follows that
$(\nabla^{s})^{*}=(-1)^{s}(\operatorname{tr}_{g}\circ\nabla)^{s}.$
It remains to verify the local formula (5). Fix a chart $(U,x^{1},\dots,x^{n})$ and a local frame $(e_{1},\dots,e_{r})$ for $E|_{U}$. Let $m\geq 0$ and let $H\in\Gamma\bigl(T^{(0,m+1)}(TM)\otimes E\bigr)$ with components $H^{a}_{J_{m}q}=H^{a}_{j_{1}\dots j_{m}q}$. By definition, the components of the trace are
$((\operatorname{tr}_{g}\circ\nabla)H)^{a}_{J_{m}}=g^{pq}\,(\nabla_{\partial_{p}}H)^{a}_{J_{m}q}.$
Applying Lemma 1 (with $k=0$, $l=m+1$), we expand the covariant derivative:
$\begin{split}((\operatorname{tr}_{g}\circ\nabla)H)^{a}_{J_{m}}&=g^{pq}\partial_{p}\bigl(H^{a}_{J_{m}q}\bigr)-\sum_{\ell=1}^{m}\Gamma^{\alpha}_{pj_{\ell}}\,g^{pq}\,H^{a}_{j_{1}\dots j_{\ell-1}\alpha j_{\ell+1}\dots j_{m}q}\\ &\quad-\Gamma^{\alpha}_{pq}\,g^{pq}\,H^{a}_{J_{m}\alpha}+\sum_{b=1}^{r}(A_{p})^{a}_{b}\,g^{pq}H^{b}_{J_{m}q}.\end{split}$ (6)
We recall the identity $\Gamma_{\alpha p}^{\alpha}=\frac{1}{\sqrt{\det(g)}}\partial_{p}\left(\sqrt{\det(g)}\right)$ and the formula for the derivative of the inverse metric, $\partial_{p}(g^{pq})=-\Gamma^{p}_{p\alpha}g^{\alpha q}-\Gamma^{q}_{p\alpha}g^{p\alpha}$. These imply the following divergence identity for any collection of functions $(Y_{q})_{q=1}^{n}$:
$g^{pq}\partial_{p}\bigl(Y_{q}\bigr)-\Gamma^{\alpha}_{pq}g^{pq}Y_{\alpha}=\frac{1}{\sqrt{\det(g)}}\,\partial_{p}\Big(\sqrt{\det(g)}\,g^{pq}Y_{q}\Big).$ (7)
Indeed, expanding the right-hand side of (7):
$\frac{1}{\sqrt{\det(g)}}\partial_{p}$ $\left(\sqrt{\det(g)}g^{pq}Y_{q}\right)$
$=\frac{1}{\sqrt{\det(g)}}\bigg[\partial_{p}\left(\sqrt{\det(g)}\right)g^{pq}Y_{q}+\sqrt{\det(g)}\,\partial_{p}(g^{pq})Y_{q}+\sqrt{\det(g)}\,g^{pq}\partial_{p}(Y_{q})\bigg]$
$=\Gamma_{\alpha p}^{\alpha}g^{pq}Y_{q}+\partial_{p}(g^{pq})Y_{q}+g^{pq}\partial_{p}(Y_{q})$
$=\Gamma_{\alpha p}^{\alpha}g^{pq}Y_{q}+\left(-\Gamma_{p\alpha}^{p}g^{\alpha q}Y_{q}-\Gamma_{p\alpha}^{q}g^{p\alpha}Y_{q}\right)+g^{pq}\partial_{p}(Y_{q}).$
Observing that the first term cancels with the second (after relabeling indices), and relabeling $\alpha\leftrightarrow q$ in the third term, we recover $g^{pq}\partial_{p}(Y_{q})-\Gamma_{pq}^{\alpha}g^{pq}Y_{\alpha}$.
Applying (7) with $Y_{q}=H^{a}_{J_{m}q}$, equation (6) simplifies to:
$\begin{split}((\operatorname{tr}_{g}\circ\nabla)H)^{a}_{J_{m}}&=\frac{1}{\sqrt{\det(g)}}\,\partial_{p}\Big(\sqrt{\det(g)}\,g^{pq}H^{a}_{J_{m}q}\Big)\\ &\quad-\sum_{\ell=1}^{m}\Gamma^{\alpha}_{pj_{\ell}}\,g^{pq}\,H^{a}_{j_{1}\dots j_{\ell-1}\alpha j_{\ell+1}\dots j_{m}q}+\sum_{b=1}^{r}(A_{p})^{a}_{b}\,g^{pq}H^{b}_{J_{m}q}.\end{split}$ (8)
We now prove by induction on $s\geq 1$ that for any $G\in\Gamma\bigl(T^{(0,s)}(TM)\otimes E\bigr)$, the following holds:
$\begin{split}\bigl((\operatorname{tr}_{g}\circ\nabla)^{s}G\bigr)^{a}&=\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(\sqrt{\det(g)}\,g^{I_{s}J_{s}}\,G^{a}_{J_{s}}\Big)\\ &\quad+\sum_{b=1}^{r}\sum_{|\beta|\leq s-1}\bigl(C_{\beta}\bigr)^{a\,J_{s}}_{b}\,\partial^{\beta}\!\bigl(G^{b}_{J_{s}}\bigr).\end{split}$ (9)
For $s=1$, setting $m=0$ in (8) yields
$((\operatorname{tr}_{g}\circ\nabla)G)^{a}=\frac{1}{\sqrt{\det(g)}}\,\partial_{p}\Big(\sqrt{\det(g)}\,g^{pq}G^{a}_{q}\Big)+\sum_{b=1}^{r}(A_{p})^{a}_{b}\,g^{pq}G^{b}_{q},$
which matches the form of (9) with $|\beta|=0$.
Now, assume (9) holds for some $s\geq 1$. We prove the case for $s+1$. Let $G\in\Gamma\bigl(T^{(0,s+1)}(TM)\otimes E\bigr)$ and define $H:=(\operatorname{tr}_{g}\circ\nabla)G\in\Gamma\bigl(T^{(0,s)}(TM)\otimes E\bigr)$. By the inductive hypothesis applied to $H$:
$\begin{split}\bigl((\operatorname{tr}_{g}\circ\nabla)^{s}H\bigr)^{a}&=\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(\sqrt{\det(g)}\,g^{I_{s}J_{s}}\,H^{a}_{J_{s}}\Big)\\ &\quad+\sum_{b=1}^{r}\sum_{|\beta|\leq s-1}\bigl(C_{\beta}\bigr)^{a\,J_{s}}_{b}\,\partial^{\beta}\!\bigl(H^{b}_{J_{s}}\bigr).\end{split}$ (10)
Since $(\operatorname{tr}_{g}\circ\nabla)^{s}H=(\operatorname{tr}_{g}\circ\nabla)^{s+1}G$, we expand $H^{a}_{J_{s}}$ in terms of $G$ using (8) with $m=s$:
$\begin{split}H^{a}_{J_{s}}&=\frac{1}{\sqrt{\det(g)}}\,\partial_{p}\Big(\sqrt{\det(g)}\,g^{pq}G^{a}_{J_{s}q}\Big)\\ &\quad-\sum_{\ell=1}^{s}\Gamma^{\alpha}_{pj_{\ell}}\,g^{pq}\,G^{a}_{j_{1}\dots j_{\ell-1}\alpha j_{\ell+1}\dots j_{s}q}+\sum_{b=1}^{r}(A_{p})^{a}_{b}\,g^{pq}G^{b}_{J_{s}q}.\end{split}$ (11)
Substituting (11) into the principal term of (10) splits it into three parts:
$\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(\sqrt{\det(g)}\,g^{I_{s}J_{s}}\,H^{a}_{J_{s}}\Big)=\mathrm{(I)}+\mathrm{(II)}+\mathrm{(III)},$ (12)
where
$\mathrm{(I)}$ $:=\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(g^{I_{s}J_{s}}\partial_{p}\big(\sqrt{\det(g)}\,g^{pq}G^{a}_{J_{s}q}\big)\Big),$
$\mathrm{(II)}$ $:=-\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(\sqrt{\det(g)}\,g^{I_{s}J_{s}}\sum_{\ell=1}^{s}\Gamma^{\alpha}_{pj_{\ell}}\,g^{pq}\,G^{a}_{j_{1}\dots j_{\ell-1}\alpha j_{\ell+1}\dots j_{s}q}\Big),$
$\mathrm{(III)}$ $:=\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(\sqrt{\det(g)}\,g^{I_{s}J_{s}}\sum_{b=1}^{r}(A_{p})^{a}_{b}\,g^{pq}G^{b}_{J_{s}q}\Big).$
For term $\mathrm{(I)}$, we utilize the identity $\phi\partial_{p}f=\partial_{p}(\phi f)-(\partial_{p}\phi)f$. Applying this with $\phi=g^{I_{s}J_{s}}$ and $f=\sqrt{\det(g)}\,g^{pq}G^{a}_{J_{s}q}$, we have:
$g^{I_{s}J_{s}}\partial_{p}\big(\sqrt{\det(g)}\,g^{pq}G^{a}_{J_{s}q}\big)=\partial_{p}\Big(g^{I_{s}J_{s}}\sqrt{\det(g)}\,g^{pq}G^{a}_{J_{s}q}\Big)-\partial_{p}\big(g^{I_{s}J_{s}}\big)\sqrt{\det(g)}\,g^{pq}G^{a}_{J_{s}q}.$
Substituting this back into $\mathrm{(I)}$:
$\mathrm{(I)}$ $=\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\partial_{p}\Big(\sqrt{\det(g)}\,g^{I_{s}J_{s}}g^{pq}\,G^{a}_{J_{s}q}\Big)$ (13)
$\quad-\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s}}\Big(\partial_{p}\big(g^{I_{s}J_{s}}\big)\sqrt{\det(g)}\,g^{pq}G^{a}_{J_{s}q}\Big).$
The first term of (13) is exactly the principal term required for the order $s+1$. If we define the block indices for order $s+1$ as $I_{s+1}=(I_{s},p)$ and $J_{s+1}=(J_{s},q)$, then $g^{I_{s+1}J_{s+1}}=g^{I_{s}J_{s}}g^{pq}$ and $\partial_{I_{s+1}}=\partial_{I_{s}}\partial_{p}$. Thus, this term is
$\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s+1}}\Big(\sqrt{\det(g)}\,g^{I_{s+1}J_{s+1}}\,G^{a}_{J_{s+1}}\Big).$
The second term of (13) involves $\partial_{I_{s}}$ acting on a term containing $G$ with no derivatives. By the Leibniz rule, the highest order of derivatives acting on $G$ in this term is $s$. Thus, there exist smooth coefficients $\bigl(B_{\beta}\bigr)$ such that this remainder term is of the form:
$\sum_{c=1}^{r}\sum_{|\beta|\leq s}\bigl(B_{\beta}\bigr)^{a\,J_{s+1}}_{c}\,\partial^{\beta}\!\bigl(G^{c}_{J_{s+1}}\bigr).$ (14)
For $\mathrm{(II)}$, the term inside $\partial_{I_{s}}$ contains $G$ algebraically (zero derivatives). Thus, applying $\partial_{I_{s}}$ yields terms with at most $s$ derivatives on $G$. Similarly, for $\mathrm{(III)}$, the dependence on $G$ is linear without additional derivatives, so expanding yields terms of order at most $s$. Combining these, there exist coefficients $\bigl(\widetilde{B}_{\beta}\bigr)$ such that:
$\mathrm{(II)}+\mathrm{(III)}=\sum_{c=1}^{r}\sum_{|\beta|\leq s}\bigl(\widetilde{B}_{\beta}\bigr)^{a\,J_{s+1}}_{c}\,\partial^{\beta}\!\bigl(G^{c}_{J_{s+1}}\bigr).$ (15)
Finally, consider the lower-order term from (10):
$\sum_{b=1}^{r}\sum_{|\beta|\leq s-1}\bigl(C_{\beta}\bigr)^{a\,J_{s}}_{b}\,\partial^{\beta}\!\bigl(H^{b}_{J_{s}}\bigr).$
Substituting (11) into $H^{b}_{J_{s}}$, we see that $H$ involves first-order derivatives of $G$. Therefore, applying $\partial^{\beta}$ (where $|\beta|\leq s-1$) to $H$ yields derivatives of $G$ of order at most (s1)+1=s(s-1)+1=s. Thus, there exist coefficients $\bigl(\overline{D}_{\beta}\bigr)$ such that this sum reduces to:
$\sum_{c=1}^{r}\sum_{|\beta|\leq s}\bigl(\overline{D}_{\beta}\bigr)^{a\,J_{s+1}}_{c}\,\partial^{\beta}\!\bigl(G^{c}_{J_{s+1}}\bigr).$ (16)
Combining the principal term from (13) with the remainder terms (14), (15), and (16), we conclude that there exist smooth coefficients $\bigl(C_{\beta}\bigr)$ with $|\beta|\leq s$ such that
$\bigl((\operatorname{tr}_{g}\circ\nabla)^{s+1}G\bigr)^{a}$ $=\frac{1}{\sqrt{\det(g)}}\,\partial_{I_{s+1}}\Big(\sqrt{\det(g)}\,g^{I_{s+1}J_{s+1}}\,G^{a}_{J_{s+1}}\Big)$
$\quad+\sum_{c=1}^{r}\sum_{|\beta|\leq s}\bigl(C_{\beta}\bigr)^{a\,J_{s+1}}_{c}\,\partial^{\beta}\!\bigl(G^{c}_{J_{s+1}}\bigr),$
which is exactly the form of (9) for $s+1$. This completes the induction. ∎
4 Local-to-Global Estimates and Structural Theorems
Having established the global integration by parts formula and the local structure of the formal adjoint, we now turn to the fundamental structural theorems of the theory. Our primary objective in this section is to establish a characterization of Sobolev sections on a coordinate chart in terms of their component functions with respect to a local frame (Lemma 5); this allows us to reduce intrinsic analytic problems to their Euclidean counterparts. To achieve this, we require a result ensuring norm equivalence for fiber metrics on vector bundles over compact manifolds. A similar equivalence is well known for the tangent bundle (see, e.g., [7]). This behavior extends to the Sobolev setting, where different choices of metric and connection yield equivalent Sobolev norms on compact manifolds, as we have previously established in [8, Theorem 2] for Sobolev spaces defined via completions.
Lemma 4.
Let $M$ be a smooth compact manifold (with or without boundary), and let $E\to M$ be a smooth vector bundle of finite rank. Let $h$ and $\widetilde{h}$ be two smooth fiber metrics on $E$. Then, there exist constants $c,C>0$ such that for all $p\in M$ and all $v\in E_{p}$,
$c\,h_{p}(v,v)\leq\widetilde{h}_{p}(v,v)\leq C\,h_{p}(v,v).$
Proof.
Let $p\in M$. Consider the unit sphere in the fiber $E_{p}$ with respect to the metric $h$:
$S_{p}:=\{v\in E_{p}\mid h_{p}(v,v)=1\}.$
Since $E_{p}$ is a finite-dimensional vector space, $S_{p}$ is a closed and bounded set, and thus compact. The function $f_{p}:S_{p}\longrightarrow\mathbb{R}$ defined by $f_{p}(v)=\widetilde{h}_{p}(v,v)$ is continuous, so it attains a maximum. Consequently, the function $f:M\longrightarrow\mathbb{R}$ given by
$f(p):=\max_{v\in S_{p}}f_{p}(v)$
is well-defined. Since the bundle and the metrics are smooth, $f$ is continuous on $M$. As $M$ is compact, there exists a point $p_{0}\in M$ such that $f(p)\leq f(p_{0})=:C$ for all $p\in M$.
Moreover,
$C=f(p_{0})=\max_{v\in S_{p_{0}}}\widetilde{h}_{p_{0}}(v,v)>0,$
since $\widetilde{h}$ is positive definite and $v\neq 0$ for all $v\in S_{p_{0}}$. Thus, for all $p\in M$ and all $v\in S_{p}$, we have
$\widetilde{h}_{p}(v,v)\leq C.$
By an analogous argument, considering the minimum, there exists a constant $c>0$ such that for all $p\in M$ and all $v\in S_{p}$,
$c\leq\widetilde{h}_{p}(v,v).$
Now, let $v\in E_{p}\setminus\{0\}$. By homogeneity, the vector $\frac{v}{|v|_{h_{p}}}$ belongs to $S_{p}$ (where $|v|_{h_{p}}=\sqrt{h_{p}(v,v)}$). Therefore,
$c\leq\widetilde{h}_{p}\!\left(\frac{v}{|v|_{h_{p}}},\frac{v}{|v|_{h_{p}}}\right)\leq C.$
Using the bilinearity of the metric, this implies
$c\,|v|^{2}_{h_{p}}\leq\frac{1}{|v|^{2}_{h_{p}}}\widetilde{h}_{p}(v,v)\leq C,$
or equivalently,
$c\,h_{p}(v,v)\leq\widetilde{h}_{p}(v,v)\leq C\,h_{p}(v,v).$
The case $v=0$ holds trivially, as all terms vanish. ∎
With the independence of the fiber metric established for compact domains, we now turn to the local analytical structure of the Sobolev spaces. The following lemma provides the crucial bridge between the intrinsic definition of $W^{m,p}(E)$—defined via weak covariant derivatives—and the standard Sobolev spaces of component functions in local coordinates. This equivalence allows us to import local analytic results, such as approximation theorems and embeddings, from Euclidean space into the manifold setting.
Lemma 5 (Local Norm Equivalence).
Let $(M,g)$ be a Riemannian manifold with Levi–Civita connection $\nabla$. Let $m\geq 1$, let $(U,\phi)$ be a regular coordinate ball in $M$, and let $E\to M$ be a smooth vector bundle of rank $r$ equipped with a fiber metric $h_{E}$ and a compatible connection $\nabla^{E}$. Let $F\in L^{1}_{\operatorname{loc}}(E)$ and let $\{e_{1},\dots,e_{r}\}$ be a local frame for $E$ over $U$, such that $F$ is expressed locally as $F=\displaystyle\sum_{a=1}^{r}F^{a}e_{a}$.
Then,
$F|_{U}\in W^{m,p}(E|_{U})\quad\text{if and only if}\quad F^{a}\circ\phi^{-1}\in W^{m,p}(\phi(U))$
for all $1\leq a\leq r$.
Furthermore, there exist constants $C_{m},c_{m}>0$ (independent of $F$) such that for all $F\in W^{m,p}(E|_{U})$,
$c_{m}\sum_{a=1}^{r}\left\|F^{a}\circ\phi^{-1}\right\|_{W^{m,p}(\phi(U))}\leq\|F\|_{W^{m,p}(E|_{U})}\leq C_{m}\sum_{a=1}^{r}\left\|F^{a}\circ\phi^{-1}\right\|_{W^{m,p}(\phi(U))}.$
Proof.
($\Rightarrow$) First, assume that $F|_{U}\in W^{m,p}(E|_{U})$. Then, for each $s\leq m$, the $s$-th weak covariant derivative of $F|_{U}$ exists; let us denote it by $G\in L^{p}\big(T^{(0,s)}(TU)\otimes E|_{U}\big)$. By definition, for all $\Psi\in\Gamma_{c}\big(T^{(0,s)}(TU)\otimes E|_{U}\big)$, the following holds:
$\int_{U}\langle F,(\nabla^{s})^{*}\Psi\rangle_{g,h_{E}}\,d\lambda_{g}=\int_{U}\langle G,\Psi\rangle_{g,h_{E}}\,d\lambda_{g}.$
In local coordinates, writing $F=F^{a}e_{a}$ and $\Psi=\Psi^{c}_{J_{s}}\,dx^{J_{s}}\otimes e_{c}$, this identity reads:
$\int_{U}h_{bc}\,F^{b}\,\big((\nabla^{s})^{*}\Psi\big)^{c}\,d\lambda_{g}=\int_{U}g^{I_{s}J_{s}}\,h_{bc}\,G^{b}_{I_{s}}\,\Psi^{c}_{J_{s}}\,d\lambda_{g}.$
Fix a target block $T_{s}=(t_{1},\dots,t_{s})$ and a fixed fiber index $a$. Given $\psi\in C_{c}^{\infty}(\phi(U))$, we define a convenient test tensor field $\Psi$ by its components:
$\Psi^{c}_{J_{s}}=\frac{1}{\sqrt{\det(g)}}\,h^{ca}\,g_{J_{s}T_{s}}\,(\psi\circ\phi).$ (17)
By Corollary 3, the formal adjoint acts on $\Psi$ as
$\big((\nabla^{s})^{*}\Psi\big)^{c}=(-1)^{s}\left[\frac{1}{\sqrt{\det(g)}}\partial_{I_{s}}\left(\sqrt{\det(g)}g^{I_{s}J_{s}}\Psi^{c}_{J_{s}}\right)+\sum_{|\beta|\leq s-1}\big(D_{\beta}\big)^{c\,J_{s}}_{d}\partial^{\beta}\Psi^{d}_{J_{s}}\right].$
Substituting (17) into this expression and using the facts that $\frac{\sqrt{\det(g)}}{\sqrt{\det(g)}}=1$ and $g^{I_{s}J_{s}}g_{J_{s}T_{s}}=\delta^{I_{s}}_{T_{s}}$, we obtain
$\big((\nabla^{s})^{*}\Psi\big)^{c}=(-1)^{s}\left[\frac{1}{\sqrt{\det(g)}}\partial_{T_{s}}\left(h^{ca}(\psi\circ\phi)\right)+\sum_{|\beta|\leq s-1}\big(D_{\beta}\big)^{c\,J_{s}}_{d}\partial^{\beta}\left(\frac{h^{da}g_{J_{s}T_{s}}}{\sqrt{\det(g)}}(\psi\circ\phi)\right)\right].$
Let $I_{1}:=\int_{U}\langle F,(\nabla^{s})^{*}\Psi\rangle_{g,h_{E}}\,d\lambda_{g}$ and $I_{2}:=\int_{U}\langle G,\Psi\rangle_{g,h_{E}}\,d\lambda_{g}$. Substituting (17) into $I_{2}$ and noting that $g^{I_{s}J_{s}}g_{J_{s}T_{s}}=\delta^{I_{s}}_{T_{s}}$, we obtain
$I_{2}=\int_{U}\frac{1}{\sqrt{\det(g)}}\delta^{I_{s}}_{T_{s}}\delta_{b}^{a}G^{b}_{I_{s}}(\psi\circ\phi)\,d\lambda_{g}=\int_{\phi(U)}(G^{a}_{T_{s}}\circ\phi^{-1})\psi\,d\lambda_{n}.$ (18)
For $I_{1}$, we use the Leibniz formula on $\partial_{T_{s}}$:
$\partial_{T_{s}}\left(h^{ca}(\psi\circ\phi)\right)=h^{ca}\partial_{T_{s}}(\psi\circ\phi)+\sum_{\alpha<\beta_{T_{s}}}\binom{\beta_{T_{s}}}{\alpha}\partial^{\alpha}(\psi\circ\phi)\partial^{\beta_{T_{s}}-\alpha}\left(h^{ca}\right).$
Substituting back into the integral $I_{1}$, we obtain
$I_{1}$ $=(-1)^{s}\int_{U}\frac{1}{\sqrt{\det(g)}}F^{a}\partial_{T_{s}}(\psi\circ\phi)\,d\lambda_{g}$
$\quad+(-1)^{s}\sum_{\alpha<\beta_{T_{s}}}\int_{U}\frac{1}{\sqrt{\det(g)}}h_{bc}F^{b}\binom{\beta_{T_{s}}}{\alpha}\partial^{\alpha}(\psi\circ\phi)\partial^{\beta_{T_{s}}-\alpha}\left(h^{ca}\right)\,d\lambda_{g}$
$\quad+\sum_{|\beta|\leq s-1}(-1)^{s}\int_{U}h_{bc}F^{b}\big(D_{\beta}\big)^{c\,J_{s}}_{d}\partial^{\beta}\left(\frac{h^{da}g_{J_{s}T_{s}}}{\sqrt{\det(g)}}(\psi\circ\phi)\right)\,d\lambda_{g}.$
We now proceed by induction on $s$. For $s=1$, we have $T_{1}=(t_{1})$, $J_{1}=(j_{1})$, so the expression reduces to
$I_{1}$ $=-\int_{\phi(U)}(F^{a}\circ\phi^{-1})\partial_{t_{1}}\psi\,d\lambda_{n}-\int_{\phi(U)}((h_{bc}F^{b})\circ\phi^{-1})\partial_{t_{1}}\left(h^{ca}\circ\phi^{-1}\right)\psi\,d\lambda_{g}$
$\quad-\int_{\phi(U)}\left(\left(h_{bc}F^{b}(D_{0})^{c\,j_{1}}_{d}h^{da}g_{j_{1}t_{1}}\right)\circ\phi^{-1}\right)\psi\,d\lambda_{g}.$
Define
$H_{t_{1}}^{a}:=h_{bc}F^{b}\partial_{t_{1}}\left(h^{ca}\right)+h_{bc}F^{b}(D_{0})^{c\,j_{1}}_{d}h^{da}g_{j_{1}t_{1}}.$
Since $F|_{U}\in W^{m,p}(E|_{U})$ implies that $F^{b}\circ\phi^{-1}\in L^{p}(\phi(U))$, and the metric coefficients are smooth on the compact set $\overline{\phi(U)}$, it follows that $H_{t_{1}}^{a}\circ\phi^{-1}\in L^{p}(\phi(U))$. Since $I_{1}=I_{2}$, we have
$\int_{\phi(U)}(F^{a}\circ\phi^{-1})\partial_{t_{1}}\psi\,d\lambda_{n}=\int_{\phi(U)}\big((G_{t_{1}}^{a}\circ\phi^{-1})+(H^{a}_{t_{1}}\circ\phi^{-1})\big)\psi\,d\lambda_{n},$
which shows that $F^{a}\circ\phi^{-1}\in W^{1,p}(\phi(U))$.
Now assume the statement holds up to order s1s-1 for $s\leq m$. The inductive hypothesis guarantees that the components of $F$ in $U$ have weak derivatives of order $\leq s-1$, so we can integrate by parts in the weak sense for the lower-order terms.
First, consider the terms coming from the Leibniz expansion in $I_{1}$:
$(-1)^{s}$ $\sum_{\alpha<\beta_{T_{s}}}\int_{U}\frac{h_{bc}F^{b}}{\sqrt{\det(g)}}\binom{\beta_{T_{s}}}{\alpha}\partial^{\alpha}(\psi\circ\phi)\partial^{\beta_{T_{s}}-\alpha}\left(h^{ca}\right)\,d\lambda_{g}$
$=(-1)^{s}\sum_{\alpha<\beta_{T_{s}}}\int_{\phi(U)}\left[\left(h_{bc}F^{b}\binom{\beta_{T_{s}}}{\alpha}\partial^{\beta_{T_{s}}-\alpha}\left(h^{ca}\right)\right)\circ\phi^{-1}\right]\partial^{\alpha}\psi\,d\lambda_{n}.$
Since $|\alpha|<|\beta_{I_{s}}|=s$, we have $|\alpha|\leq s-1$. Integrating by parts (weakly) $|\alpha|$ times, we transfer the derivatives from $\psi$ to the coefficient (which involves $F^{b}$), so the previous term equals
$\sum_{\alpha<\beta_{T_{s}}}(-1)^{s+|\alpha|}\int_{\phi(U)}\partial^{\alpha}\left[\left(h_{bc}F^{b}\binom{\beta_{T_{s}}}{\alpha}\partial^{\beta_{T_{s}}-\alpha}\left(h^{ca}\right)\right)\circ\phi^{-1}\right]\psi\,d\lambda_{n}.$
Second, consider the terms in $I_{1}$ involving the connection coefficients ($D_{\beta}$). Using the same change of variables and recalling that $|\beta|\leq s-1$, we get
$\sum_{|\beta|\leq s-1}(-1)^{s}\int_{U}h_{bc}F^{b}\big(D_{\beta}\big)^{c\,J_{s}}_{d}\partial^{\beta}\left(\frac{h^{da}g_{J_{s}T_{s}}}{\sqrt{\det(g)}}(\psi\circ\phi)\right)\,d\lambda_{g}$
$=(-1)^{s}\sum_{|\beta|\leq s-1}\int_{\phi(U)}\left[\left(h_{bc}F^{b}\big(D_{\beta}\big)^{c\,J_{s}}_{d}\sqrt{\det(g)}\right)\circ\phi^{-1}\right]\partial^{\beta}\left(\left(\frac{h^{da}g_{J_{s}T_{s}}}{\sqrt{\det(g)}}\circ\phi^{-1}\right)\psi\right)\,d\lambda_{n}.$
Integrating by parts $|\beta|$ times yields
$(-1)^{s}\sum_{|\beta|\leq s-1}(-1)^{|\beta|}\int_{\phi(U)}\partial^{\beta}\left[\left(h_{bc}F^{b}\big(D_{\beta}\big)^{c\,J_{s}}_{d}\sqrt{\det(g)}\right)\circ\phi^{-1}\right]\left(\frac{h^{da}g_{J_{s}T_{s}}}{\sqrt{\det(g)}}\circ\phi^{-1}\right)\psi\,d\lambda_{n}.$
We now define $\widetilde{G}^{a}_{T_{s}}$ on $\phi(U)$ by grouping $G^{a}_{T_{s}}\circ\phi^{-1}$ (from $I_{2}$) and moving the terms calculated above from $I_{1}$ to the other side (introducing an extra factor of 1-1, which becomes $(-1)^{s+1}$):
$\widetilde{G}^{a}_{T_{s}}$ $:=G^{a}_{T_{s}}\circ\phi^{-1}+\sum_{\alpha<\beta_{T_{s}}}(-1)^{s+1+|\alpha|}\partial^{\alpha}\left[\left(h_{bc}F^{b}\binom{\beta_{T_{s}}}{\alpha}\partial^{\beta_{T_{s}}-\alpha}\left(h^{ca}\right)\right)\circ\phi^{-1}\right]$
$\quad+\sum_{|\beta|\leq s-1}(-1)^{s+1+|\beta|}\partial^{\beta}\left[\left(h_{bc}F^{b}\big(D_{\beta}\big)^{c\,J_{s}}_{d}\sqrt{\det(g)}\right)\circ\phi^{-1}\right]\left(\frac{h^{da}g_{J_{s}T_{s}}}{\sqrt{\det(g)}}\circ\phi^{-1}\right).$
The inductive hypothesis states that $F^{b}\circ\phi^{-1}\in W^{s-1,p}(\phi(U))$ for all $b\in\{1,\dots,r\}$. Since $\widetilde{G}_{T_{s}}^{a}$ consists of smooth bounded functions (defined on $\overline{\phi(U)}$) and weak derivatives of $F^{b}\circ\phi^{-1}$ of order at most s1s-1, we conclude that $\widetilde{G}_{T_{s}}^{a}\in L^{p}(\phi(U))$. Using the equality $I_{1}=I_{2}$ and the definition of $\widetilde{G}$, we obtain
$(-1)^{s}\int_{\phi(U)}(F^{a}\circ\phi^{-1})\partial_{T_{s}}\psi\,d\lambda_{n}=\int_{\phi(U)}\widetilde{G}_{T_{s}}^{a}\psi\,d\lambda_{n}.$
This shows that $F^{a}\circ\phi^{-1}$ possesses a weak derivative $\partial_{T_{s}}(F^{a}\circ\phi^{-1})=\widetilde{G}_{T_{s}}^{a}\in L^{p}(\phi(U))$. Thus, $F^{a}\circ\phi^{-1}\in W^{s,p}(\phi(U))$. Since $s\leq m$ was arbitrary, we conclude that $F^{a}\circ\phi^{-1}\in W^{m,p}(\phi(U))$ for all $a\in\{1,\dots,r\}$.
($\Leftarrow$) Now assume that $F^{a}\circ\phi^{-1}\in W^{m,p}(\phi(U))$ for all $a\in\{1,\dots,r\}$. Fix $s\leq m$ and let $\Psi\in\Gamma_{c}\big(T^{(0,s)}(TU)\otimes E|_{U}\big)$. We have
$\int_{U}\langle F,(\nabla^{s})^{*}\Psi\rangle_{g,h_{E}}\,d\lambda_{g}=\int_{U}h_{ab}\,F^{a}\big((\nabla^{s})^{*}\Psi\big)^{b}\,d\lambda_{g}.$
By Corollary 3, the adjoint in coordinates takes the form
$\big((\nabla^{s})^{*}\Psi\big)^{b}=(-1)^{s}\left[\frac{1}{\sqrt{\det(g)}}\partial_{I_{s}}\left(\sqrt{\det(g)}g^{I_{s}J_{s}}\Psi^{b}_{J_{s}}\right)+\sum_{|\beta|\leq s-1}\big(D_{\beta}\big)^{b\,J_{s}}_{c}\,\partial^{\beta}\Psi^{c}_{J_{s}}\right].$
Then the integral splits into two terms, $I_{1}$ and $I_{2}$:
$\int_{U}\langle F,(\nabla^{s})^{*}\Psi\rangle_{g,h_{E}}\,d\lambda_{g}$ $=\underbrace{(-1)^{s}\int_{U}h_{ab}F^{a}\frac{1}{\sqrt{\det(g)}}\partial_{I_{s}}\left(\sqrt{\det(g)}g^{I_{s}J_{s}}\Psi^{b}_{J_{s}}\right)\,d\lambda_{g}}_{I_{1}}$
$\quad+\underbrace{(-1)^{s}\sum_{|\beta|\leq s-1}\int_{U}h_{ab}F^{a}\big(D_{\beta}\big)^{b\,J_{s}}_{c}\,\partial^{\beta}\Psi^{c}_{J_{s}}\,d\lambda_{g}}_{I_{2}}.$
Changing the domain of integration to $\phi(U)$ and writing the compositions explicitly:
$I_{1}$ $=(-1)^{s}\int_{\phi(U)}\left[\left(h_{ab}F^{a}\right)\circ\phi^{-1}\right]\partial_{I_{s}}\left(\left(\sqrt{\det(g)}g^{I_{s}J_{s}}\Psi^{b}_{J_{s}}\right)\circ\phi^{-1}\right)\,d\lambda_{n},$
$I_{2}$ $=(-1)^{s}\sum_{|\beta|\leq s-1}\int_{\phi(U)}\left[\left(h_{ab}F^{a}(D_{\beta})^{b\,J_{s}}_{c}\sqrt{\det(g)}\right)\circ\phi^{-1}\right]\partial^{\beta}(\Psi^{c}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}.$
We now integrate by parts weakly in $I_{1}$ (valid since $F^{a}\circ\phi^{-1}\in W^{m,p}(\phi(U))$ and $s\leq m$):
$I_{1}=\int_{\phi(U)}\partial_{I_{s}}\left[\left(h_{ab}\circ\phi^{-1}\right)(F^{a}\circ\phi^{-1})\right]\left(\sqrt{\det(g)}g^{I_{s}J_{s}}\Psi^{b}_{J_{s}}\circ\phi^{-1}\right)\,d\lambda_{n}.$
Applying the Leibniz rule to the bracketed term:
$I_{1}$ $=\int_{\phi(U)}\Bigl[\bigl(h_{ab}g^{I_{s}J_{s}}\sqrt{\det(g)}\circ\phi^{-1}\bigr)\partial_{I_{s}}(F^{a}\circ\phi^{-1})\Bigr](\Psi^{b}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}$
$\quad+\int_{\phi(U)}\Bigg[\sum_{\begin{subarray}{c}I_{s}\\ \alpha<\beta_{I_{s}}\end{subarray}}\binom{\beta_{I_{s}}}{\alpha}\partial^{\beta_{I_{s}}-\alpha}\bigl(h_{ab}\circ\phi^{-1}\bigr)\partial^{\alpha}(F^{a}\circ\phi^{-1})$
$\qquad\qquad\qquad\times\left(\bigl(g^{I_{s}J_{s}}\sqrt{\det(g)}\Psi^{b}_{J_{s}}\bigr)\circ\phi^{-1}\right)\Bigg]\,d\lambda_{n}.$
Define the coefficient functions $(B_{\gamma})_{abI_{s}}$ on $\phi(U)$ for $|\gamma|\leq s-1$ as
$(B_{\gamma})_{ab}^{J_{s}}:=\sum_{\begin{subarray}{c}I_{s}\\ \gamma<\beta_{I_{s}}\end{subarray}}\binom{\beta_{I_{s}}}{\gamma}\partial^{\beta_{I_{s}}-\gamma}\bigl(h_{ab}\circ\phi^{-1}\bigr)\left(\bigl(g^{I_{s}J_{s}}\sqrt{\det(g)}\bigr)\circ\phi^{-1}\right).$
Substituting this definition into $I_{1}$:
$I_{1}$ $=\int_{\phi(U)}\Bigg[\bigl(h_{ab}g^{I_{s}J_{s}}\sqrt{\det(g)}\circ\phi^{-1}\bigr)\partial_{I_{s}}(F^{a}\circ\phi^{-1})$
$\qquad\qquad+\sum_{|\gamma|\leq s-1}(B_{\gamma})_{ab}^{J_{s}}\partial^{\gamma}(F^{a}\circ\phi^{-1})\Bigg](\Psi^{b}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}.$
For $I_{2}$, integrating by parts $|\beta|$ times yields
$I_{2}$ $=\sum_{|\beta|\leq s-1}(-1)^{s+|\beta|}\int_{\phi(U)}\partial^{\beta}\Bigl[\Bigl(\bigl(h_{ab}(D_{\beta})^{b\,J_{s}}_{c}\sqrt{\det(g)}\bigr)\circ\phi^{-1}\Bigr)$
$\qquad\qquad\qquad\qquad\qquad\times(F^{a}\circ\phi^{-1})\Bigr](\Psi^{c}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}.$
Applying Leibniz rule again to expand the derivative $\partial^{\beta}$ of the product:
$I_{2}$ $=\int_{\phi(U)}\sum_{|\beta|\leq s-1}(-1)^{s+|\beta|}\sum_{\gamma\leq\beta}\binom{\beta}{\gamma}\partial^{\beta-\gamma}\Bigl(\bigl(h_{ab}(D_{\beta})^{b\,J_{s}}_{c}\sqrt{\det(g)}\bigr)\circ\phi^{-1}\Bigr)$
$\qquad\qquad\qquad\times\partial^{\gamma}(F^{a}\circ\phi^{-1})(\Psi^{c}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}.$
Define the coefficient functions $(C_{\gamma})^{J_{s}}_{ac}$ on $\phi(U)$ for each $|\gamma|\leq s-1$ as
$(C_{\gamma})^{J_{s}}_{ac}:=\sum_{\begin{subarray}{c}|\beta|\leq s-1\\ \gamma\leq\beta\end{subarray}}(-1)^{s+|\beta|}\binom{\beta}{\gamma}\partial^{\beta-\gamma}\left(\left(h_{ab}(D_{\beta})^{b\,J_{s}}_{c}\sqrt{\det(g)}\right)\circ\phi^{-1}\right).$
Substituting this definition into $I_{2}$:
$I_{2}=\int_{\phi(U)}\sum_{|\gamma|\leq s-1}(C_{\gamma})^{J_{s}}_{ac}\partial^{\gamma}(F^{a}\circ\phi^{-1})(\Psi^{c}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}.$
Combining both results, and renaming the fiber index $b\to c$ in $I_{1}$ (using $h_{ac}$):
$\int_{U}\langle F,(\nabla^{s})^{*}\Psi\rangle\,d\lambda_{g}$ $=\int_{\phi(U)}\Bigg[\left(h_{ac}g^{I_{s}J_{s}}\sqrt{\det(g)}\circ\phi^{-1}\right)\partial_{I_{s}}(F^{a}\circ\phi^{-1})$
$\quad+\sum_{|\gamma|\leq s-1}\left((B_{\gamma})_{ac}^{J_{s}}+(C_{\gamma})^{J_{s}}_{ac}\right)\partial^{\gamma}(F^{a}\circ\phi^{-1})\Bigg](\Psi^{c}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}.$
We now define the correction coefficients $(\widehat{D}_{\gamma})$ on $U$ by pulling back $B$ and $C$ and inverting the metric weights:
$\big(\widehat{D}_{\gamma}\big)^{a}_{b\,I_{s}}:=\frac{g_{I_{s}\tilde{J}_{s}}h^{a\tilde{c}}}{\sqrt{\det(g)}}\Big[\big((B_{\gamma})^{\tilde{J}_{s}}_{b\tilde{c}}\circ\phi\big)+\big((C_{\gamma})^{\tilde{J}_{s}}_{b\tilde{c}}\circ\phi\big)\Big].$
Let $G\in L^{p}(T^{(0,s)}(TM)\otimes E|_{U})$ be given in coordinates by:
$G^{a}_{I_{s}}=\partial_{I_{s}}F^{a}+\sum_{|\gamma|\leq s-1}\big(\widehat{D}_{\gamma}\big)^{a}_{b\,I_{s}}\partial^{\gamma}F^{b}.$ (19)
We verify that $G$ is the required weak derivative:
$\int_{U}\langle G,\Psi\rangle_{g,h_{E}}\,d\lambda_{g}$ $=\int_{U}h_{ac}g^{I_{s}J_{s}}G_{I_{s}}^{a}\Psi_{J_{s}}^{c}\,d\lambda_{g}$
$=\int_{U}h_{ac}g^{I_{s}J_{s}}\left(\partial_{I_{s}}F^{a}+\sum_{|\gamma|\leq s-1}(\widehat{D}_{\gamma})^{a}_{b\,I_{s}}\partial^{\gamma}F^{b}\right)\Psi^{c}_{J_{s}}\,d\lambda_{g}.$
The first term matches the principal part. For the lower-order terms, substituting $\widehat{D}_{\gamma}$:
$\int_{U}\sum_{|\gamma|\leq s-1}h_{ac}g^{I_{s}J_{s}}\left(\frac{g_{I_{s}\tilde{J}_{s}}h^{a\tilde{c}}}{\sqrt{\det(g)}}\Big[\big((B_{\gamma})^{\tilde{J}_{s}}_{b\tilde{c}}\circ\phi\big)+\big((C_{\gamma})^{\tilde{J}_{s}}_{b\tilde{c}}\circ\phi\big)\Big]\right)\partial^{\gamma}F^{b}\Psi^{c}_{J_{s}}\,d\lambda_{g}$
$=\int_{U}\sum_{|\gamma|\leq s-1}\frac{1}{\sqrt{\det(g)}}(h_{ac}h^{a\tilde{c}})(g^{I_{s}J_{s}}g_{I_{s}\tilde{J}_{s}})\Big[\big((B_{\gamma})^{\tilde{J}_{s}}_{b\tilde{c}}\circ\phi\big)+\big((C_{\gamma})^{\tilde{J}_{s}}_{b\tilde{c}}\circ\phi\big)\Big]\partial^{\gamma}F^{b}\Psi^{c}_{J_{s}}\,d\lambda_{g}.$
Using the identities $h_{ac}h^{a\tilde{c}}=\delta^{\tilde{c}}_{c}$ and $g^{I_{s}J_{s}}g_{I_{s}\tilde{J}_{s}}=\delta^{J_{s}}_{\tilde{J}_{s}}$, we obtain
$=\int_{U}\sum_{|\gamma|\leq s-1}\frac{1}{\sqrt{\det(g)}}\Big[\big((B_{\gamma})^{J_{s}}_{bc}\circ\phi\big)+\big((C_{\gamma})^{J_{s}}_{bc}\circ\phi\big)\Big]\partial^{\gamma}F^{b}\Psi^{c}_{J_{s}}\,d\lambda_{g}$
$=\int_{\phi(U)}\sum_{|\gamma|\leq s-1}\left[(B_{\gamma})^{J_{s}}_{bc}+(C_{\gamma})^{J_{s}}_{bc}\right]\partial^{\gamma}(F^{b}\circ\phi^{-1})(\Psi^{c}_{J_{s}}\circ\phi^{-1})\,d\lambda_{n}.$
This expression matches the expansion of $\int_{U}\langle F,(\nabla^{s})^{*}\Psi\rangle\,d\lambda_{g}$ exactly. Therefore, $G$ is the $s$-th weak covariant derivative of $F$ in $U$. Moreover, $G\in L^{p}(T^{(0,s)}(TU)\otimes E|_{U})$ because each coordinate component is a linear combination of weak derivatives of the components of $F$ (up to order $s$) multiplied by smooth, bounded functions. Since $s\leq m$ was arbitrary, $F|_{U}\in W^{m,p}\big(E|_{U}\big)$.
Finally, we establish the norm estimates. The Sobolev norm is defined as
$\|F\|_{W^{m,p}(E|_{U})}:=\left(\sum_{s=0}^{m}\int_{U}|\nabla^{s}_{w}F|_{g,h_{E}}^{p}\,d\lambda_{g}.\right)^{\frac{1}{p}}$
From (19), for any $s\leq m$ and multi-index $I_{s}=(i_{1},\dots,i_{s})$:
$(\nabla^{s}_{w}F)^{a}_{I_{s}}=\partial_{I_{s}}F^{a}+\sum_{|\gamma|\leq s-1}\big(\widehat{D}_{\gamma}\big)^{a}_{b\,I_{s}}\,\partial^{\gamma}F^{b}.$
Define a fiber metric $\overline{h}_{E}$ on $E|_{U}$ by $\overline{h}_{E}(u,v)=\displaystyle\sum_{a=1}^{r}u^{a}v^{a}$ for sections $u=\displaystyle\sum_{a=1}^{r}u^{a}e_{a},v=\sum_{a=1}^{r}v^{a}e_{a}$. Also consider the Riemannian metric $\delta:=\phi^{-1}\overline{g}$ on $\overline{U}$. Since $\overline{U}$ is compact, Lemma 4 provides constants $c_{0},C_{0}>0$ for the pointwise equivalence of the fiber norms, uniform for all $s\in\{0,\dots,m\}$. By integrating these estimates and observing that the Riemann–Lebesgue measures $d\lambda_{g}$ and $d\lambda_{\overline{g}}$ are comparable on compact sets through positive constants [8, Lemma 2], we conclude that there exist constants $c,C>0$ such that the following equivalence of Sobolev norms holds:
$c\|F\|_{W^{m,p}(E|_{U}),\overline{g},\overline{h}_{E}}\leq\|F\|_{W^{m,p}(E|_{U})}\leq C\|F\|_{W^{m,p}(E|_{U}),\overline{g},\overline{h}_{E}}.$ (20)
We express the auxiliary norm as:
$\|F\|_{W^{m,p}(E|_{U}),\overline{g},\overline{h}_{E}}$ $=\Bigg(\sum_{s=0}^{m}\int_{U}\Bigg(\sum_{a=1}^{r}\sum_{I_{s}}\bigg(\partial_{I_{s}}F^{a}+\sum_{|\gamma|\leq s-1}\big(\widehat{D}_{\gamma}\big)^{a}_{b\,I_{s}}\partial^{\gamma}F^{b}\bigg)^{2}\Bigg)^{\frac{p}{2}}d\lambda_{\overline{g}}\Bigg)^{\frac{1}{p}}$
$=\Bigg(\sum_{s=0}^{m}\int_{\phi(U)}\Bigg(\sum_{a=1}^{r}\sum_{I_{s}}\bigg(\partial_{I_{s}}(F^{a}\circ\phi^{-1})$
$\qquad\qquad+\sum_{|\gamma|\leq s-1}\left(\big(\widehat{D}_{\gamma}\big)^{a}_{b\,I_{s}}\circ\phi^{-1}\right)\partial^{\gamma}(F^{b}\circ\phi^{-1})\bigg)^{2}\Bigg)^{\frac{p}{2}}d\lambda_{n}\Bigg)^{\frac{1}{p}}.$
To bound this from above, using $(A+B)^{2}\leq 2A^{2}+2B^{2}$ and defining $K:=\max\|\widehat{D}\|_{\infty}$, we obtain:
$\|F\|_{W^{m,p}(E|_{U}),\overline{g},\overline{h}_{E}}$ $\leq\sqrt{2}\Bigg(\sum_{s=0}^{m}\int_{\phi(U)}\Bigg(\sum_{a=1}^{r}\sum_{|\alpha|=s}\frac{s!}{\alpha!}\left(\partial^{\alpha}(F^{a}\circ\phi^{-1})\right)^{2}$
$\quad+rn^{s}\left(\sum_{b=1}^{r}\sum_{|\gamma|\leq s-1}K\partial^{\gamma}(F^{b}\circ\phi^{-1})\right)^{2}\Bigg)^{\frac{p}{2}}d\lambda_{n}\Bigg)^{\frac{1}{p}}.$
Using $\|\cdot\|_{2}\leq\|\cdot\|_{1}$ to introduce square roots on coefficients, and then applying $(\sum|x_{i}|)^{p}\leq N^{p-1}\sum|x_{i}|^{p}$, all multiplicative constants absorb into a constant $C^{\prime}>0$:
$\leq C^{\prime}\left(\sum_{s=0}^{m}\int_{\phi(U)}\sum_{a=1}^{r}\left(\sum_{|\alpha|=s}\left|\partial^{\alpha}(F^{a}\circ\phi^{-1})\right|^{p}+\sum_{|\gamma|\leq s-1}\left|\partial^{\gamma}(F^{a}\circ\phi^{-1})\right|^{p}\right)d\lambda_{n}\right)^{\frac{1}{p}}$
$\leq C^{\prime\prime}\left(\sum_{a=1}^{r}\sum_{|\beta|\leq m}\int_{\phi(U)}\left|\partial^{\beta}(F^{a}\circ\phi^{-1})\right|^{p}d\lambda_{n}\right)^{\frac{1}{p}}=C^{\prime\prime}\left(\sum_{a=1}^{r}\left\|F^{a}\circ\phi^{-1}\right\|_{W^{m,p}(\phi(U))}^{p}\right)^{\frac{1}{p}}.$
Using $\|\cdot\|_{p}\leq\|\cdot\|_{1}$, we obtain the upper bound:
$\|F\|_{W^{m,p}(E|_{U})}\leq C_{m}\sum_{a=1}^{r}\left\|F^{a}\circ\phi^{-1}\right\|_{W^{m,p}(\phi(U))}.$
For the lower bound, from (19), we isolate the partial derivative:
$\partial_{I_{s}}(F^{a}\circ\phi^{-1})=(\nabla^{s}_{w}F)^{a}_{I_{s}}\circ\phi^{-1}-\sum_{|\gamma|\leq s-1}\left(\big(\widehat{D}_{\gamma}\big)^{a}_{b\,I_{s}}\circ\phi^{-1}\right)\partial^{\gamma}(F^{b}\circ\phi^{-1}).$ (21)
The $L^{p}$ norm of the first term is bounded by the bundle Sobolev norm. Taking $L^{p}(\phi(U))$ norms and summing over components and orders:
$\sum_{a=1}^{r}\|F^{a}\circ\phi^{-1}\|_{W^{m,p}(\phi(U))}\leq\widetilde{C}\left(\|F\|_{W^{m,p}(E|_{U}),\overline{g},\overline{h}_{E}}+\sum_{a=1}^{r}\|F^{a}\circ\phi^{-1}\|_{W^{m-1,p}(\phi(U))}\right).$ (22)
We proceed by induction on $m$. For $m=1$, the $L^{p}$ norms are clearly bounded. Assume the bound holds for m1m-1. Then
$\sum_{a=1}^{r}\|F^{a}\circ\phi^{-1}\|_{W^{m-1,p}(\phi(U))}\leq K_{m-1}\|F\|_{W^{m-1,p}(E|_{U}),\overline{g},\overline{h}_{E}}.$
Substituting this into (22):
$\sum_{a=1}^{r}\|F^{a}\circ\phi^{-1}\|_{W^{m,p}(\phi(U))}\leq\widetilde{C}\left(1+K_{m-1}\right)\|F\|_{W^{m,p}(E|_{U}),\overline{g},\overline{h}_{E}}.$
Combining this with (20), we obtain the lower bound constant $c_{m}>0$. ∎
Since the local characterization in Lemma 5 relies on the Sobolev spaces defined on the restricted bundle $E|_{U}$, we first ensure that the restriction of a global Sobolev section is well-defined and consistent with the local theory:
Proposition 6 (Restriction to open subsets).
Let $(M,g)$ be a Riemannian manifold and $E\to M$ a smooth vector bundle equipped with a fiber metric $h_{E}$. If $U\subseteq M$ is an open subset and $u\in W^{m,p}(E)$, then $u|_{U}\in W^{m,p}(E|_{U})$.
Proof.
Let $s\leq m$. To show that $u|_{U}$ possesses a weak covariant derivative of order $s$ in $L^{p}(T^{(0,s)}(TU)\otimes E|_{U})$, let $\sigma\in\Gamma_{c}(T^{(0,s)}(TU)\otimes E|_{U})$ be an arbitrary test section. We extend $\sigma$ to a section $\widetilde{\sigma}\in\Gamma_{c}(T^{(0,s)}(TM)\otimes E)$ by setting it to zero outside $U$.
Since $u\in W^{m,p}(E)$, it possesses a global $s$-th weak covariant derivative $\nabla_{w}^{s}u$ satisfying:
$\int_{M}\langle\nabla_{w}^{s}u,\widetilde{\sigma}\rangle_{g,h_{E}}d\lambda_{g}=\int_{M}\langle u,(\nabla^{s})^{*}\widetilde{\sigma}\rangle_{h_{E}}d\lambda_{g}.$
Because $(\nabla^{s})^{*}$ is a local operator, the support of $(\nabla^{s})^{*}\widetilde{\sigma}$ is contained in $\operatorname{supp}(\widetilde{\sigma})=\operatorname{supp}(\sigma)\subseteq U$. Furthermore, on the open set $U$, we have $(\nabla^{s})^{*}\widetilde{\sigma}\equiv(\nabla^{s})^{*}\sigma$. It follows that:
$\int_{U}\langle(\nabla_{w}^{s}u)|_{U},\sigma\rangle_{g,h_{E}}d\lambda_{g}$ $=\int_{M}\langle\nabla_{w}^{s}u,\widetilde{\sigma}\rangle_{g,h_{E}}d\lambda_{g}$
$=\int_{M}\langle u,(\nabla^{s})^{*}\widetilde{\sigma}\rangle_{h_{E}}d\lambda_{g}$
$=\int_{U}\langle u|_{U},(\nabla^{s})^{*}\sigma\rangle_{h_{E}}d\lambda_{g}.$
This identity identifies $(\nabla_{w}^{s}u)|_{U}$ as the $s$-th weak covariant derivative of $u|_{U}$. Since $\nabla_{w}^{s}u\in L^{p}$, its restriction remains in $L^{p}(T^{(0,s)}(TU)\otimes E|_{U})$. As $s\leq m$ was arbitrary, we conclude that $u|_{U}\in W^{m,p}(E|_{U})$. ∎
Another fundamental property is that the weak covariant derivative is a local operator:
Proposition 7 (Locality of the Weak Covariant Derivative).
Let $(M,g)$ be a Riemannian manifold and let $E\to M$ be a smooth vector bundle with fiber metric $h_{E}$ and a connection $\nabla^{E}$. Let $u\in L^{1}_{\operatorname{loc}}(E)$ be a section possessing a weak $s$-th covariant derivative $\nabla^{s}_{w}u$. Let $U\subseteq M$ be an open set. If $u=0$ a.e. in $U$, then $\nabla^{s}_{w}u=0$ a.e. in $U$.
Proof.
Let $\sigma\in\Gamma_{c}(T^{(0,s)}(TU)\otimes E|_{U})$ be an arbitrary test section with support in $U$. We extend $\sigma$ by zero to the whole manifold $M$, denoting the extension by $\widetilde{\sigma}$. Since $\operatorname{supp}(\sigma)\subset U$, $\widetilde{\sigma}$ is smooth and compactly supported on $M$. By the definition of the weak derivative,
$\int_{M}\langle\nabla_{w}^{s}u,\widetilde{\sigma}\rangle_{g,h_{E}}d\lambda_{g}=\int_{M}\langle u,(\nabla^{s})^{*}\widetilde{\sigma}\rangle_{h_{E}}d\lambda_{g}.$
Since $(\nabla^{s})^{*}$ is a local operator (being a differential operator with smooth coefficients), $\operatorname{supp}((\nabla^{s})^{*}\widetilde{\sigma})\subseteq\operatorname{supp}(\widetilde{\sigma})\subset U$. Since $u$ vanishes almost everywhere on $U$, the right-hand side is zero. Thus,
$\int_{U}\langle\nabla_{w}^{s}u,\sigma\rangle_{g,h_{E}}d\lambda_{g}=0$
for all $\sigma\in\Gamma_{c}(T^{(0,s)}(TU)\otimes E|_{U})$. By the fundamental lemma of the calculus of variations extended to vector bundles (density of test sections in $L^{1}_{loc}$), we conclude that $\nabla_{w}^{s}u=0$ almost everywhere in $U$. ∎
5 Global Approximation Results
We are now in a position to prove the main approximation result of this paper. The following theorem generalizes the classical Meyers-Serrin theorem to the context of vector bundles over Riemannian manifolds, stating that smooth sections are dense in the Sobolev space $W^{m,p}(E)$.
Theorem 8 (Meyers–Serrin for Vector Bundles).
Let $(M,g)$ be a Riemannian manifold and let $\pi_{E}:E\to M$ be a smooth vector bundle equipped with a fiber metric $h_{E}$ and a compatible connection $\nabla^{E}$. Then
$H^{m,p}(E)=W^{m,p}(E).$
Proof.
Since $M$ is a Lindelöf space, it admits a countable open cover by regular coordinate balls $(U_{\alpha},\phi_{\alpha})_{\alpha\in\mathbb{N}}$. Let $(\psi_{\alpha})_{\alpha\in\mathbb{N}}$ be a smooth partition of unity subordinate to this cover.
Consider a section $F\in W^{m,p}(E)$. We can decompose it as
$F=\sum_{\alpha=1}^{\infty}\psi_{\alpha}F,$
where the sum is locally finite. For each $\alpha$, by Proposition 6, $\psi_{\alpha}F|_{U_{\alpha}}$ belongs to $W^{m,p}(E|_{U_{\alpha}})$, since $\psi_{\alpha}$ is smooth and has compact support contained in $U_{\alpha}$.
Fix $\alpha$ and work in the chart $(U_{\alpha},\phi_{\alpha})$, equipped with a local frame $(e_{1},\dots,e_{r})$ of $E$ over $U_{\alpha}$. Applying Lemma 5 to the section $(\psi_{\alpha}F)|_{U_{\alpha}}$, we deduce that for every fiber index $a\in\{1,\dots,r\}$,
$\big(\psi_{\alpha}F\big)^{a}\circ\phi_{\alpha}^{-1}\in W^{m,p}(\phi_{\alpha}(U_{\alpha})).$
Moreover, there exist constants $C_{m,\alpha}>0$ such that, for any section $H\in W^{m,p}(E|_{U_{\alpha}})$,
$\|H\|_{W^{m,p}(E|_{U_{\alpha}})}\leq C_{m,\alpha}\sum_{a=1}^{r}\big\|H^{a}\circ\phi_{\alpha}^{-1}\big\|_{W^{m,p}(\phi_{\alpha}(U_{\alpha}))}.$
Let $\Omega_{\alpha}:=\phi_{\alpha}(U_{\alpha})\subset\mathbb{R}^{n}$. By the classical Meyers–Serrin theorem on domains in $\mathbb{R}^{n}$, for each $\alpha\in\mathbb{N}$ and each component $a\in\{1,\dots,r\}$, there exists a sequence
$(\widetilde{G}_{\alpha,k})^{a}\in C_{c}^{\infty}(\Omega_{\alpha}),\qquad k\in\mathbb{N},$
such that
$\sum_{a=1}^{r}\left\|(\widetilde{G}_{\alpha,k})^{a}-\big(\psi_{\alpha}F\big)^{a}\circ\phi_{\alpha}^{-1}\right\|_{W^{m,p}(\Omega_{\alpha})}\xrightarrow[k\to\infty]{}0.$
Using these smooth functions, we define a smooth section $G_{\alpha,k}\in\Gamma(E|_{U_{\alpha}})$ by
$G_{\alpha,k}=\sum_{a=1}^{r}\Big[(\widetilde{G}_{\alpha,k})^{a}\circ\phi_{\alpha}\Big]\;e_{a}.$
Then, by the norm estimate from Lemma 5, it follows that
$\|G_{\alpha,k}-(\psi_{\alpha}F)|_{U_{\alpha}}\|_{W^{m,p}(E|_{U_{\alpha}})}\xrightarrow[k\to\infty]{}0.$
Extending each $G_{\alpha,k}$ by zero outside $U_{\alpha}$, we may view it as a global section with compact support, i.e., $G_{\alpha,k}\in\Gamma_{c}(E)$. Furthermore, the norm on the restriction is given by
$\|G_{\alpha,k}-(\psi_{\alpha}F)|_{U_{\alpha}}\|_{W^{m,p}(E|_{U_{\alpha}})}=\left(\sum_{s=0}^{m}\int_{U_{\alpha}}|\nabla_{w}^{s}(G_{\alpha,k}-(\psi_{\alpha}F)|_{U_{\alpha}})|^{p}_{g,h_{E}}d\lambda_{g}\right)^{\frac{1}{p}}.$
Since $\operatorname{supp}(G_{\alpha,k}-\psi_{\alpha}F)\subseteq U_{\alpha}$ and $\nabla^{s}_{w}$ is a local operator (Proposition 7), the integral over $U_{\alpha}$ coincides with the global integral:
$\left(\sum_{s=0}^{m}\int_{M}|\nabla_{w}^{s}(G_{\alpha,k}-\psi_{\alpha}F)|^{p}_{g,h_{E}}d\lambda_{g}\right)^{\frac{1}{p}}.$
Consequently,
$\|G_{\alpha,k}-\psi_{\alpha}F\|_{W^{m,p}(E)}\xrightarrow[k\to\infty]{}0.$
Now, fix $\alpha\in\mathbb{N}$ and an integer $q\in\mathbb{N}$. By the convergence established above, there exists an index $k_{\alpha,q}\in\mathbb{N}$ such that
$\|\psi_{\alpha}F-G_{\alpha,k_{\alpha,q}}\|_{W^{m,p}(E|_{U_{\alpha}})}<\frac{1}{2^{\alpha+q}}.$
Since $\operatorname{supp}(G_{\alpha,k})\subseteq U_{\alpha}$ for all $\alpha,k\in\mathbb{N}$ and the cover $(U_{\alpha})_{\alpha\in\mathbb{N}}$ is locally finite, the sum
$G_{q}:=\sum_{\alpha=1}^{\infty}G_{\alpha,k_{\alpha,q}}$
is locally finite and therefore defines a global smooth section $G_{q}\in\Gamma(E)$.
We now show that $G_{q}$ approximates $F$ in the Sobolev norm. Using the triangle inequality:
$\|F-G_{q}\|_{W^{m,p}(E)}\leq\sum_{\alpha=1}^{\infty}\|\psi_{\alpha}F-G_{\alpha,k_{\alpha,q}}\|_{W^{m,p}(E)}\leq\sum_{\alpha=1}^{\infty}\frac{1}{2^{\alpha+q}}=\frac{1}{2^{q}}.$
In particular,
$\|G_{q}\|_{W^{m,p}(E)}\leq\|F\|_{W^{m,p}(E)}+\|F-G_{q}\|_{W^{m,p}(E)}\leq\|F\|_{W^{m,p}(E)}+2^{-q}<\infty,$
which implies that $G_{q}\in W^{m,p}(E)\cap\Gamma(E)$ for all $q\in\mathbb{N}$.
Since $2^{-q}\to 0$ as $q\to\infty$, we conclude that $G_{q}\to F$ in $W^{m,p}(E)$. Thus, every section in $W^{m,p}(E)$ can be approximated in the Sobolev norm by smooth sections, proving that
$H^{m,p}(E)=W^{m,p}(E).$
Having established the equivalence between the completion-based spaces and those defined via weak derivatives, we conclude this section by addressing the geometric robustness of the theory. The following theorem ensures that, on compact manifolds, the Sobolev spaces $W^{m,p}(E)$ are intrinsic to the smooth structure of the bundle and the manifold, remaining invariant under changes of the underlying metrics and connection.
Theorem 9 (Geometric Invariance on Compact Manifolds).
Let $M$ be a compact Riemannian manifold and $E\to M$ a smooth vector bundle. The Sobolev spaces $W^{m,p}(E)$ are independent of the choice of the Riemannian metric $g$, the fiber metric $h_{E}$, and the compatible connection $\nabla^{E}$.
Proof.
By Theorem 8, $W^{m,p}(E)$ coincides with the completion-based space $H^{m,p}(E)$. The independence of $H^{m,p}(E)$ from the underlying geometric data on compact manifolds follows by generalizing the analytic arguments established for the scalar-valued case in [8, Theorem 2] to the vector bundle setting.∎
6 Sobolev Embeddings and the Rellich–Kondrashov Theorem
Sobolev embeddings are of paramount importance when addressing partial differential equations. Just as in the case of scalar-valued functions, it is natural to inquire about the validity of these embeddings within the more general framework we have defined for smooth vector bundles. It turns out that if $M$ is a compact manifold, the classical results remain valid in this setting, as we demonstrate in this section.
Theorem 10 (Sobolev Embedding for Vector Bundles).
Let $(M,g)$ be a compact Riemannian manifold of dimension $n$, and let $E\longrightarrow M$ be a smooth vector bundle equipped with a fiber metric $h_{E}$ and a compatible connection $\nabla^{E}$. Then, for any integers $j\geq 0$ and $1\leq m<n$, and real numbers $p\in[1,n/m)$ and $q\in[1,p^{*}_{m}]$, we have
$W^{j+m,p}(E)\subseteq W^{j,q}(E),$
and this inclusion is continuous. Here $p^{*}_{m}=np/(n-mp)$ denotes the critical Sobolev exponent.
Proof.
Let $q\in[1,p_{m}^{*}]$ be arbitrary. Consider a finite open cover $(U_{\alpha},\phi_{\alpha})_{\alpha=1}^{N}$ of $M$ consisting of regular coordinate balls, each being the domain of a local frame $(e_{1,\alpha},\dots,e_{r,\alpha})$, and let $(\psi_{\alpha})_{\alpha=1}^{N}$ be a partition of unity subordinate to this cover. By the classical Sobolev embedding theorem in Euclidean domains, for each $\alpha\in\{1,\dots,N\}$, we have the continuous inclusion:
$W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))\subseteq W^{j,q}(\phi_{\alpha}(U_{\alpha})).$ (23)
Thus, there exists $C_{\alpha}>0$ such that $\|\cdot\|_{W^{j,q}(\phi_{\alpha}(U_{\alpha}))}\leq C_{\alpha}\|\cdot\|_{W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))}$.
Let $F\in W^{j+m,p}(E)$ and assume that $F=F^{a}_{\alpha}e_{a,\alpha}$ in $U_{\alpha}$. Note that by Proposition 6, $(\psi_{\alpha}F)|_{U_{\alpha}}\in W^{j+m,p}(E|_{U_{\alpha}})$. By Lemma 5, we know that for all $a\in\{1,\dots,r\}$ and all $\alpha\in\{1,\dots,N\}$, the function $\psi_{\alpha}F^{a}_{\alpha}\circ\phi_{\alpha}^{-1}$ belongs to $W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))$. Consequently, by (23), it also belongs to $W^{j,q}(\phi_{\alpha}(U_{\alpha}))$ and satisfies
$\|\psi_{\alpha}F^{a}_{\alpha}\circ\phi_{\alpha}^{-1}\|_{W^{j,q}(\phi_{\alpha}(U_{\alpha}))}\leq C_{\alpha}\|\psi_{\alpha}F^{a}_{\alpha}\circ\phi_{\alpha}^{-1}\|_{W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))}.$
Applying Lemma 5 again, since the component functions are in $W^{j,q}$, we conclude that $(\psi_{\alpha}F)|_{U_{\alpha}}\in W^{j,q}(E|_{U_{\alpha}})$. Furthermore, there exist constants $K_{\alpha,j,m},K_{\alpha,j}>0$ such that
$\|\psi_{\alpha}F\|_{W^{j,q}(E|_{U_{\alpha}})}\leq K_{\alpha,j}\sum_{a=1}^{r}\|\psi_{\alpha}F^{a}_{\alpha}\circ\phi^{-1}_{\alpha}\|_{W^{j,q}(\phi_{\alpha}(U_{\alpha}))},$
and
$K_{\alpha,j,m}\sum_{a=1}^{r}\|\psi_{\alpha}F^{a}_{\alpha}\circ\phi_{\alpha}^{-1}\|_{W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))}\leq\|\psi_{\alpha}F\|_{W^{j+m,p}(E|_{U_{\alpha}})}.$
Now, for each $s\leq m$, $(\psi_{\alpha}F)|_{U_{\alpha}}$ possesses a weak covariant derivative $\nabla_{w}^{s}(\psi_{\alpha}F)|_{U_{\alpha}}\in L^{q}(T^{(0,s)}(TU_{\alpha})\otimes E|_{U_{\alpha}})$. For each $\alpha$, we extend this derivative by zero outside $U_{\alpha}$ and denote the extension by $\nabla^{s}_{w}(\psi_{\alpha}F)$. By Proposition 7 (locality), $\operatorname{supp}(\nabla_{w}^{s}(\psi_{\alpha}F))\subseteq\operatorname{supp}((\psi_{\alpha}F)|_{U_{\alpha}})\subseteq U_{\alpha}$.
We claim that $\nabla^{s}_{w}(\psi_{\alpha}F)\in L^{q}(T^{(0,s)}(TM)\otimes E)$ is indeed the global weak derivative of $\psi_{\alpha}F$. Let $\Psi\in\Gamma_{c}(T^{(0,s)}(TM)\otimes E)$. Since $\operatorname{supp}(\psi_{\alpha}F)$ is a closed set contained in $U_{\alpha}$, there exists an open set $V_{\alpha}$ such that $\operatorname{supp}(\psi_{\alpha}F)\subseteq V_{\alpha}\subseteq\overline{V_{\alpha}}\subseteq U_{\alpha}$. Let $f_{\alpha}\in C^{\infty}(M)$ be a bump function such that $f_{\alpha}\equiv 1$ on $\overline{V_{\alpha}}$ and $\operatorname{supp}(f_{\alpha})\subseteq U_{\alpha}$. Define $\widetilde{\Psi}:=\Psi|_{U_{\alpha}}f_{\alpha}$. Note that $\operatorname{supp}(\widetilde{\Psi})\subseteq U_{\alpha}$ and $\widetilde{\Psi}\in\Gamma_{c}(T^{(0,s)}(TU_{\alpha})\otimes E|_{U_{\alpha}})$. Then
$\int_{M}\langle\nabla_{w}^{s}(\psi_{\alpha}F),\Psi\rangle_{g,h_{E}}d\lambda_{g}$ $=\int_{U_{\alpha}}\langle\nabla_{w}^{s}(\psi_{\alpha}F),\widetilde{\Psi}\rangle_{g,h_{E}}d\lambda_{g}$
$=\int_{U_{\alpha}}\langle\psi_{\alpha}F,(\nabla^{s})^{*}\widetilde{\Psi}\rangle_{h_{E}}d\lambda_{g}$
$=\int_{M}\langle\psi_{\alpha}F,(\nabla^{s})^{*}\Psi\rangle_{h_{E}}d\lambda_{g},$
where we used the fact that $\Psi\equiv\widetilde{\Psi}$ on $V_{\alpha}$ and the locality of $(\nabla^{s})^{*}$ implies $(\nabla^{s})^{*}\Psi\equiv(\nabla^{s})^{*}\widetilde{\Psi}$ on $V_{\alpha}$ (since $\psi_{\alpha}F$ vanishes outside $V_{\alpha}$). Thus, $\psi_{\alpha}F\in W^{j,q}(E)$. Since $F=\displaystyle\sum_{\alpha=1}^{N}\psi_{\alpha}F$, it follows that $F\in W^{j,q}(E)$.
Finally, we estimate the norm:
$\|F\|_{W^{j,q}(E)}$ $=\left\|\sum_{\alpha=1}^{N}\psi_{\alpha}F\right\|_{W^{j,q}(E)}\leq\sum_{\alpha=1}^{N}\|\psi_{\alpha}F\|_{W^{j,q}(E)}=\sum_{\alpha=1}^{N}\|\psi_{\alpha}F\|_{W^{j,q}(E|_{U_{\alpha}})}$
$\leq\sum_{\alpha=1}^{N}\sum_{a=1}^{r}K_{\alpha,j}\|\psi_{\alpha}F^{a}_{\alpha}\circ\phi_{\alpha}^{-1}\|_{W^{j,q}(\phi_{\alpha}(U_{\alpha}))}$
$\leq\sum_{\alpha=1}^{N}\sum_{a=1}^{r}K_{\alpha,j}C_{\alpha}\|\psi_{\alpha}F^{a}_{\alpha}\circ\phi_{\alpha}^{-1}\|_{W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))}.$
Defining $K:=\displaystyle\max_{1\leq\alpha\leq N}\{K_{\alpha,j},C_{\alpha}\}$ and using the lower bound for the local Sobolev norm:
$\|F\|_{W^{j,q}(E)}\leq K\sum_{\alpha=1}^{N}\frac{1}{K_{\alpha,j,m}}\|\psi_{\alpha}F\|_{W^{j+m,p}(E|_{U_{\alpha}})}\leq C\sum_{\alpha=1}^{N}\|\psi_{\alpha}F\|_{W^{j+m,p}(E)},$
where $C$ is a constant independent of $F$.
To conclude the continuity, we use the density of smooth sections (Theorem 8). Suppose $F\in\Gamma(E)\cap W^{j+m,p}(E)$. By the Leibniz rule for the tensor product connection:
$|\nabla^{s}(\psi_{\alpha}F)|_{g,h_{E}}\leq\sum_{t=0}^{s}\binom{s}{t}|\nabla^{s-t}\psi_{\alpha}|_{g}|\nabla^{t}F|_{g,h_{E}}.$
Taking $L^{p}$ norms, there exists $C^{\prime}>0$ such that
$\|\psi_{\alpha}F\|_{W^{j+m,p}(E)}\leq C^{\prime}\|F\|_{W^{j+m,p}(E)},\qquad\forall\alpha\in\{1,\dots,N\}.$ (24)
Thus,
$\|F\|_{W^{j,q}(E)}\leq NCC^{\prime}\|F\|_{W^{j+m,p}(E)}.$
Extending this result by density, we conclude that the inclusion is continuous. ∎
We also state and prove the Rellich–Kondrashov compactness theorem for vector bundles.
Theorem 11 (Rellich–Kondrashov for Vector Bundles).
Let $(M,g)$ be a compact Riemannian manifold of dimension $n$, and let $E\longrightarrow M$ be a smooth vector bundle with fiber metric $h_{E}$ and compatible connection $\nabla^{E}$. Then, for any integers $j\geq 0$ and $1\leq m<n$, and real numbers $p\in[1,n/m)$ and $q\in[1,p^{*}_{m})$, the inclusion
$W^{j+m,p}(E)\subseteq W^{j,q}(E)$
is compact.
Proof.
Let $q\in[1,p_{m}^{*})$ be arbitrary. Consider the same finite cover $(U_{\alpha},\phi_{\alpha})_{\alpha=1}^{N}$ and partition of unity $(\psi_{\alpha})_{\alpha=1}^{N}$ as before. By Theorem 10, we have the continuous embedding $W^{j+m,p}(E)\subseteq W^{j,q}(E)$.
Let $(F_{k})_{k\in\mathbb{N}}\subseteq W^{j+m,p}(E)$ be a bounded sequence; that is, there exists $C_{0}>0$ such that $\|F_{k}\|_{W^{j+m,p}(E)}\leq C_{0}$ for all $k$.
For each $k\in\mathbb{N}$ and each $\alpha\in\{1,\dots,N\}$, we have $(\psi_{\alpha}F_{k})|_{U_{\alpha}}\in W^{j+m,p}(E|_{U_{\alpha}})$ and $(\psi_{\alpha}F_{k})|_{U_{\alpha}}\in W^{j,q}(E|_{U_{\alpha}})$. By Lemma 5, there exist constants $K_{\alpha,j,m}>0$ and $K_{\alpha,j}>0$ such that the following norm estimates hold:
$\begin{split}K_{\alpha,j,m}\sum_{a=1}^{r}\|\psi_{\alpha}(F_{k})_{\alpha}^{a}\circ\phi_{\alpha}^{-1}\|_{W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))}&\leq\|\psi_{\alpha}F_{k}\|_{W^{j+m,p}(E|_{U_{\alpha}})},\\ \|\psi_{\alpha}F_{k}\|_{W^{j,q}(E|_{U_{\alpha}})}&\leq K_{\alpha,j}\sum_{a=1}^{r}\|\psi_{\alpha}(F_{k})_{\alpha}^{a}\circ\phi_{\alpha}^{-1}\|_{W^{j,q}(\phi_{\alpha}(U_{\alpha}))}.\end{split}$ (25)
By combining the first inequality of (25) with the estimate (24), it follows that for each $\alpha\in\{1,\dots,N\}$ and each component index $a\in\{1,\dots,r\}$, the sequence of functions $\left(\psi_{\alpha}(F_{k})_{\alpha}^{a}\circ\phi_{\alpha}^{-1}\right)_{k\in\mathbb{N}}$ is bounded in $W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))$:
$\|\psi_{\alpha}(F_{k})_{\alpha}^{a}\circ\phi_{\alpha}^{-1}\|_{W^{j+m,p}(\phi_{\alpha}(U_{\alpha}))}\leq\frac{1}{K_{\alpha,j,m}}\|\psi_{\alpha}F_{k}\|_{W^{j+m,p}(E)}\leq\frac{C^{\prime}C_{0}}{K_{\alpha,j,m}}.$
By the classical Rellich–Kondrachov theorem for Euclidean domains, for each chart index $\alpha\in\{1,\dots,N\}$ and each fiber component index $a\in\{1,\dots,r\}$, the sequence $\left(\psi_{\alpha}(F_{k})_{\alpha}^{a}\circ\phi_{\alpha}^{-1}\right)_{k\in\mathbb{N}}$ admits a subsequence converging in $W^{j,q}(\phi_{\alpha}(U_{\alpha}))$. Since there are finitely many charts $\alpha\in\{1,\dots,N\}$ and finitely many fiber components $a\in\{1,\dots,r\}$, we may perform successive subsequence extractions. More precisely, by iteratively refining the sequence $(F_{k})$ a finite number of times, we obtain a subsequence (still denoted by $(F_{k})$) such that, for each chart $\alpha\in\{1,\dots,N\}$ and each fiber component $a\in\{1,\dots,r\}$, the sequence $\psi_{\alpha}(F_{k})_{\alpha}^{a}\circ\phi_{\alpha}^{-1}$ converges —and hence is Cauchy— in $W^{j,q}(\phi_{\alpha}(U_{\alpha}))$.
Let $\varepsilon>0$. Since the subsequence is Cauchy component-wise, there exists $N^{\prime}\in\mathbb{N}$ such that for all $k,l\geq N^{\prime}$,
$\sum_{a=1}^{r}\|\psi_{\alpha}(F_{k}-F_{l})_{\alpha}^{a}\circ\phi^{-1}_{\alpha}\|_{W^{j,q}(\phi_{\alpha}(U_{\alpha}))}<\frac{\varepsilon}{NK_{\alpha,j}},\qquad\forall\alpha\in\{1,\dots,N\}.$
Then, estimating the global norm:
$\|F_{k}-F_{l}\|_{W^{j,q}(E)}$ $=\left\|\sum_{\alpha=1}^{N}\psi_{\alpha}(F_{k}-F_{l})\right\|_{W^{j,q}(E)}\leq\sum_{\alpha=1}^{N}\|\psi_{\alpha}(F_{k}-F_{l})\|_{W^{j,q}(E)}$
$\leq\sum_{\alpha=1}^{N}K_{j,\alpha}\sum_{a=1}^{r}\|\psi_{\alpha}(F_{k}-F_{l})_{\alpha}^{a}\circ\phi_{\alpha}^{-1}\|_{W^{j,q}(\phi_{\alpha}(U_{\alpha}))}$
$<\sum_{\alpha=1}^{N}\frac{\varepsilon}{N}=\varepsilon.$
Thus, $(F_{k})_{k\in\mathbb{N}}$ is a Cauchy sequence in the Banach space $W^{j,q}(E)$, and therefore converges. We conclude that the embedding is compact. ∎
The results presented in this work establish the foundational properties of Sobolev spaces of sections of vector bundles over Riemannian manifolds. By bridging the gap between intrinsic geometric definitions and local analytical estimates, we have extended the classical Meyers–Serrin and Rellich–Kondrashov theorems to this more general setting. Finally, we must note that the global integration by parts formula derived in Section 3 (Theorem 2) remains valid for manifolds with boundary, provided the corresponding boundary integrals are taken into account. While the treatment of such boundary cases and the associated trace theorems lies beyond the immediate scope of this paper, the framework developed herein provides the necessary tools for addressing these problems in a purely intrinsic manner in future research.
7 Applications: Green’s Formula and Norm Equivalence
In the Euclidean setting, it is well known that the integration by parts formula is a fundamental pillar of Sobolev space theory and nonlinear analysis, particularly in the study of partial differential equations (PDEs). A classic application of this formula is establishing the equivalence between the standard Sobolev norm and a structured norm based on powers of the Laplacian. For a bounded domain $\Omega\subset\mathbb{R}^{n}$ and any function $u\in H_{0}^{m}(\Omega)$, the standard Sobolev norm is strictly equal to
$\|u\|_{H^{m}(\Omega)}^{2}=\displaystyle\sum_{j=0}^{m}\displaystyle\int_{\Omega}|\nabla^{j}u|^{2}\,d\lambda_{n}=\displaystyle\sum_{\begin{subarray}{c}j=0\\ j\text{ even}\end{subarray}}^{m}\displaystyle\int_{\Omega}|\Delta^{\frac{j}{2}}u|^{2}\,d\lambda_{n}+\displaystyle\sum_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{m}\displaystyle\int_{\Omega}\big|\nabla\Delta^{\frac{j-1}{2}}u\big|^{2}\,d\lambda_{n},$ (26)
where $\Delta=\displaystyle\sum_{i=1}^{n}\partial_{i}^{2}$ is the standard Euclidean Laplacian. While the counterpart of this equivalence on Riemannian manifolds is widely accepted and heavily used in the geometric PDE literature, see for example [9],[10],[11], an explicit, purely geometric proof is surprisingly scarce. In this section, we utilize the integration by parts formula (Theorem 2) to generalize Green’s identity and subsequently provide a rigorous proof of this norm equivalence on vector bundles with a closed manifold as the base space, explicitly accounting for the intrinsic curvature. Finally, as a Corollary of this result, we show the norm equivalence in closed manifolds, which is a well known result in the literature but lacks any rigorous proof.
7.1 Green’s Formula for the Bochner Laplacian
An immediate consequence of our main theorem is the standard Green’s formula in the context of vector bundles.
Corollary 12 (Green’s Formula for the Bochner Laplacian).
Let $(M,g)$ be a Riemannian manifold with boundary $\partial M$, and let $\pi_{E}:E\to M$ be a smooth vector bundle equipped with a fiber metric $h_{E}$ and a compatible connection $\nabla^{E}$. Let $\nu$ denote the outward unit normal to $\partial M$.
Then, for any $u,v\in\Gamma_{c}(E)$, the following identity holds:
$\int_{M}\langle\nabla^{E}u,\nabla^{E}v\rangle_{g,h_{E}}\,d\lambda_{g}=\displaystyle\int_{M}\langle u,\Delta_{B}v\rangle_{h_{E}}\,d\lambda_{g}+\displaystyle\int_{\partial M}\big\langle u,\nabla^{E}_{\nu}v\big\rangle_{h_{E}}\,d\lambda_{\widetilde{g}}$,
where the Bochner Laplacian is defined by $\Delta_{B}:=\nabla^{E*}\nabla^{E}$. In particular, if $\partial M=\varnothing$, then
$\int_{M}\langle\nabla^{E}u,\nabla^{E}v\rangle_{g,h_{E}}\,d\lambda_{g}=\displaystyle\int_{M}\langle u,\Delta_{B}v\rangle_{h_{E}}\,d\lambda_{g}.$
Proof.
We apply Theorem 2 with $s=1$ and $k=l=0$, setting
$F=u\in\Gamma_{c}(E),\quad\text{and}\quad G=\nabla^{E}v\in\Gamma_{c}(T^{*}M\otimes E).$
The identity from the theorem then takes the form:
$\int_{M}\langle\nabla^{E}u,\nabla^{E}v\rangle_{g,h_{E}}\,d\lambda_{g}=-\displaystyle\int_{M}\big\langle u,(\operatorname{tr}_{g}\circ\nabla)(\nabla^{E}v)\big\rangle_{h_{E}}\,d\lambda_{g}+\displaystyle\int_{\partial M}\Big\langle u,\,\iota_{\nu}(\nabla^{E}v)\Big\rangle_{h_{E}}\,d\lambda_{\widetilde{g}}$. (27)
Under the sign convention adopted for the Bochner Laplacian, we have
$\Delta_{B}=-(\operatorname{tr}_{g}\circ\nabla)\circ\nabla^{E},$
and therefore, the interior term can be rewritten as
$\big\langle u,(\operatorname{tr}_{g}\circ\nabla)(\nabla^{E}v)\big\rangle_{h_{E}}=\langle u,\Delta_{B}v\rangle_{h_{E}}.$
It remains to identify the boundary term. In local coordinates $(x^{1},\dots,x^{n})$, we write the normal vector as $\nu=\nu^{i}\partial_{i}$. By the definition of the interior multiplication $\iota_{\nu}$ (acting on the last covariant index), we have $\iota_{\nu}(\nabla^{E}v)=\nu^{i}(\nabla^{E}v)_{i}$. Furthermore, since $(\nabla^{E}v)_{i}=\nabla^{E}_{\partial_{i}}v$, we obtain $\iota_{\nu}(\nabla^{E}v)=\nu^{i}\nabla^{E}_{\partial_{i}}v$. Finally, using the $C^{\infty}(M)$-linearity of the connection with respect to the vector field, it follows that
$\nabla^{E}_{\nu}v=\nabla^{E}_{\nu^{i}\partial_{i}}v=\nu^{i}\nabla^{E}_{\partial_{i}}v,$
which implies $\iota_{\nu}(\nabla^{E}v)=\nabla^{E}_{\nu}v$. Substituting these results into (27) yields the desired formula. ∎
7.2 Technical Preliminaries and the Commutator Lemma
To systematically handle the complexity of tensor contractions and curvature terms when derivatives are commuted, we adopt the standard * notation.
Notation 1.
Let $E_{1},\dots,E_{r},H\to M$ be smooth vector bundles equipped with fiber metrics. For sections $u_{i}\in\Gamma(E_{i})$, the expression $T=u_{1}*\dots*u_{r}\in\Gamma(H)$ denotes any section obtained from the tensor product $u_{1}\otimes\dots\otimes u_{r}$ via a finite linear combination of permutations, dualities, and contractions induced by the metrics.
A standard consequence of this definition is that on any compact subset $K\subset M$, there exists a constant $C_{K}>0$ such that $|u_{1}*\dots*u_{r}|\leq C_{K}|u_{1}|\dots|u_{r}|$. Furthermore, since the connection is compatible with the metrics, it commutes with natural contractions, yielding a generalized Leibniz rule:
$\nabla_{X}(u_{1}*\dots*u_{r})=\displaystyle\sum_{i=1}^{r}u_{1}*\dots*(\nabla_{X}u_{i})*\dots*u_{r}.$
To study norm equivalence, we must understand how the Bochner Laplacian commutes with higher-order covariant derivatives. Unlike the Euclidean Laplacian, this commutator is non-zero and introduces curvature terms.
Lemma 13 (Commutator of the Laplacian and Covariant Derivatives).
Let $(M,g)$ be a Riemannian manifold and $E\to M$ a smooth vector bundle with a compatible connection $\nabla^{E}$. For any integer $m\geq 1$, there exists a finite linear combination of terms of the form $\nabla^{k}\mathrm{Rm}*\nabla^{m-k}u$ for $0\leq k\leq m$, such that for all $u\in\Gamma(E)$:
$\Delta_{B}(\nabla^{m}u)=\nabla^{m}(\Delta_{B}u)+\displaystyle\sum_{k=0}^{m}\nabla^{k}\mathrm{Rm}*\nabla^{m-k}u.$
Here, $\mathrm{Rm}$ denotes both the Riemannian curvature of $M$ and the curvature of the connection on $E$.
Proof.
We proceed by induction on $m$. First, we establish the identity for an arbitrary tensor field $u\in\Gamma(T^{(0,s)}M\otimes E)$ for $s\geq 0$:
$\Delta_{B}(\nabla u)-\nabla(\Delta_{B}u)=\mathrm{Rm}*\nabla u+\nabla\mathrm{Rm}*u.$ (28)
Fix $p\in M$ and choose geodesic normal coordinates centered at $p$, where $g_{ij}(p)=\delta_{ij}$ and the Christoffel symbols vanish. At $p$, the Bochner Laplacian simplifies to $\Delta_{B}v(p)=\displaystyle\sum_{i=1}^{n}\nabla_{\partial_{i}}\nabla_{\partial_{i}}v(p)$. Evaluating the commutator at $p$, we can add and subtract the mixed derivative term $\nabla_{\partial_{i}}\nabla(\nabla_{\partial_{i}}u)$ to obtain:
$\big(\Delta_{B}(\nabla u)-\nabla(\Delta_{B}u)\big)(p)$ $=\displaystyle\sum_{i=1}^{n}\Big(\nabla_{\partial_{i}}\nabla_{\partial_{i}}(\nabla u)-\nabla(\nabla_{\partial_{i}}\nabla_{\partial_{i}}u)\Big)(p)$
$=\displaystyle\sum_{i=1}^{n}\Big(\nabla_{\partial_{i}}\bigl([\nabla_{\partial_{i}},\nabla]u\bigr)+[\nabla_{\partial_{i}},\nabla](\nabla_{\partial_{i}}u)\Big)(p).$
To analyze the commutator $\nabla_{\partial_{i}},\nabla]$ acting on the tensor-valued section $u$, we expand its action on vector fields $X_{1},\dots,X_{s},X$. Because the connection is torsion-free and we are evaluating at the center of normal coordinates, the derivative terms cancel gracefully, leaving only the curvature operators. Intrinsically, this yields:
$\nabla_{\partial_{i}},\nabla]u=R^{E}(\partial_{i},\cdot)u-\displaystyle\sum_{\alpha=1}^{s}\iota^{(\alpha)}_{R^{M}(\partial_{i},\cdot)}u,$
where $\iota^{(\alpha)}$ denotes the interior multiplication on the $\alpha$-th component. In the language of our * operator, this shows that $\nabla_{\partial_{i}},\nabla]u=\mathrm{Rm}*u$. Applying the covariant derivative $\nabla_{\partial_{i}}$ to this expression via the Leibniz rule generates terms of the form $\nabla\mathrm{Rm}*u+\mathrm{Rm}*\nabla u$. The term $\nabla_{\partial_{i}},\nabla](\nabla_{\partial_{i}}u)$ is straightforwardly of the form $\mathrm{Rm}*\nabla u$. Substituting these back into the sum over $i$ completes the proof for the base case (28).
Assume the result holds for some $m\geq 1$. Applying the base case (28) to the tensor field $\nabla^{m}u$:
$\Delta_{B}(\nabla^{m+1}u)=\nabla\bigl(\Delta_{B}(\nabla^{m}u)\bigr)+\mathrm{Rm}*\nabla^{m+1}u+\nabla\mathrm{Rm}*\nabla^{m}u.$
By the inductive hypothesis:
$\nabla\bigl(\Delta_{B}(\nabla^{m}u)\bigr)=\nabla\left(\nabla^{m}(\Delta_{B}u)+\displaystyle\sum_{k=0}^{m}\nabla^{k}\mathrm{Rm}*\nabla^{m-k}u\right).$
Applying the Leibniz rule to the sum, the derivative distributes over the curvature and the section terms. Re-indexing the summations naturally yields a combination of terms of the form $\nabla^{k}\mathrm{Rm}*\nabla^{m+1-k}u$ for $0\leq k\leq m+1$. This completes the induction. ∎
7.3 Equivalence of Norms on Closed Manifolds
Theorem 14 (Equivalence of Norms via the Bochner Laplacian).
Let $(M,g)$ be a closed Riemannian manifold and $E\to M$ a smooth vector bundle equipped with a fiber metric $h_{E}$ and a compatible connection $\nabla^{E}$. For any integer $m\in\mathbb{N}$ and section $u\in H^{m}(E)$, we define the structured norm:
$\|u\|_{\Delta_{B}}^{2}:=\displaystyle\sum_{\begin{subarray}{c}j=0\\ j\text{ even}\end{subarray}}^{m}\displaystyle\int_{M}|\Delta_{B}^{\frac{j}{2}}u|^{2}\,d\lambda_{g}+\displaystyle\sum_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{m}\displaystyle\int_{M}\big|\nabla\Delta_{B}^{\frac{j-1}{2}}u\big|^{2}\,d\lambda_{g}.$
Then there exist constants $c,C>0$ such that
$c\|u\|_{H^{m}(E)}^{2}\leq\|u\|_{\Delta_{B}}^{2}\leq C\|u\|_{H^{m}(E)}^{2},\qquad\forall u\in H^{m}(E).$
Proof.
Since $M$ is a closed Riemannian manifold ($\partial M=\varnothing$), and $C^{\infty}(E)$ is dense in $H^{m}(E)$ (Theorem 8), it suffices to prove both inequalities for smooth sections $u\in C^{\infty}(E)$. In what follows, all integration by parts are performed without boundary terms.
Step 1: $\|u\|_{\Delta_{B}}^{2}\leq C\|u\|_{H^{m}(E)}^{2}$.
By definition, the Bochner Laplacian is the metric trace of the second covariant derivative: $\Delta_{B}=-\operatorname{tr}_{g}(\nabla^{2})$. A crucial observation is that, because the connection $\nabla$ is compatible with the Riemannian metric $g$ (i.e., $\nabla g^{-1}=0$), the covariant derivative commutes with the metric contraction operator $\operatorname{tr}_{g}$. Explicitly, for any tensor field $T$, the Leibniz rule implies $\nabla(\operatorname{tr}_{g}T)=\nabla(g^{pq}T_{pq})=(\nabla g^{pq})T_{pq}+g^{pq}\nabla T_{pq}=\operatorname{tr}_{g}(\nabla T)$.
Applying this commutativity iteratively for any integer $k\geq 1$, we obtain:
$\Delta_{B}^{k}u=(-1)^{k}\operatorname{tr}_{g}^{(k)}(\nabla^{2k}u),$
where $\operatorname{tr}_{g}^{(k)}$ denotes the composition of $k$ successive metric contractions. This implies that $\Delta_{B}^{k}u$ is simply a specific tensor contraction of $\nabla^{2k}u$, which in our notation can be written as $\Delta_{B}^{k}u=g^{-1}*\dots*g^{-1}*\nabla^{2k}u$. Since $M$ is compact, the metric components are uniformly bounded, so there exists $C_{k}>0$ such that $|\Delta_{B}^{k}u|_{h_{E}}\leq C_{k}|\nabla^{2k}u|_{g,h_{E}}$. Squaring and integrating over $M$:
$\int_{M}|\Delta_{B}^{k}u|_{h_{E}}^{2}\,d\lambda_{g}\leq C_{k}^{2}\displaystyle\int_{M}|\nabla^{2k}u|_{g,h_{E}}^{2}\,d\lambda_{g}.$
Similarly, for the odd terms, the commutativity of $\nabla$ with $\operatorname{tr}_{g}$ implies $\nabla\Delta_{B}^{k}u=(-1)^{k}\operatorname{tr}_{g}^{(k)}(\nabla^{2k+1}u)$. Bounding this in the exact same manner yields:
$\int_{M}|\nabla\Delta_{B}^{k}u|_{g,h_{E}}^{2}\,d\lambda_{g}\leq(C_{k}^{\prime})^{2}\displaystyle\int_{M}|\nabla^{2k+1}u|_{g,h_{E}}^{2}\,d\lambda_{g}.$
Summing over all $j\in\{0,\dots,m\}$ (separating into even and odd $j$), we conclude that each term in $\|u\|_{\Delta_{B}}^{2}$ is bounded by the corresponding $j$-th term in the standard Sobolev norm, establishing the first inequality.
Step 2: $\|u\|_{H^{m}(E)}^{2}\leq C\|u\|_{\Delta_{B}}^{2}$.
We proceed by induction on $j$ to show that $\int_{M}|\nabla^{j}u|^{2}\,d\lambda_{g}\leq C_{j}\|u\|_{\Delta_{B}}^{2}$ for all $j\in\{0,\dots,m\}$.
For $j=0$ and $j=1$, the terms $\int_{M}|u|^{2}\,d\lambda_{g}$ and $\int_{M}|\nabla u|^{2}\,d\lambda_{g}$ appear explicitly in the definition of $\|u\|_{**}^{2}$ for the even sum ($j=0$) and odd sum ($j=1$), respectively.
For $j=2$, applying Green’s formula (Corollary 12) to the bundle $T^{*}M\otimes E$ yields:
$\int_{M}|\nabla^{2}u|_{g,h_{E}}^{2}\,d\lambda_{g}=\displaystyle\int_{M}\langle\nabla u,\Delta_{B}(\nabla u)\rangle_{g,h_{E}}\,d\lambda_{g}.$
Using the Commutator Lemma 13 with $m=1$, we substitute $\Delta_{B}(\nabla u)=\nabla(\Delta_{B}u)+\mathrm{Rm}*\nabla u$:
$\int_{M}|\nabla^{2}u|^{2}\,d\lambda_{g}=\displaystyle\int_{M}\langle\nabla u,\nabla(\Delta_{B}u)\rangle\,d\lambda_{g}+\displaystyle\int_{M}\langle\nabla u,\mathrm{Rm}*\nabla u\rangle\,d\lambda_{g}.$
Applying Corollary 12 again (now to the bundle $E$) transforms the first integral into $\int_{M}|\Delta_{B}u|^{2}\,d\lambda_{g}$. For the second integral, since $M$ is compact, $|\mathrm{Rm}*\nabla u|\leq C|\nabla u|$. Thus:
$\int_{M}|\nabla^{2}u|^{2}\,d\lambda_{g}\leq\displaystyle\int_{M}|\Delta_{B}u|^{2}\,d\lambda_{g}+C\displaystyle\int_{M}|\nabla u|^{2}\,d\lambda_{g}.$
Both terms on the right are explicitly controlled by $\|u\|_{\Delta_{B}}^{2}$.
Assume now $j\geq 3$ and that the bound holds for all integers $l<j$. We write $\nabla^{j}u=\nabla^{2}(\nabla^{j-2}u)$ and apply the $j=2$ case just proved to the tensor field $\nabla^{j-2}u$:
$\int_{M}|\nabla^{j}u|^{2}\,d\lambda_{g}\leq\displaystyle\int_{M}|\Delta_{B}(\nabla^{j-2}u)|^{2}\,d\lambda_{g}+C\displaystyle\int_{M}|\nabla^{j-1}u|^{2}\,d\lambda_{g}.$
By the inductive hypothesis, the second term is bounded by $C\|u\|_{\Delta_{B}}^{2}$. To estimate the first term, we apply Lemma 13 with m=j2m=j-2:
$\Delta_{B}(\nabla^{j-2}u)=\nabla^{j-2}(\Delta_{B}u)+\displaystyle\sum_{k=0}^{j-2}\nabla^{k}\mathrm{Rm}*\nabla^{j-2-k}u.$
Squaring this expression and integrating, and noting that all covariant derivatives of $\mathrm{Rm}$ are uniformly bounded on the compact manifold $M$, we obtain:
$\int_{M}|\Delta_{B}(\nabla^{j-2}u)|^{2}\,d\lambda_{g}\leq C\displaystyle\int_{M}|\nabla^{j-2}(\Delta_{B}u)|^{2}\,d\lambda_{g}+C\displaystyle\sum_{l=0}^{j-2}\displaystyle\int_{M}|\nabla^{l}u|^{2}\,d\lambda_{g}.$
By the inductive hypothesis, the sum over $l$ is bounded by $C\|u\|_{\Delta_{B}}^{2}$. Therefore,
$\int_{M}|\nabla^{j}u|^{2}\,d\lambda_{g}\leq C\displaystyle\int_{M}|\nabla^{j-2}(\Delta_{B}u)|^{2}\,d\lambda_{g}+C\|u\|_{\Delta_{B}}^{2}.$
We now apply the inductive hypothesis directly to the section $\Delta_{B}u$, since j2<jj-2<j:
$\int_{M}|\nabla^{j-2}(\Delta_{B}u)|^{2}\,d\lambda_{g}\leq C\left(\displaystyle\sum_{\begin{subarray}{c}\ell=0\\ \ell\text{ even}\end{subarray}}^{j-2}\displaystyle\int_{M}\bigl|\Delta_{B}^{\frac{\ell}{2}}(\Delta_{B}u)\bigr|^{2}\,d\lambda_{g}+\displaystyle\sum_{\begin{subarray}{c}\ell=1\\ \ell\text{ odd}\end{subarray}}^{j-2}\displaystyle\int_{M}\bigl|\nabla\Delta_{B}^{\frac{\ell-1}{2}}(\Delta_{B}u)\bigr|^{2}\,d\lambda_{g}\right).$
Re-indexing these sums via $r=\ell+2$ (for the even sum) and $s=\ell+2$ (for the odd sum), we find:
$C\left(\displaystyle\sum_{\begin{subarray}{c}r=2\\ r\text{ even}\end{subarray}}^{j}\displaystyle\int_{M}\bigl|\Delta_{B}^{\frac{r}{2}}u\bigr|^{2}\,d\lambda_{g}+\displaystyle\sum_{\begin{subarray}{c}s=3\\ s\text{ odd}\end{subarray}}^{j}\displaystyle\int_{M}\bigl|\nabla\Delta_{B}^{\frac{s-1}{2}}u\bigr|^{2}\,d\lambda_{g}\right).$
These terms are exactly those that appear in the definition of $\|u\|_{\Delta_{B}}^{2}$. Thus, $\int_{M}|\nabla^{j}u|^{2}\,d\lambda_{g}\leq C_{j}\|u\|_{\Delta_{B}}^{2}$. Summing over $j\in\{0,\dots,m\}$ completes the induction and the proof. ∎
An immediate and highly useful consequence of this theorem is the corresponding norm equivalence for scalar functions on closed manifolds, which recovers the natural generalization of (26) via the Laplace–Beltrami operator.
Corollary 15 (Norm Equivalence for the Laplace–Beltrami Operator).
Let $(M,g)$ be a closed Riemannian manifold. For any integer $m\in\mathbb{N}$ and any scalar function $u\in H^{m}(M)$, the standard scalar Sobolev norm
$\|u\|_{H^{m}(M)}^{2}:=\displaystyle\sum_{j=0}^{m}\displaystyle\int_{M}|\nabla^{j}u|_{g}^{2}\,d\lambda_{g}$
is equivalent to the structured norm defined via the Laplace–Beltrami operator $\Delta_{g}$:
$\|u\|_{\Delta_{g}}^{2}:=\displaystyle\sum_{\begin{subarray}{c}j=0\\ j\text{ even}\end{subarray}}^{m}\displaystyle\int_{M}|\Delta_{g}^{\frac{j}{2}}u|^{2}\,d\lambda_{g}+\displaystyle\sum_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{m}\displaystyle\int_{M}\big|\nabla\Delta_{g}^{\frac{j-1}{2}}u\big|_{g}^{2}\,d\lambda_{g}.$
Proof.
Consider the trivial line bundle $E=M\times\mathbb{R}$ equipped with the standard Euclidean metric on the fibers and the standard flat connection. Smooth sections of $E$ are naturally identified with smooth scalar functions $u\in C^{\infty}(M)$. Under this identification, the compatible connection $\nabla^{E}$ reduces to the standard Riemannian connection (with $\nabla^{E}u$ corresponding to the differential $du$, whose pointwise norm coincides with that of the gradient $\nabla u$), and the fiber metric $h_{E}$ is simply standard scalar multiplication.
Consequently, the Bochner Laplacian $\Delta_{B}=\nabla^{E*}\nabla^{E}=-\operatorname{tr}_{g}(\nabla^{2})$ acting on sections of $M\times\mathbb{R}$ coincides exactly with the positive-definite Laplace–Beltrami operator $\Delta_{g}=-\operatorname{div}\circ\nabla$ acting on scalar functions. The result then follows immediately from applying Theorem 14 to this specific bundle. ∎
Declarations
  • Funding: The authors did not receive support from any organization for the submitted work.
  • Conflict of interest/Competing interests: There is no conflict of interest related to the work submitted for publication.
  • Ethics approval and consent to participate: All authors have approved and consented to participate in the work submitted for publication.
  • Consent for publication: All authors have approved and consented to submit the work for publication.
  • Data availability: Not applicable.
  • Materials availability: Not applicable.
  • Code availability: Not applicable.
  • Author contribution: All authors contributed to the study conception and design of the work submitted for publication. The first draft of the manuscript was written by Carlos D. Velázquez-Mendoza and all authors commented on previous versions of the manuscript. All authors read and approved the final manuscript.